A Complete Bibliography of Publications in the Fisheries and Aquaculture Journal

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

10 May 2024 Version 1.07

Title word cross-reference

2 [200]. _{4} [200]. × [154, 120].

* [200].

-A [342].

1 [194].

2020 [259, 251]. 2021 [258]. 2150-3508 [235].

9th [257].

Abalone [129]. Abay [201]. Aberrations [220]. Abundance [201, 229, 300, 321, 103, 5, 30]. Access [37]. Access-Point [37].

Acclimation [359]. Acid [154, 236]. Acids [63]. Activation [311]. Active [164]. Activities [120]. Activity [245, 249, 225]. Acute [171, 200]. Adamawa [240]. Adaptability [317]. Adaptations [327, 335, 330]. Added [249]. Additives [209]. Adhesive [108]. Adjacent [211]. Adult [208, 260, 312]. Adults [172]. Advancement [279]. Advances [262]. Advantages [310]. aegyptensis [106]. aequippinatus [80]. Aeromonas [118]. Affects [149, 191]. AFLP [61]. Africa [195, 56]. African [32, 22, 20, 227, 178, 165, 197, 10, 172, 78, 100]. Africocypha [178]. After [121]. Against [220, 3]. Agbokim [10]. Age [164, 24, 14, 183]. Agriculture [190]. Agro [211]. Agro-System [211]. Ahaha [188]. Akomoje [39]. Al [13]. Al-Seeb [13]. Alabama [29]. Algal [132]. Allahabad [214]. alliacea [178]. Alloherpesvirus [127]. Almix [136]. Alone [51]. Along [363, 79]. Alternatives [170]. Alters [246]. Aluminosilicate [32]. Amazon [102]. American [311, 119]. americanus [119]. Amino [236]. Ammonia [171]. Among [252, 53, 24, 144]. Amplified [207]. Anabantid [165]. Anabas [152, 136]. Anabus [90]. Analyses [79, 102]. Analysis [344, 308, 259, 114, 243, 372, 61, 365, 302, 221, 145]. Anatomy [332, 336]. Anchovy [162]. Andassa [201]. Andhra [303]. Angelfish [358]. anguillaris [189]. Anicut [183, 182]. Animal [127, 67, 51]. Animals [81, 326]. anisocapsularis [195]. annulus [174]. Antenna [288, 314]. Anterior [108]. Antibacterial [245]. Antibiotic [206]. Antioxidant [231]. Any [127]. Apex [339]. Aphanomyces [3]. Apparatus [108, 266]. Apparent [219]. Applicable [34]. Application [211, 308, 73, 25, 155, 92]. Applications [46, 296]. Applied [48]. Approach [19, 115, 342]. Aqua [67]. Aquaculture [211, 176, 54, 86, 82, 293, 280, 309, 279, 212, 168, 281, 197, 243, 372, 366, 368, 47, 8, 87, 62, 295, 190, 256, 370, 316, 98, 297, 44, 73, 12, 89, 257, 296, 306, 289, 324, 294, 320, 248, 216, 235, 342, 6, 346, 278, 84, 74, 310, 274, 268, 173, 52, 252, 99, 203, 92, 178, 265, 258]. Aquaponic [140, 283]. Aquatic [362, 127, 318, 81, 116, 50, 93, 172, 326, 333, 57, 51]. Arabian [11]. Ardibo [199]. Area [305]. Areas [168, 367]. arriving [258]. Article [347]. Artisanal [56, 228]. Asa [175]. Ascocotyle [77]. Asian [54]. Aspects [201, 242]. Assam [114]. Assay [42]. Assemblages [96]. Assembly [145]. Assesment [110]. Assess [115]. Assessing [233]. Assessment [369, 56, 281, 322, 159, 25]. Assessments [228]. Associations [173]. Astaxanthin [121]. Aulia [161]. auritus [193]. Austria [257]. Authority [49]. Authorization [67]. Availabilities [63]. Award [256, 251]. axelrodii [79]. Azadirachtin [3].

Bacteria [304, 202, 156]. Bacterial [2, 295, 198, 356]. Bagridae [188]. Bahir [153]. Bangladesh [176, 123, 177, 192, 300, 321, 103, 126, 115, 117, 75, 216, 166]. Bangladeshi [254]. Bardawil [223]. Bark [206]. Barley [150]. Barracuda [205, 339]. Barramundi [1]. barteri [206]. Based [88, 249, 174, 9, 49, 48, 74, 125, 142, 236]. Basin [201, 237]. Bass [242, 24].

Be [84, 258]. Bead [8]. Behavior [337]. Behavioral [330]. Behaviour [338, 359, 178]. Behavioural [355]. being [180]. Beluga [28]. Beneficial [132]. benefit [114]. Benefits [309, 346, 333]. Bengal [138]. Benthic [194, 229, 158]. Benue [167]. Bervoortia [145]. Beta [90]. Better [342]. between [76]. Bhola [216]. Big [205]. Bigeye [193]. Bio [350, 305, 8, 152, 84]. Bio-Filter [8]. Bio-Manipulation [84]. Bio-Synthesis [152]. Bioaccumulation [175, 177]. Biochemical [3, 146, 186]. Biodiversity [369, 19, 196]. Biohazards [35]. Biological [201, 57]. Biology [13, 352, 128, 10, 182, 204]. Biomarkers [111]. Biomass [33, 220, 104]. Biosecurity [297]. Biota [318]. Biotechnological [46]. Bivalve [62]. Bivalvia [208]. Bivalvulida [195]. Bladder [336]. Bloch [137, 136]. Blooms [132]. Blue [308, 201]. Bodies [322, 98, 307, 41]. Body [88, 167, 23, 142, 288, 314]. Bomb [187]. Bone [220]. Both [34]. Botswana [58]. Brachydeuterus [193]. Brachyura [15]. Brazil [305, 102]. Bred [311, 186]. Breeding [72]. Brief [277, 315, 275, 290, 285, 338, 332, 273, 276, 274]. Brown [224, 284]. Budgets [29]. Building [316]. Burchell [32, 20, 227, 207, 91, 142]. buthupogon [175]. Butipearl^{T M} [154]. Butyric [154]. By-catch [56]. By-Products [213].

C [245]. C-Type [245]. Ca [152]. Ca/P [152]. Cadmium [166]. Cage [168]. Cake [32]. calcarifer [1]. Calcium [32]. California [233]. Cambaridae [88, 55]. Cameroon [130, 195, 56, 165, 237, 240]. Campus [300, 321]. Can [84]. Candidate [87]. Capacity [231]. Cape [194]. capito [225]. Captivity [2]. Capture [79]. Captured [137]. Carbohydrate [71]. Carbohydrates [184]. Carcass [219]. Cardinal [79]. carota [90]. Carotenoids [154]. Carp [150, 71, 10, 17, 182, 72, 215, 183]. carpio [146, 113, 150, 74, 198, 215]. Carps [181]. Casamance [194]. Case [218, 212, 281, 222, 157, 348, 199]. Caspian [7]. caspicus [7]. Cast [363]. Castor [227]. Catch [10, 191, 56]. Catching [161, 114]. Catfish [32, 22, 20, 227, 311, 154, 197, 263, 38, 140, 172, 78, 282, 323, 100, 317, 52, 178]. Catla [315, 181]. Caught [198]. Caused [135]. CBNRM [49]. Cell [91]. Cells [220]. Census [37]. Center [172]. Central [195, 217, 139]. Centuries [157]. cephalus [18, 139]. Ceratomyxa [36, 40]. Cessation [121]. Chain [101, 365, 282, 323]. Challenges [141, 318, 250]. Change [252]. Changes [164, 38, 7, 3, 14, 107, 179, 353]. Changing [191, 348]. Channel [311]. Chanos [16]. Chaos [101]. Chapai [176]. Characteristics [363, 240, 7, 338, 331, 332]. Characterization [106, 96]. Characters [368, 319]. Chefa [31]. Chelonia [2]. Chemical [39, 33, 240, 228, 96, 221, 107]. Chhattisgarh [302]. Chile [59]. Chilean [59]. chilensis [59]. China [50, 92]. Chinook [36, 40]. Chittagong [300, 321]. Chloramin [135]. Chloramin-T [135]. Chromosomal [220]. Chrysichthys [228, 188]. Chub [18]. Chum [131]. Cichlidae [260, 312, 358]. Circadian [162]. Circles [66]. Circulating [8]. Cirrhina [3]. Cirrhinus [164, 128, 183, 182, 181]. City [75, 365]. Clade [218]. Clarias [32, 20, 227, 178, 175, 207, 154, 305, 189, 38, 228, 172, 91, 206, 78, 142, 100, 198]. Classical [58]. Clearwater [55]. Climate [187, 348, 252, 353]. Climatic [43]. Climbing [152]. Clymene [221]. Cnidaria [237]. Coast [105, 11, 205, 60, 30, 110, 139]. Coastal [168, 228, 222, 115, 221, 160, 148, 217]. Cod [318]. Coefficient [120]. Cold [51]. Collected [106, 175, 168, 185, 225]. Collection [19]. Color [90]. Coloration [121]. Colossoma [184]. columnaris [304]. Combining [102]. Commentary [280, 349, 354, 361, 270, 262, 356]. Commercial [105, 272]. Commercialization [173]. commerson [191]. Common [150, 357, 34, 215]. Commonly [166, 198]. Communal [49]. communis [227, 113]. Communities [316, 144]. Community [299, 246, 115, 241, 49]. Comparative [344, 185, 186]. Comparison [37, 23]. Comparisons [24]. compete [26]. Components [208]. Composition [201, 219, 33, 300, 321, 271, 313, 185, 103, 142, 95]. Comprehensive [369]. Concentration [192, 261, 166]. Concentrations [221]. Concept [34]. Concrete [25]. Condition [88, 189, 137, 188, 152, 241, 42]. Conditions [43]. Conference [370, 257]. Conservation [318, 250, 325, 128, 329, 268, 160]. Constrains [267]. Constraints [144]. Construction [278, 203]. Consumed [166]. Consumer [50]. Consumption [50]. Containing [3]. Content [28, 188, 221]. Contents [71]. Contribution [133]. Control [279, 294, 178]. Controlling [101]. Conventional [71]. cooperative [302]. Cooperatives [322]. Cope [298]. Copper [200]. Coptodon [242]. Coral [353]. Correlation [28]. Cost [114, 144, 51]. Cost-benefit [114]. Costal [267]. Cottonseed [219]. coubie [10]. County [212]. Coupled [283]. Cowry [174]. Crab [61, 179]. Crabs [327]. Craft [123]. Crayfish [88, 55]. Creature [298]. Creek [222]. Creel [37]. Criteria [108]. Critical [49]. Croaker [234]. crocea [234]. Crops [46]. Cross [188]. Crustacean [134]. Ctenopharyngodon [72]. Ctenophores [95]. Ctenopoma [165]. Cultivated [46]. Cultivation [320]. Cultural [94]. Culture [290, 156, 129, 217]. Cultured [207, 63, 25, 42]. Cumilla [241]. Current [83, 322, 50, 181, 92]. CuSO [200]. Cutthroat [26]. Cuvier [205]. CYP19 [74]. Cyperus [20]. Cypraea [174]. Cypraeidae [174]. Cyprinid [74]. Cyprinus [150, 146, 113, 74, 198, 215]. Cyprus [31].

D [163]. D. [106]. Dal [113]. Dam [240]. Dams [96]. Danio [171, 155, 138, 80]. Dar [153]. Data [255]. date [258]. Daucus [90]. Dawn [46]. Decapoda [88, 222, 15, 55]. Decapterus [159]. Decline [113]. Deep [212, 273]. Deep-Sea [273]. Degraded [95]. Delta [190, 58]. Denmark [95]. Densities [95, 70]. Density [78, 217]. Desalination [132]. Description [195, 237]. Design [363]. Detection [21]. Determinants [252]. Determination [71, 117, 74]. Developing [212]. Development [222, 249, 316, 98, 6, 15]. Developmental [208]. Devices [114]. Devil [288, 314]. Dhaka [123]. Dhanusha [345]. Dhonagoda [196]. Diagnostic [279]. diesing [108]. Diet [32, 169, 3, 246, 91, 142, 183]. Dietary [22, 20, 121, 154, 146, 260, 312, 120, 231]. Diets [18, 219, 249, 17, 184, 100, 236]. Differences [368, 76]. Different [16, 63, 169, 224, 190, 185, 155, 70, 217, 184, 166, 310, 107]. Digestibility [219, 120, 231]. Diplectanum [108]. Diplodus [110]. Diplozoon [106]. Direction [254]. Disease [351, 277, 279, 85, 93, 294, 179, 347, 345, 343, 304, 303]. Diseases [356]. Disinfectant [135]. Distribution [165, 10, 30]. District [176, 202, 216, 138, 302, 345, 344]. Diversification [141, 317]. Diversity [369, 207, 201, 123, 27, 300, 340, 321, 246, 238, 60]. DNA [207, 207, 152]. Dolphin [221]. Dorsal [24]. Downstream [181]. dragonfly [178]. Dry [18]. Duckweed [100]. due [305]. During [7, 95, 229, 42, 358]. Dynamic [110]. Dynamics [193, 4, 86, 33, 36, 40, 11, 34, 160].

Early [186]. Earthen [142]. East [105, 303, 60, 30, 343]. Eastern [5, 80]. Ebonyi [229, 229]. ECM [208]. Ecological [315, 19, 329, 326, 319]. Ecology [328, 360]. Economic [325, 228, 320, 302, 241, 102]. Economics [210]. Economy [308, 159]. Ecosystem [362, 26, 116, 194, 48]. Ecosystem-Based [48]. Ecosystems [367, 57]. Editor [122]. Editorial [362, 27, 272, 238, 235]. edulis [42]. Edwards [119]. Edwardsiella [143, 226]. Eel [320]. Effect [150, 16, 154, 220, 240, 1, 7, 90, 155, 248, 216, 109, 261, 215, 203]. Effective [144]. Effectiveness [8]. Effectors [249]. Effects [161, 171, 147, 2, 219, 62, 224, 146, 9, 206, 124, 70, 236, 353, 231]. Efficacy [114]. Efficiencies [1]. Efficiency [190]. Egg [311]. Eggs [172, 30]. Egypt [53, 223, 225]. Electric [354]. Element [12]. Eleyele [137, 180]. Emerging [87, 85]. Emphasis [88, 128]. Encapsulated [154]. Energy [1, 184]. Engraulis [162]. Enteric [202]. Environment [327, 62, 76, 97, 348]. Environmental [2, 305, 275, 53, 84]. Environmentally [65, 95]. Enzyme [225, 120]. Epigenetic [81]. Eriocheir [179]. esculentus [20]. Estimating [24]. Estimation [199, 255, 234]. Estuarine [4]. Ethiopia [141, 199, 201, 281, 250, 33, 322, 202, 143, 226, 172, 133, 261, 260, 312, 198, 31, 203, 343, 153]. Ethiopian [368]. Europe [259]. European [372, 67]. Evaluating [152]. Evaluation [41, 311, 220, 271, 313, 91, 153]. Evidence [298]. Examine [266]. Excellent [163]. Existing [176]. experimental [118]. Exploitation [130, 60, 233, 148]. Exposure [136]. Expression [124]. External [327, 343]. Extracellular [208]. Extract [213, 231, 178]. Extracts [271, 313]. Eye [205].

Faced [326]. Facility [42]. Factor [189, 137, 188]. Factors [33, 179, 124]. Failing [68]. Failure [247]. Fake [187]. Family [94]. Far [54]. fario [224, 284]. Farm [170, 177, 72]. Farmed [177, 192, 163, 239]. Farmer [173]. Farmers [176, 252]. Farming [232, 349, 263, 187, 292, 291, 262, 310, 345]. Farms [117, 345, 303, 119]. fasciata [129]. Fasting [358]. Fatty [63]. Fauna [344, 229, 158]. Faunal [238]. feasibility [59]. Features [327, 162]. Fecundity [189]. Fed [227, 142, 184]. Federal [147]. Feed [253, 209, 71, 219, 146, 142, 272]. Feeding [128, 249, 214, 284, 113, 248].

Feeds [177, 192, 239, 67]. Fermented [287]. Fertilization [311]. Field [16]. Fillet [121]. Fillets [77]. Filter [8]. Fin [24]. findings [118]. Finfish [30]. Fingerlings [22, 91, 206, 152, 120]. Fins [335]. Fish [211, 253, 344, 161, 176, 35, 351, 147, 167, 277, 232, 45, 83, 349, 287, 318, 335, 168, 114, 111, 8, 213, 87, 27, 135, 240, 112, 355, 128, 228, 202, 158, 340, 130, 285, 338, 246, 226, 210, 146, 291, 365, 85, 117, 134, 12, 307, 113, 5, 142, 273, 270, 272, 320, 144, 286, 283, 72, 302, 341, 80, 269, 310, 334, 173, 319, 198, 125, 252, 215, 288, 314, 99, 356, 345, 343, 153, 200, 242, 301]. Fish-Handler [277]. Fisheries [141, 199, 56, 37, 259, 244, 372, 367, 151, 370, 58, 98, 49, 257, 64, 6, 84, 266, 69, 235]. Fisherman [241]. Fishermen [322, 210, 126, 133, 302]. Fishery [41, 35, 43, 250, 157, 172, 191, 233, 102]. Fishes [106, 195, 201, 23, 354, 75, 337, 329, 166]. Fishing [369, 123, 299, 164, 275, 371, 157, 97, 144, 264, 125]. Fishpond [203]. Fishway [357]. Flatfish [76]. Floating [8]. Floodplain [58]. Flora [238]. Food [35, 101, 214, 25, 113, 333]. Foods [129]. Forecast [131]. Forehead [288, 314]. Formulated [17, 239]. Forsskål [16]. forsteri [205]. Four [130]. Foxhunt [299]. Framework [98]. Free [364, 347]. Fresh [83, 228]. Freshwater [195, 199, 128, 202, 146, 329, 215, 200, 358]. Friendly [19, 65]. Fry [135, 78, 178]. Full [288, 314]. Function [230]. Functional [209]. Functions [335, 336, 341]. Fungus [172]. furcatus [188]. Future [116, 254].

Gache [49, 49]. galbana [63]. Gambia [64]. Ganga [164]. gariepinus [32, 20, 227, 178, 207, 154, 305, 228, 172, 91, 206, 78, 142, 100, 198]. Gastropods [174]. Gear [123, 164]. Gene [124]. Genetic [207, 364, 61, 234]. Genetically [239]. Genetics [4, 270, 268]. Genotoxicology [116]. Geographically [103]. Ghana [193, 19, 348, 221]. GIFT [239]. Gilgel [201]. Gill [108, 130]. Gillnet [205]. Gills [165, 341]. Ginger [271, 313]. Girard [88]. Global [99]. GLO^{T M} [154]. Godavari [303]. Gold [174]. Golden [118]. Goldlined [11]. Gomastapur [176]. Gonads [38]. Gonialosa [214]. Good [66]. Government [147]. Gracilaria [129]. Gradient [130]. Gram [202, 198]. Gram-Negative [198]. Grass [72]. Gray [221]. Greater [126]. Greek [69]. Green [2]. Grey [223]. Groundnut [32, 219]. groups [302]. grow [109]. grow-out [109]. Growing [333]. Growth [150, 227, 16, 154, 18, 219, 164, 169, 76, 137, 146, 185, 17, 25, 78, 142, 230, 124, 109, 70, 184, 148, 120, 234, 231, 242, 183, 119]. Grunt [193]. Guam [364]. Guapimirim [305]. Guder [281]. Guinea [348]. Gulf [204, 48, 233, 348]. Günther [237, 188, 180]. guptai [106]. Gut [188]. Gynogenetic [155].

H [200]. Habit [214]. Habitat [55]. Habitation [348]. Habitats [104, 276]. Habits [113]. Haemato [146]. Haemato-Biochemical [146]. Haematological [206, 14]. Halibut [332]. Haliotis [129]. Ham [214]. Hamilton [128, 17, 138, 183, 182]. Handler [277]. Haramaya [202]. Harmful [132]. Harvest [41, 230, 153]. Harvesting [361, 292, 101, 57, 112]. Harvests [46]. harveyi [218]. Hatchery [59, 105, 61, 172, 270, 186].

Hatchery-Bred [186]. Hatchlings [70]. Hayiq [198]. Health [45]. Heat [155, 77]. Heavy [177, 192, 117, 166]. Helminthiasis [202]. Help [367]. Hematological [3]. Henneguya [165]. Hepatic [120]. Hepatocytes [124]. Hepatopancreas [179]. Hepsetus [137]. Heptahydrate [200]. Herbal [3]. Herpesvirus [127]. Herpesviruses [127]. Heterobranchus [175]. hidden [301]. High [1, 288, 314]. Highland [261]. Hilsa [126]. Histological [38]. Histopathological [136]. Holding [42]. Homarus [119]. Homeostasis [283]. Honduras [230]. Hormone [72]. Horticulture [372]. host [304]. Housing [51]. Hupe [59]. Huso [28]. Hybrid [154, 120]. Hybrids [26]. Hydrated [32]. hydrophila [118]. Hydroponic [150]. Hypersaline [132]. Hypophthalmichthys [72]. hypophthalmus [177]. Hypophysial [286]. Hypothalamic [286].

Ibadan [282, 323]. Ice [215]. Iceland [232]. Ichthyofauna [305]. Ichthyofaunal [369]. Ichyofaunal [123]. Ictalurus [311]. idella [72]. Identification [364, 143, 172]. II [200]. Ikwori [5]. ilisha [126]. Ilorin [175]. imaginis [34]. Immune [154]. Imo [158]. Impact [56, 299, 275, 159, 359, 17, 95]. Impacts [53, 125]. Implementation [196]. Implications [193, 88, 97, 320, 160]. Importance [315, 325, 326, 319, 99]. Important [228, 103, 341]. Imported [105]. Imprints [81]. Improved [41, 210, 239]. Improvement [31]. Impulse [286]. IMTA [342]. In-season [131]. Inactivation [149]. Inclusion [227]. Income [210]. Index [55, 42]. Indexes [186]. India [106, 369, 114, 164, 105, 214, 107, 205, 138, 60, 183, 182, 80, 30]. Indian [13, 128, 222, 159, 17, 152]. Indicating [76]. Indicators [125]. Indices [28, 300, 321]. indicus [222]. Indigenous [253]. Indonesia [308, 217]. Induced [149, 72]. Indus [181]. Industry [43, 232, 310]. Inefficient [157]. Infecting [195]. Infection [218, 36, 40, 135, 226, 14, 118]. Infections [127]. Infectious [94, 294]. Infestation [53]. Influence [18, 78, 5, 120]. Influences [266]. Injury [149]. Injury-Induced [149]. Inland [58, 98, 12, 29, 244]. Innovations [318]. Insights [275, 264]. Installation [59]. Insulin [124]. Insulin-like [124]. Integrated [309, 190, 342]. Intellectual [44]. Intensive [211]. Interaction [307]. interactions [304]. Intermediaries [173]. Internal [343]. International [370, 257]. Intestine [136]. Introduction [210, 116]. intrusive [23]. invadans [3]. Invasion [305]. Invasive [104]. Inventory [123]. Invertebrate [246, 89]. Investigating [284]. Investigation [49, 196]. Investigations [53]. Ionic [28, 7]. ionized [171]. IPNV [135]. Iran [284]. Iranian [151]. IRIS [194]. IRIS-1 [194]. Iro [125]. Islands [174]. Iso [63]. Isochrysis [63]. Isolation [143, 172]. Isotopic [76]. Issue [87]. Issues [44]. Italy [85]. Items [71].

J&K [72]. Jalpaiguri [138]. Japanese [162, 34]. japonicus [162]. Java [217]. Jebel [161]. Jellyfish [95]. Jemma [141]. Journal [235]. Juba [365]. Juvenile [2, 219, 36, 40, 1, 100, 260, 312, 234]. Juveniles [169].

Karala [138]. Kariba [49]. Kashmir [107, 113, 72, 106]. kashmirensis [106]. Kemissie [31]. Kenya [222, 242, 307, 204]. kept [2]. Kerala [369]. KHV [127]. Khyber [344]. Kilifi [222]. Killifish [9]. kingsleyae [237]. Knowledge [176, 44]. Koi [127]. Kokanee [255]. Kunnu [22]. Kuriftu [33].

L. [20, 227, 368, 185, 25, 74, 260, 312, 42]. Labeo [169, 10, 17, 181]. Laboratory [16, 111]. Lagoon [223]. Lagos [228, 144]. Lake [369, 189, 137, 225, 107, 113, 5, 204, 199, 111, 33, 143, 210, 242, 133, 260, 312, 198, 125]. Lakes [37, 202, 261, 343, 368, 59]. Land [217]. Landing [288, 314]. Lands [49]. Langano [143]. Lannea [206]. Large [234]. Largemouth [242, 24]. Larimichthys [234]. Larvae [18, 15]. Larval [19]. Lates [1]. Latest [210]. latifolia [31]. lazera [38]. Lead [166]. Learning [68]. Leaves [206]. Lectin [245]. Left [76]. Leg [217]. Leiomyoma [247]. Lemna [100]. Length [88, 75, 188, 180]. Lentic [26]. Lessons [58]. Leuciscus [18]. Levels [227, 16, 2, 169, 146, 184, 125, 231]. Liberia [160]. Liberian [148]. lichenoides [129]. Lichtenstein [358]. Life [350, 172, 330]. like [124]. Likelihood [255]. Limfjorden [95]. Limitations [357]. Limiting [12, 144]. Limnology [244]. Linnaeus [110]. Lip [223]. Lipid [71, 169]. Lipids [184]. Litopenaeus [170, 249, 105, 109, 217]. Live [18, 269]. Livelihood [141, 126, 133, 31, 241]. Liver [225, 152]. Living [369, 298]. Liza [223]. Loach [245, 231]. Load [79]. lobster [119]. Local [185]. Locally [311]. Locations [185]. Logit [252]. longa [77]. longiceps [13]. longifilis [175]. Longitudinal [130]. Loose [109]. Low [12, 144, 51]. Low-Cost [144]. Lowe [148]. Lower [363, 39, 167, 183, 182]. LSS [109]. lucioperca [14].

Mackerel [191]. Macro [238]. Macrobenthic [115]. macrocephalus [154]. Macrophyte [246]. macropomum [184]. maderensis [148]. magai [195]. Maggot [91]. Mahseer [118]. Main [281]. Major [112, 17, 134, 267, 181]. Making [31]. Makurdi [167]. Malaysia [197]. Mammal [348]. Mammals [56]. Managed [37]. Management [193, 88, 250, 213, 58, 44, 93, 49, 11, 223, 48, 334, 252, 160, 303]. Manasbal [72, 107]. Mangrove [238]. Manipulation [84]. manmina [214]. Manzala [111]. MAPE [240]. Mardan [344]. Marginal [217]. Mariculture [267, 333]. Marine [352, 23, 350, 194, 367, 360, 159, 12, 142, 276, 348, 262]. Mariotteya [53]. Markers [61]. Market [66, 259, 243, 372, 266]. Marketing [133]. Markets [228]. Marrow [220]. Mass [53, 80]. Matrix [208]. Mattress [31]. Maturation [105, 182]. Maximum [255]. maya [70]. Meal [20, 87, 91, 100]. Meals [219]. Means [141, 31]. Measurements [167]. Measures [279, 297]. Measuring [23]. Meat [224]. Media [311]. Mediterranean [69]. melanostictus [34]. melanotheron [228]. Melen [130]. Menhaden [145]. Metabolism [184]. Metal [175]. Metals [177, 192, 117, 166]. Methods [369, 23, 116, 371, 292, 24, 97]. Metropolis [282, 323]. Mexican [139]. Micro [8]. Micro-Bead [8]. Microbial [228, 153]. Microbiological [83]. Microemulsified [154]. Microfauna [194]. Microfloral [27]. Microhabitat [138]. Microorganisms [280]. Micropterus [242]. Microsatellite [207, 364]. Migration [7, 286]. Milkfish [16]. Milne [119]. Mini [218, 279, 6, 65]. Mini-mussels [65]. Mini-Review [279, 6]. Misconceptions [357]. Mismanagement [69]. Mitigation [62]. Mitten [179]. Mobility [311]. Model [127, 170, 98, 101, 149, 252]. Modeled [37]. Modulation [147]. Molecular [106]. molitrix [72]. Mollucas [174]. Monitoring [116]. monodon [104]. Monogenean [108]. Monogeneans [130]. Monosex [74]. Moringa [206]. Mormyridae [237]. Morphogenesis [21]. Morphological [106, 174, 339]. Morphometrics [88]. Morphometry [167]. Mortalities [53]. Mortality [135, 148]. mossambicus [25, 120]. Motility [28, 162]. Mouldy [32]. mrigala [164, 3, 181]. Mud [78, 100, 178]. Mugil [139, 225, 77]. Mugilidae [139]. Mullet [223, 77]. Multi [309, 216, 342]. Multi-Ownership [216]. Multi-Trophic [309, 342]. Multinomial [252]. Multitrophic [119]. Mungeli [302]. Muradnagar [241]. Muscadomestica [91]. Muscle [152]. Mushroom [91]. Mussel [59, 187, 119]. mussels [65]. Muthupettai [30]. Mycobateriosis [85]. mydas [2]. Myeloperoxidase [149]. mykiss [135, 359]. Mytilus [59, 42]. Myxidium [195]. Myxobolidae [237]. Myxobolus [195, 165, 240]. Myxosporea [195]. Myxosporean [240]. Myxosporidia [237].

Nadu [183, 182]. Nagaland [114]. Naivasha [242, 210]. Nan’ao [212]. Nanotechnology [6]. National [159, 172, 72]. Natural [150, 129, 49, 25]. Nawabgonj [176]. near [75]. Necrosis [179]. Need [58]. Negative [202, 198]. Negro [102]. Nepal [345]. Net [363, 191]. Neurosecretory [286]. Neutral [42]. Neutrophil [149]. Newfoundland [42]. News [187]. NFSF [72]. Nickel [166]. Nigeria [39, 253, 175, 207, 147, 167, 197, 189, 229, 228, 158, 10, 137, 5, 104, 282, 323, 180, 144, 52, 252]. nigrodigitatus [228]. Nile [161, 201, 281, 368, 210, 185, 230, 260, 312]. niloticus [192, 368, 240, 156, 130, 210, 185, 53, 230, 74, 260, 312, 198, 120]. Nitrate [63]. Nitrogen [29, 57]. Noakhali [126]. Nodaviruses [168]. Non [71, 23, 67]. Non-Conventional [71]. Non-intrusive [23]. Non-Ruminant [67]. Nonlethal [24]. North [110, 223, 343, 141, 133, 80]. Northeastern [42]. Northern [199, 284, 138, 55]. Note [362, 351, 352, 293, 277, 122, 315, 275, 263, 290, 285, 338, 306, 332, 324, 273, 276, 272, 346, 274]. Nov [195]. Novel [6]. Novo [145]. Nutrient [219, 185]. Nutrients [261]. Nutrigenomics [47]. Nutrition [209, 301]. Nutritional [9, 152]. Nutritive [16]. Nyanza [204]. nymphs [178]. Nyong [237].

O [200]. obscura [180]. Obtain [230]. Obtained [167]. Occurrence [307, 286, 345]. Ocean [212, 367, 222]. Oceanography [251]. Oceans [316]. October [257]. Octopus [331, 70]. odoe [137]. off [13, 139]. Offshore [348]. Ogun [39]. Oguta [189]. Oil [13, 194, 213, 285]. Okavango [58]. oleifera [206]. OLS [230]. Oman [13, 11]. Omega [213]. Omega-3 [213]. on-growth [119]. Oncorhynchus [135, 359]. Ondo [52]. one [203]. Online [50]. Only [12]. Ontogenesis [186]. Operated [363]. Opportunities [65]. Optimization [220]. Optimum [230]. Orconectes [88, 55]. Oreochoromis [192]. Oreochromis [368, 240, 156, 130, 210, 185, 53, 25, 230, 74, 260, 312, 198, 120]. Organic [366]. Organs [354, 166]. Ornamental [8, 80, 102]. Oromia [33, 143, 41, 322]. Osmolality [28]. Otamiri [158]. Otoliths [76]. Outlining [348]. Ovarian [222]. Ovatide [72]. Overview [287, 340, 289, 64]. Ownership [216]. Oxidative [45]. Oxytetracycline [206]. Oyo [282, 323]. Oyster [290, 157].

P [152]. P. [303]. Pacific [139, 34]. Padma [75]. Pagasitikos [48]. Pakhtunkhwa [344]. Pakistan [344]. Pakistani [159]. Pangasius [177]. Parachanna [180]. Paracherodon [79]. Paradigm [58]. Parameter [234]. Parameters [39, 154, 2, 281, 158, 146, 206, 307, 107, 104]. Paramormyrops [237]. Parasite [79, 240, 108]. Parasites [351, 307, 237]. Parasitic [53, 134, 14, 343]. Parentage [364]. Park [212]. Part [133]. Participation [286]. Participatory [41]. Pathogen [364]. Pathogenesis [21]. Pathogenic [156, 172]. Pathogens [295, 198]. Pathological [118, 179]. Patterns [50, 137]. pausicostata [100]. Pectoral [24]. pelagicus [15]. Penaeid [222]. Penaeidae [94, 60]. Penaeus [222, 364, 104]. Pentahydrate [200]. Percentage [103]. Perch [152, 236]. Percocypris [200]. Percoidei [139]. Performance [227, 16, 219, 185, 25, 142, 120, 231]. Performances [150]. Period [79, 15]. Periods [358]. Perspective [45, 84]. perspectives [116]. Pesticides [211]. petherici [165]. pethericii [165]. Petivera [178]. Phagicola [77]. Phenetic [174]. Phenotypic [368, 234, 358]. Philippine [43, 86]. Phosphorus [29]. Physical [39, 33]. Physico [240, 107, 96]. Physico-Chemical [240, 96, 107]. Physicochemical [192, 229, 158]. Physiologic [186]. Physiological [28]. Physiology [16]. Phytochemical [271, 313]. Phytoplankton [33, 104, 96]. Pigmentation [154]. pingi [200]. Pisces [358, 260, 312]. Plankton [300, 321, 5]. Plant [150, 87, 100, 236, 301]. Plant-based [236]. Plants [132]. Plasma [28, 7]. Plasticity [358]. platanus [77]. platensis [220]. Platform [194]. PLS [19]. Point [37]. Points [223]. Policy [147, 196]. Pollutants [221]. Pollution [111, 53, 115, 84, 347]. Polyculture [71]. Polymorphic [207]. Polysaccharides [93]. Pomfret [338]. Pond [211, 71, 216, 278]. Ponds [300, 321, 361, 142, 109, 347, 29]. Population [110, 193, 4, 207, 168, 226, 11, 270, 95, 34, 268, 160]. Populations [61]. Portunidae [15]. Portunus [61, 15]. Pose [43]. Possible [62]. Post [41, 19, 112, 78, 153]. Post-harvesting [112]. Potassium [12, 283]. Potential [199, 119, 301]. Potentials [22, 175]. Practical [32]. Practices [345]. Pradesh [303]. Pre [59]. Pre-feasibility [59]. Precision [24]. Predation [95]. Predator [339]. Predatory [330, 178]. Preferences [50]. Preliminary [198, 203]. Preparation [278]. Present [116, 254, 103]. Prestigious [251]. Prevalence [307, 343]. Prevention [35, 294]. Prey [70]. Price [147]. Prioritization [98]. Probiotic [9, 25, 248]. Probiotic-Based [9]. Probiotics [3, 73, 296, 306, 324, 303]. Problems [112]. Procedure [105, 230]. Process [360]. Processed [67]. Processing [365, 144, 153]. Product [153]. Production [147, 299, 112, 360, 140, 113, 155, 216, 124, 109, 65, 302, 74, 333]. Productions [87]. Products [35, 287, 213, 87, 50, 282, 323]. Profitability [52]. Promoting [333]. Propagation [80]. Properties [192, 229, 215]. Property [44]. propinquus [55]. Prospective [6]. Prospects [197, 92]. Protected [367]. Protecting [316]. Protection [305]. Protective [220]. Protein [71, 1, 146, 17, 91, 152, 100]. Proteins [67, 215, 301]. Proximate [185, 129]. Pseudotolithus [160]. Pterophyllum [358]. punctatus [311]. Purification [124]. Purworejo [217]. Put [37]. Put-and-Take [37]. putitora [118].

Qarun [225]. Quality [211, 281, 224, 228, 25, 307, 5, 153, 303]. Quantitative [79]. Quantum [154].

Rainbow [37, 121, 26, 163]. Rainy [229]. Rajshahi [75]. ramada [223]. Random [207]. RAPD [207]. Rate [10, 15]. rather [76]. Ratio [152]. Ration [16]. Ratios [7]. Rays [24]. Razole [303]. Re [8, 67]. Re-Authorization [67]. Re-Circulating [8]. Rearing [70]. Reason [113]. Reba [183, 182, 128, 183, 182]. Recirculating [274, 92]. Recombinant [124]. Recreational [264]. Recruitment [149]. Red [325, 42, 120, 190]. Redbelly [242]. Reduce [135]. Reference [223]. Regency [217]. Regimen [284]. Region [59, 33, 240, 102]. Regions [126]. Related [112, 84]. Relation [167, 33, 14, 15]. Relationship [88, 174, 14, 188]. Relationships [75, 180, 55]. Relative [201]. Reproduction [328, 139, 9]. Reproductive [13, 128, 10, 331, 182]. rerio [171, 155, 138]. Research [47, 370, 172, 77, 51]. Researchers [251]. Reservoir [180, 182, 183, 161]. resistance [303]. Resource [199, 49, 342]. Resources [253, 250, 254]. Response [169, 42]. Retention [121]. Return [131]. Reuse [132]. Revealed [125]. Review [308, 232, 279, 212, 81, 62, 94, 128, 73, 134, 6, 99, 218]. Rhabdosargus [11]. Rheological [215]. Rhythm [162]. Ricinus [227]. Rift [368]. Right [76]. Ring [174]. Rio [102]. Rising [288, 314]. Risk [305, 179, 252]. Risks [21]. River [39, 175, 167, 165, 123, 281, 305, 114, 164, 229, 190, 158, 284, 188, 196, 138, 181, 214, 75, 196]. Rivers [201, 141]. RNA [152]. RNA/DNA [152]. Roach [7]. rohita [169, 17, 181]. Rohu [169]. Role [209, 350, 316, 49, 294, 173]. root [178]. root-extract [178]. Rule [298]. Ruminant [67]. russelli [159]. rusticus [88]. Rusty [88]. rutilus [7, 7].

S. [228]. Safety [35]. salar [359]. Saline [25, 29]. Salinity [16, 12, 15]. Salmo [224, 359, 284]. salmoides [242]. Salmon [232, 212, 213, 36, 40, 163, 255, 131]. Salmonids [359]. Sand [217]. Sander [14]. Sange [165]. sarba [11]. Sardine [13, 34]. Sardinella [13, 148]. Sardinops [34]. sargus [110]. Save [258]. Scad [159]. scalare [358]. Scale [125]. Scales [24, 334]. schedule [257]. Schizothoracine [113]. Schooling [337]. Scientific [57]. Scomberomorus [191]. Screening [368]. Sea [56, 232, 191, 273, 11]. Seabream [110, 11]. Seafood [50]. Seahorse [328]. Season [229, 261, 131]. Seasonal [161, 38, 103]. Seasonally [225]. Seasons [5]. Seatrout [4]. Seawater [2]. Sector [64]. Sediment [117]. Sediments [177, 192]. Seeb [13]. Seed [227, 59, 72, 269]. Selected [41, 177, 322, 228, 202, 307, 144, 52, 343, 153]. Selectivity [164, 205]. Self [170]. Self-Sustaining [170]. Semen [28]. Seminal [28, 7]. Senegal [194]. senegalensis [160]. Set [191]. Set-Net [191]. Several [186]. Sex [74]. Sexual [88]. Shark [103]. Sharp [288, 314]. shasta [36, 40]. Shell [208, 174, 109]. Shewa [141]. SHG [302]. Shift [58, 125]. shifts [260, 312]. Shoals [319]. Shock [155]. Shore [288, 314]. Short [219, 94]. Short-Term [219]. Shrimp [19, 222, 254, 292, 60, 233, 217, 29, 105]. Shrimps [94, 104, 21]. Siani [110]. Side [76]. Significance [337, 329]. Silver [72]. Similar [96]. Sinai [223]. Sindh [181]. sinensis [179]. Single [91]. Sites [41, 168, 107, 119]. Size [23, 164]. Skirring [194]. skunk [178]. Small [37, 23, 307, 125]. Smoking [213]. Smoothing [131]. Snails [111]. Snapper [325]. Social [102]. Socio [241]. Socio-Economic [241]. Sockeye [255]. Sodium [32]. Somatolactin [124]. Some [195, 201, 177, 33, 228, 242, 223, 14, 120, 108]. Sorghum [120]. Sources [163]. South [5, 105, 365, 60, 30, 343]. South-Eastern [5]. Southeast [205, 54]. Southeastern [305]. Southern [143]. Southwest [137]. Southwestern [180]. Soybean [249]. Sp [195, 108]. Spanish [191]. Sparidae [11]. Spatial [165, 96]. Spawners [255]. Spawning [7, 105]. Species [201, 167, 127, 33, 228, 300, 321, 130, 226, 5, 95, 326, 74, 198, 31, 125, 237]. Species-Based [125]. Specific [364]. Specific-Pathogen-Free [364]. Specifications [363]. Specificity [127]. specularis [146]. Sperm [311, 28, 7, 162]. Spermatogenia [247]. Sphyraena [205]. Spines [24]. Spirulina [220]. Spline [131]. Spoilage [83]. Sport [264]. Spot [21]. Spotted [4]. Spring [42]. Sprout [150]. squamata [129]. Squids [330]. Stage [88]. Stand [51]. Stand-Alone [51]. Starch [120]. State [39, 167, 229, 228, 158, 282, 323, 180, 144, 52]. Station [130]. Status [369, 197, 322, 151, 254, 226, 103, 126, 181, 241, 116]. Stem [206]. Stenella [221]. Step [230]. Stock [110, 364, 159, 34]. Stocking [78, 217]. Stomach [221, 136]. Storage [79, 215]. Stored [42]. Strains [368, 185]. Strange [288, 314]. Strategic [212, 98]. Strategies [250, 233, 252]. Stream [255, 53]. Stress [45]. Stretches [363]. Structure [305, 115, 341]. Studies [16, 229, 186, 204]. Study [369, 176, 218, 83, 59, 212, 33, 50, 174, 157, 117, 136, 317, 203, 343]. Stunning [224]. Sturgeons [186]. Styrling [156]. Subcontinent [128]. Substitutes [87]. Sudan [161, 365]. suecica [63]. Suitability [55]. Sulfate [200]. Summit [372]. Supplement [22, 20, 100]. Supplementation [3]. Supplements [236]. Surface [108, 191]. Surrounding [327]. Survey [194, 50, 255, 104, 60, 198]. Surveys [37]. Survival [18, 78, 15, 70].

Sustainability [66, 340]. Sustainable [293, 316, 97, 342, 333]. Sustained [157]. Sustaining [170]. Swabi [344]. Swim [336]. Swimming [61, 359]. Syndrome [247, 109, 21]. Synergistic [154]. Synodontis [204]. Synopsis [291]. Synthesis [152]. Synthetic [72]. System [211, 8, 130, 331, 140, 286, 51]. Systems [309, 190, 342, 283, 274, 92].

T [63, 135]. T-Iso [63]. Taiwan [212, 231]. Take [37]. Tambaqui [184]. Tamil [183, 182]. Tana [133]. Tanks [25]. tarda [143]. Targeted [223]. Technical [357]. Techniques [224, 364, 361, 289, 320, 310, 153]. Technological [279]. Technologies [41, 144]. Technology [176, 371, 248, 301]. Teeth [288, 314]. Teleost [355]. Temperature [191, 15]. Temperatures [1]. Temporal [33, 96]. Tenualosa [126]. Term [219]. Testes [145]. Testicular [247]. testudineus [90, 152, 136]. Tetra [79]. tetraodoni [195]. Tetraselmis [63]. Texas [4]. Textural [215]. Their [28, 167, 26, 62, 160, 327, 154, 114, 97, 248, 329]. Theory [58]. Therapeutic [279]. There [127, 58]. Thin [223]. Thoothukudi [205]. Thorns [288, 314]. Threat [43]. Threatened [128]. Threats [196, 326]. Three [23, 368, 300, 321, 185, 96, 181, 261, 166, 203, 237]. Three-in-one [203]. Threshold [101]. Thrips [34]. Thysanoptera [34]. Tigernut [20]. Tigers [104]. Tilapia [209, 167, 192, 219, 368, 156, 210, 242, 185, 25, 230, 124, 260, 312, 239, 120, 228, 225]. tilapiae [240]. Time [230]. Tissues [177, 208]. Tonekabon [284]. Tool [48, 6]. Tools [116]. Topography [108]. Tor [118]. Towkak [114]. Town [153]. Toxicity [200]. Toxicology [355]. Toxins [350]. Traditional [49]. Tragedy [69]. Traits [339, 234]. Transcriptome [145]. Transgenic [99]. Transmission [21]. Transportation [269]. Transversal [130]. Treat [132]. Treated [77, 100]. Treatment [135]. Trends [83, 197, 50]. Tributaries [281]. trituberculatus [61]. Trophic [309, 10, 342, 125]. Tropical [265]. Trout [37, 121, 26, 224, 163, 284]. trutta [224, 284]. Tuna [349, 340]. Tunisia [168]. Turag [123]. Turtle [2]. Turtles [56]. Twelve [75]. Two [167, 37, 103, 230]. Type [245, 186]. Types [295, 334]. Typha [31]. typus [160]. tyrannus [145].

Uganda [185, 173]. Ulva [129]. Un-ionized [171]. Understanding [173]. University [300, 321]. Unlocking [301]. Upazila [176, 241]. Update [151]. Uptake [175]. Use [87, 248, 138, 119]. Used [71, 279, 297, 289]. Uses [281, 306, 324]. Using [207, 255, 61, 72, 31, 131, 301, 252]. Utilisation [1]. Utilization [20, 219, 146, 142, 144, 342, 217, 100].

Vaccines [294]. Valenciennes [193]. Valley [106, 368, 113]. Value [66, 365, 282, 323]. vannamei [170, 249, 364, 105, 303, 109, 217]. Var [156, 146, 113]. Varapuzha [369]. Variability [61, 96]. Variance [364]. Variation [161, 167, 63]. varicolor [178]. Various [177, 192]. Varying [227]. Vegetables [140]. Vembanad [369, 363]. Ventures [52]. Verification [41]. VHSV [135]. Vibrio [218]. Victoria [204]. victoriae [204]. Vienna [257].

Vietnam [190]. Viral [135, 356]. Virus [21]. Viruses [94]. Viscera [142]. Viscera-based-diet [142]. Vision [57]. Visual [255]. Vitamin [163]. vulgaris [90].

Waste [22, 213, 132, 334]. Water [211, 212, 281, 192, 135, 240, 1, 322, 228, 98, 115, 25, 12, 307, 5, 104, 132, 29, 51, 41, 303]. Waterfalls [10]. Waters [13, 4, 222, 159, 205, 348, 221, 160, 244, 148]. Weather [245]. weed [178]. Weight [167, 75, 188, 88, 180]. Welfare [283]. Well [12, 180]. Well-being [180]. West [56]. Western [133, 307]. Wet [42]. Wetland [31]. Wetlands [369, 363]. White [110, 21, 217, 161]. Wild [46, 207, 168, 61, 157, 270, 186, 138]. will [258]. Winter [42]. within [160]. Wollo [343]. Women [82, 299, 144]. Wonchit [141]. World [333].

Yamuna [214]. Yaounde [130]. Yeast [231]. Yellow [236, 234]. Yield [103]. Yilan [212]. Young [251].

Zander [14]. Zebrafish [171, 149, 155, 138]. Zeway [143]. zilli [228, 242]. zillii [225]. Zimbabwe [49]. Zinc [200]. Ziway [260, 312]. ZnSO [200]. Zone [141, 343]. Zones [103].

References

Glencross:2010:EHW

[1] Brett Glencross and Michel Bermudes. Effect of high water temperatures on the utilisation efficiencies of energy and protein by juvenile barramundi, Lates calcarifer. Fisheries and Aquaculture Journal, 1(1):1–11, ???? 2010. CODEN FAJIAO. ISSN 2150-

     3508. URL https://www.longdom.org/open-access/effect-of-high-
     water-temperatures-on-the-utilisation-efficiencies-of-energy-
     2150-3508.1000014.pdf.

Chuen-Im:2010:EEP

[2] Thanaporn Chuen-Im, Phanthitra Phengpan, and Kamolchanok Panishkan. Effects of environmental parameters on bacterial levels in seawater from juvenile green turtle (Chelonia mydas) kept in captivity. Fisheries and Aquaculture Journal, 1(1):1–8, ???? 2010. CODEN FAJIAO. ISSN

     2150-3508. URL https://www.longdom.org/open-access/effects-
     of-environmental-parameters-on-bacterial-levels-in-seawater-
     2150-3508.1000009.pdf.

Harikrishnan:2010:SDC

[3] Ramasamy Harikrishnan, Chellam Balasundaram, and MoonSoo Heo. Supplementation diet containing probiotics, herbal and azadirachtin on hematological and biochemical changes in Cirrhina mrigala against Aphanomyces invadans. Fisheries and Aquaculture Journal, 1(1):1– 11, ???? 2010. CODEN FAJIAO. ISSN 2150-3508. URL https://

www.longdom.org/open-access/supplementation-diet-containing-
probiotics-herbal-and-azadirachtin-on-hematological-2150-3508.
1000004.pdf.

Anderson:2010:PGD

[4] Joel D. Anderson and William J. Karel. Population genetics and dynamics of spotted seatrout in the estuarine waters of Texas. Fisheries and Aquaculture Journal, 1(1):1–19, ???? 2010. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/population-genetics-
and-dynamics-of-spotted-seatrout-in-the-estuarine-waters-of-
texas-2150-3508.1000002.pdf.

Offem:2011:ISW

[5] B. O. Offem, E. O. Ayotunde, G. U. Ikpi, S. N. Ochang, and F. B. Ada. Influence of seasons on water quality, abundance of fish and plankton species of Ikwori Lake, South-Eastern Nigeria. Fisheries and Aquaculture Journal, 2(1):1–18, ???? 2011. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
influence-of-seasons-on-water-quality-abundance-of-fish-and-
plankton-species-of-ikwori-lake-southeastern-nigeria-43050.html.

Rather:2011:NNT

[6] M. A. Rather, R. Sharma, M. Aklakur, S. Ahmad, N. Kumar, M. Khan, and V. L. Ramya. Nanotechnology: a novel tool for aquaculture and fisheries development. A prospective mini-review. Fisheries and Aquaculture Journal, 2(1):1–5, ???? 2011. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/nanotechnology-
a-novel-tool-for-aquaculture-and-fisheries-development-2150-
3508.1000016.pdf.

Golpour:2011:CIR

[7] A. Golpour, M. R. Imanpoor, and S. A. Hosseini. Changes in ionic ratios of seminal plasma and its effect on sperm characteristics in Caspian roach (Rutilus rutilus caspicus) during spawning migration. Fisheries and Aquaculture Journal, 2(1):1–6, ???? 2011. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
changes-in-ionic-ratios-of-seminal-plasma-and-its-effect-on-
sperm-characteristics-2150-3508.1000017.pdf.

Fadhil:2011:EFM

[8] R. Fadhil, J. Endan, S. N. B. S. Arziman, F. S. Taip, and M. S. B. H. Ja’afar. Effectiveness of floating micro-bead bio-filter for ornamental fish in a re-circulating aquaculture system. Fisheries and Aquaculture Journal, 2(1):1–8, ???? 2011. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/effectiveness-
of-floating-microbead-biofilter-for-ornamental-fish-2150-3508.
1000028.pdf.

Lombardo:2011:PBN

[9] F. Lombardo, G. Gioacchini, and O. Carnevali. Probiotic-based nutritional effects on killifish reproduction. Fisheries and Aquaculture Journal, 2(1):1–11, ???? 2011. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/probioticbased-
nutritional-effects-on-killifish-reproduction-2150-3508.1000033.
pdf.

Ikpi:2012:CRD

[10] G. U. Ikpi, A. Jenyo-Oni, and B. O. Offem. Catch rate, distribution, trophic and reproductive biology of the African carp Labeo coubie in the Agbokim Waterfalls, Nigeria. Fisheries and Aquaculture Journal, 3(1):2–14, ???? 2012. CODEN FAJIAO.

ISSN 2150-3508.  URL https://www.longdom.org/open-access/
catch-rate-distribution-trophic-and-reproductive-biology-of-
the-african-carp-2150-3508.1000038.pdf.

Mehanna:2012:PDM

[11] S. F. Mehanna, F. R. Al-Kiyumi, and L. Al-Kharusi. Population dynamics and management of goldlined seabream Rhabdosargus sarba (Sparidae) from the Oman Coast of Arabian Sea. Fisheries and Aquaculture Journal, 3(1):2–10, ???? 2012. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/population-dynamics-
and-management-of-goldlined-seabream-2150-3508.1000040.pdf.

Mourad:2012:AMF

[12] N. Mourad, S. Kreydiyyeh, J. Ghanawi, and I. P. Saoud. Aquaculture of marine fish in inland low salinity well water: Potassium is not the only limiting element. Fisheries and Aquaculture Journal, 3(1): 2–13, ???? 2012. CODEN FAJIAO. ISSN 2150-3508. URL https:

//www.longdom.org/open-access/aquaculture-of-marine-fish-in-
inland-low-salinity-well-water-2150-3508.1000042.pdf.

Al-Jufaili:2012:RBI

[13] S. M. Al-Jufaili. Reproductive biology of the Indian oil sardine Sardinella longiceps from Al-Seeb waters off Oman. Fisheries and Aquaculture Journal, 3(1):2–13, ???? 2012. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/reproductive-
biology-of-the-indian-oil-sardine-sardinella-longiceps-2150-
3508.1000044.pdf.

Movahed:2012:SHC

[14] R. Movahed, H. Khara, M. R. Hayatbakhsh, and M. Rahbar. Some haematological changes of zander (Sander lucioperca) in relation to age and its relationship with parasitic infection. Fisheries and Aquaculture Journal, 3(1):2–8, ???? 2012. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/some-haematological-
changes-of-zander-sander-lucioperca-2150-3508.1000047.pdf.

Ravi:2012:SRD

[15] R. Ravi and M. K. Manisseri. Survival rate and development period of the larvae of Portunus pelagicus (Decapoda, Brachyura, Portunidae) in relation to temperature and salinity. Fisheries and Aquaculture Journal, 3(1):2–9, ???? 2012. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/survival-rate-and-
development-period-of-the-larvae-of-portunus-pelagicus-2150-
3508.1000049.pdf.

Barman:2012:EDS

[16] U. K. Barman, S. K. Garg, and A. Bhatnagar. Effect of different salinity and ration levels on growth performance and nutritive physiology of milkfish, Chanos chanos (Forsskål) — field and laboratory studies. Fisheries and Aquaculture Journal, 3(1):2–12, ???? 2012. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/effect-
of-different-salinity-and-ration-levels-on-growth-2150-3508.
1000053.pdf.

Manivannan:2012:IFP

[17] S. Manivannan and T. S. Saravanan. Impact of formulated protein diets on growth of the Indian major carp, Labeo rohita (Hamilton). Fisheries and Aquaculture Journal, 3(1):2–7, ???? 2012. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/impact-of-
formulated-protein-diets-on-growth-of-the-indian-major-2150-
3508.1000057.pdf.

Crooks:2013:ILD

[18] Neil Crooks, Wayne Rees, Alan Black, David Hide, John R. Britton, and Alan Henshaw. Influence of live and dry diets on growth and survival of chub (Leuciscus cephalus) larvae. Fisheries and Aquaculture Journal, 4 (1):1–6, ???? 2013. CODEN FAJIAO. ISSN 2150-3508. URL https://

www.longdom.org/open-access/influence-of-live-and-dry-diets-
on-growth-and-survival-2150-3508.1000062.pdf.

Darpaah:2013:EBF

[19] George A. Darpaah. Ecological and biodiversity friendly approach to post larval shrimp (PLS) collection in Ghana. Fisheries and Aquaculture Journal, 4(1):1–11, ???? 2013. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/ecological-and-
biodiversity-friendly-approach-2150-3508.1000063.pdf.

Agbabiaka:2013:UTC

[20] Lukman A. Agbabiaka and Chukwuka F. Ezeafulukwe. Utilization of tigernut (Cyperus esculentus L.) meal as dietary supplement by African catfish (Clarias gariepinus Burchell, 1822). Fisheries and Aquaculture Journal, 4(1):1–6, ???? 2013. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/utilization-of-
tigernut-cyperus-esculentus-l-2150-3508.1000064.pdf.

Reddy:2013:MPD

[21] Alla D. Reddy, Geevaretnam Jeyasekaran, and Robinson J. Shakila. Morphogenesis, pathogenesis, detection and transmission risks of white spot syndrome virus in shrimps. Fisheries and Aquaculture Journal, 4(1):1–13, ???? 2013. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/morphogenesis-
pathogenesis-detection-and-transmission-risks-of-2150-3508.1000066.
pdf.

Agbabiaka:2013:PKW

[22] Lukman A. Agbabiaka and Chinwe U. Madubuko. Potentials of kunnu waste as dietary supplement for African catfish fingerlings. Fisheries and Aquaculture Journal, 4(1):1–5, ???? 2013. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/potentials-of-
kunnu-waste-as-dietary-supplement-2150-3508.1000068.pdf.

Belford:2013:MBS

[23] Stanton G. Belford, Nanette E. Chadwick, and Maroof A. Khalaf. Measuring body size in small marine fishes: a comparison of three non-intrusive methods. Fisheries and Aquaculture Journal, 4(1):1–8, ???? 2013. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/measuring-body-size-
in-small-marine-fishes-a-comparison-2150-3508.1000071.pdf.

Morehouse:2013:ELB

[24] Reid L. Morehouse, Steven B. Donabauer, and Angela C. Grier. Estimating largemouth bass age: Precision and comparisons among scales, pectoral fin rays, and dorsal fin spines as nonlethal methods. Fisheries and Aquaculture Journal, 4(1):1–7, ???? 2013. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/estimating-
largemouth-bass-age-precision-and-comparisons-among-2150-3508.
1000074.pdf.

Mohamed:2013:APA

[25] Asaad H. Mohamed, Rex F. M. Traifalgar, and Augusto E. Serrano, Jr. Assessment of probiotic application on natural food, water quality and growth performance of saline tilapia Oreochromis mossambicus L. cultured in concrete tanks. Fisheries and Aquaculture Journal, 4(1):1–7, ???? 2013. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/assessment-of-
probiotic-application-on-natural-food-2150-3508.1000075.pdf.

Courtney:2013:RTT

[26] Joshua M. Courtney, Amy C. Courtney, and Michael W. Courtney. Rainbow trout and their hybrids out compete cutthroat trout in a lentic ecosystem? Fisheries and Aquaculture Journal, 4(1):1–7, ???? 2013. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/do-rainbow-trout-and-their-hybrids-outcompete-cutthroat-
trout-2150-3508.1000078.pdf.

Ganguly:2013:MDF

[27] Subha Ganguly. Microfloral diversity in fish: an editorial. Fisheries and Aquaculture Journal, 4(1):1, ???? 2013. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/microfloral-
diversity-in-fish-an-editorial-2150-3508.1000e101.pdf.

Aramli:2013:SBH

[28] Mohammad Sadegh Aramli, Rajab Mohammad Nazari, Mohammad Reza Kalbassi, and Salim Aramli. Semen of beluga, Huso huso: Ionic content and osmolality of seminal plasma and their physiological correlation with sperm motility indices. Fisheries and Aquaculture Journal, 4(1):

1–5, ???? 2013. CODEN FAJIAO. ISSN 2150-3508. URL https://

www.longdom.org/open-access/semen-of-beluga-huso-huso-ionic-
content-2150-3508.1000079.pdf.

Sun:2013:PNB

[29] Wei Sun and Claude E. Boyd. Phosphorus and nitrogen budgets for inland, saline water shrimp ponds in Alabama. Fisheries and Aquaculture Journal, 4(1):1–5, ???? 2013. CODEN FAJIAO. ISSN 2150-3508. URL https://

www.longdom.org/open-access/phosphorus-and-nitrogen-budgets-
for-inland-saline-water-shrimp-ponds-in-alabama-2150-3508.1000080.
pdf.

Selvam:2013:DAF

[30] J. Selvam, D. Varadharajan, A. Babu, and T. Balasubramanian. Distribution and abundance of finfish eggs from Muthupettai, South East Coast of India. Fisheries and Aquaculture Journal, 4(1):1–19, ???? 2013. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/distribution-and-abundance-of-finfish-eggs-from-
muthupettai-south-east-coast-of-india.pdf.

Tessema:2013:MMU

[31] Assefa Tessema, Nuria Abdurohman, and K. S. Goudar. Mattress making using Typha latifolia and Cyprus species of Chefa Wetland in Kemissie, Ethiopia: a means for livelihood improvement. Fisheries and Aquaculture Journal, 4(1):1–5, ???? 2013. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
mattress-making-using-typha-latifolia-and-cyprus-species-of-
chefa-wetland-in-kemissie-ethiopia-a-means-for-livelihood-improvement-
2150-3508.1000082.pdf.

Adeniji:2013:MGC

[32] Comfort Adetutu Adeniji, Pius Abimbola Okiki, Ajani Murano Rasheed, and Rasheed Bolaji. Mouldy groundnut cake and hydrated sodium calcium aluminosilicate in practical diet for African catfish Clarias gariepinus (Burchell, 1822). Fisheries and Aquaculture Journal, 4(1):1–5, ???? 2013. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/mouldy-groundnut-cake-and-hydrated-sodium-calcium-
aluminosilicate-2150-3508.1000083.pdf.

Fayissa:2013:STD

[33] Zelalem Dessalegn Fayissa and Demeke Kifle. The study of temporal dynamics of phytoplankton biomass and species composition in relation to some physical and chemical factors, Lake Kuriftu, Oromia Region, Ethiopia. Fisheries and Aquaculture Journal, 4(1):1–7, ???? 2013. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/the-study-of-temporal-dynamics-of-phytoplankton-
biomass-and-species-2150-3508.1000084.pdf.

Sakuramoto:2013:CCP

[34] Kazumi Sakuramoto. A common concept of population dynamics applicable to both Thrips imaginis (Thysanoptera) and the Pacific stock of the Japanese sardine (Sardinops melanostictus). Fisheries and Aquaculture Journal, 4(1):1–9, ???? 2013. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/a-common-concept-of-
population-dynamics-applicable-to-both-thrips-2150-3508.1000085.
pdf.

Andrada:2013:FSP

[35] Mitchel Andrada. Food safety prevention on biohazards in fish and fishery products. Fisheries and Aquaculture Journal, 4(1):1, ???? 2013. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/food-safety-prevention-on-biohazards-in-fish-and-fishery-
products-2150-3508.1000e102.pdf.

Fujiwara:2014:DIJa

[36] Masami Fujiwara. Dynamics of infection of juvenile Chinook salmon with Ceratomyxa shasta. Fisheries and Aquaculture Journal, 5(1):1–6, ???? 2014. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/dynamics-of-infection-of-juvenile-chinook-salmon-
with-ceratomyxa.pdf.

Barnes:2014:CCC

[37] Michael E. Barnes, Greg Simpson, John Carreiro, and Jill Voorhees. A comparison of a creel census to modeled access-point creel surveys on two small lakes managed as put-and-take rainbow trout fisheries. Fisheries and Aquaculture Journal, 5(1):1–4, ???? 2014. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
a-comparison-of-a-creel-census-to-modeled-accesspoint-creel-
surveys-on-two-small-lakes-managed-as-putandtake-rainbow-trout-
fisheries-2150-3508.1000086.pdf.

Emam:2014:SHC

[38] Mahmoud Abdelghagffar Emam and Badia Abughrien. Seasonal histological changes in gonads of the catfish (Clarias lazera). Fisheries and Aquaculture Journal, 5(1):1–4, ???? 2014. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/seasonal-
histological-changes-in-gonads-of-the-catfish-clarias-lazera-
2150-3508.1000087.pdf.

Adeosun:2014:PCP

[39] F. I. Adeosun, B. J. Akin-Obasola, F. B. Oyekanmi, and J. O. Kayode. Physical and chemical parameters of Lower Ogun River Akomoje, Ogun State, Nigeria. Fisheries and Aquaculture Journal, 5(1):1–5, ???? 2014. CODEN FAJIAO. ISSN 2150-3508. URL https://www.

longdom.org/open-access/physical-and-chemical-parameters-of-
lower-ogun-river-akomoje-ogun-state-nigeria-2150-3508.1000088.
pdf.

Fujiwara:2014:DIJb

[40] Masami Fujiwara. Dynamics of infection of juvenile Chinook salmon with Ceratomyxa shasta. Fisheries and Aquaculture Journal, 5(1):0, ???? 2014. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/dynamics-of-infection-of-juvenile-chinook-salmon-
with-ceratomyxa-2150-3508.1000089.pdf.

Abdi:2014:PEV

[41] Tilahun Genet Abdi. Participatory evaluation and verification of improved post harvest fishery technologies on selected sites of Oromia water bodies. Fisheries and Aquaculture Journal, 5(1):2–7, ???? 2014. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
participatory-evaluation-and-verification-of-improved-post-harvest-
fishery-technologies-on-selected-sites-of-oromia-water-bodies-
2150-3508.1000090.pdf.

Wyatt:2014:CIN

[42] Jessica Wyatt, Sharon Kenny, Terry Mills, Dawn H. Marshall, and Harry M. Murray. Condition index and neutral red assay response of cultured Mytilus edulis L. stored in a wet holding facility during winter and spring in Northeastern Newfoundland. Fisheries and Aquaculture Journal, 5(1):2–8, ???? 2014. CODEN FAJIAO. ISSN 2150-3508. URL https://

www.longdom.org/open-access/condition-index-and-neutral-red-
assay-response-of-cultured-mytilus-edulis-l-stored-in-a-wet-
holding-facility-during-winter-and-spring-in-northeastern-newfoundland-
2150-3508.1000091.pdf.

Andrada:2014:CCP

[43] Mitchel Andrada. Climatic conditions pose a threat to Philippine fishery industry. Fisheries and Aquaculture Journal, 5(1):2, ???? 2014.

CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/climatic-conditions-pose-a-threat-to-philippine-
fishery-industry-2150-3508.1000e103.pdf.

Koonjul:2014:KMI

[44] Priyum Koonjul. Knowledge management and intellectual property issues in aquaculture. Fisheries and Aquaculture Journal, 5(1):2, ???? 2014. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/knowledge-management-and-intellectual-property-issues-
in-aquaculture-2150-3508.1000e104.pdf.

Blier:2014:FHO

[45] Pierre Blier. Fish health: an oxidative stress perspective. Fisheries and Aquaculture Journal, 5(1):2, ???? 2014. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/fish-health-an-
oxidative-stress-perspective-2150-3508.1000e105.pdf.

Alfaro:2014:DBA

[46] Andrea C. Alfaro. The dawn of biotechnological applications on wild harvests and cultivated crops. Fisheries and Aquaculture Journal, 5(1): 2–5, ???? 2014. CODEN FAJIAO. ISSN 2150-3508. URL https://www.

longdom.org/open-access/the-dawn-of-biotechnological-applications-
on-wild-harvests-and-cultivated-crops-2150-3508.1000e106.pdf.

Exadactylos:2014:NAR

[47] Athanasios Exadactylos. Nutrigenomics in aquaculture research. Fisheries and Aquaculture Journal, 5(2):1, ???? 2014. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/nutrigenomics-
in-aquaculture-research-2150-3508.10000e107.pdf.

Petihakis:2014:TEB

[48] G. Petihakis, A. Theodorou, K. Tsiaras, A. Pollani, A. Prospathopoulos, and G. Triantafyllou. A tool for ecosystem-based management applied to Pagasitikos Gulf. Fisheries and Aquaculture Journal, 5(2):1–4, ???? 2014. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/a-tool-for-ecosystembased-management-applied-to-
pagasitikos-gulf-2150-3508.1000092.pdf.

Mawere:2014:TAC

[49] Munyaradzi Mawere, Christopher M. Mabeza, and Thokozani Shava. Traditional authority in Community Based Natural Resource Management (CBNRM): a critical investigation of the role of traditional authority in fisheries management in Gache Gache Communal Lands of Kariba, Zimbabwe. Fisheries and Aquaculture Journal, 5(2):1–7, ???? 2014. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/traditional-
authority-in-community-based-natural-resource-management-cbnrm-
a-critical-investigation-of-the-role-of-traditional-authority-
in-fisheries-2150-3508.1000093.pdf.

Hu:2014:OSS

[50] Yaqin Hu, Chunhong Yuan, Kefeng Yu, Yinghong Qu, Shunsheng Chen, Xichang Wang, and Ikuo Kimura. An online survey study of consumer preferences on aquatic products in China: Current seafood consumption patterns and trends. Fisheries and Aquaculture Journal, 5(2):1–6, ???? 2014. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/an-online-
survey-study-of-consumer-preferences-on-aquatic-products-in-
china-current-seafood-consumption-patterns-and-trends-2150-3508.
1000094.pdf.

Zadlock:2014:LCS

[51] Frank J. Zadlock IV, Jianjun Feng, Henri Estanbouli, Maria Bender, and Ziping Zhang. A low cost, stand-alone, cold water aquatic animal research housing system. Fisheries and Aquaculture Journal, 5(2):1–6, ???? 2014. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/a-low-cost-standalone-cold-water-aquatic-animal-
research-housing-system-2150-3508.1000095.pdf.

Thompson:2014:PSV

[52] O. A. Thompson and T. E. Mafimisebi. Profitability of selected ventures in catfish aquaculture in Ondo State, Nigeria. Fisheries and Aquaculture Journal, 5(2):1–7, ???? 2014. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/profitability-
of-selected-ventures-in-catfish-aquaculture-in-ondo-state-nigeria-
2150-3508.1000096.pdf.

Mahmoud:2014:IMM

[53] Nisreen E. Mahmoud, M. M. Fahmy, and Mohga F. M. Badawy. Investigations on mass mortalities among Oreochromis niloticus at Mariotteya Stream, Egypt: Parasitic infestation and environmental pollution impacts. Fisheries and Aquaculture Journal, 5(2):1–7, ???? 2014. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/investigations-on-mass-mortalities-among-oreochromis-
niloticus-at-mariotteya-stream-egypt-parasitic-infestation-and-
environmental-pollution-impacts-2150-3508.1000097.pdf.

Andrada:2014:SAA

[54] Mitchel Andrada. Southeast Asian aquaculture: How far are we? Fisheries and Aquaculture Journal, 5(2):1, ???? 2014. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/southeast-
asian-aquaculture-how-far-are-we-2150-3508.1000e108.pdf.

Simon:2014:HSI

[55] Thomas P. Simon and Nicholas J. Cooper. Habitat suitability index relationships for the northern clearwater crayfish, Orconectes propinquus (Decapoda: Cambaridae). Fisheries and Aquaculture Journal, 5(3):1–7, ???? 2014. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/habitat-suitability-index-relationships-for-
the-northern-clearwater-crayfish-orconectes-propinquus-decapoda-
cambaridae-2150-3508.1000098.pdf.

Ayissi:2014:IAC

[56] I. Ayissi and T. J. E. Jiofack. Impact assessment on by-catch artisanal fisheries: Sea turtles and mammals in Cameroon, West Africa. Fisheries and Aquaculture Journal, 5(3):1–5, ???? 2014. CODEN

FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/impact-assessment-on-bycatch-artisanal-fisheries-sea-
turtles-and-mammals-in-cameroon-west-africa-2150-3508.1000099.
pdf.

Yadav:2014:BNH

[57] R. C. Yadav. Biological nitrogen harvesting from aquatic ecosystems — a new scientific vision. Fisheries and Aquaculture Journal, 5(3):1–8, ???? 2014. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/biological-nitrogen-harvesting-from-aquatic-
ecosystemsa-new-scientific-vision-2150-3508.1000100.pdf.

K:2014:CFT

[58] Mosepele K. Classical fisheries theory and inland (floodplain) fisheries management; is there need for a paradigm shift? Lessons from the Okavango Delta, Botswana. Fisheries and Aquaculture Journal, 5(3): 1–8, ???? 2014. CODEN FAJIAO. ISSN 2150-3508. URL https:/

/www.longdom.org/open-access/classical-fisheries-theory-and-
inland-floodplain-fisheries-management-is-there-need-for-a-paradigm-
shift-lessons-from-the-okavango-delta-botswana-2150-3508.1000101.
pdf.

Carrasco:2014:PFS

[59] A. V. Carrasco, M. Astorga, A. Cisterna, A. Farías, V. Espinoza, and I. Uriarte. Pre-feasibility study for the installation of a Chilean mussel Mytilus chilensis (Hupe, 1854) seed hatchery in the Lakes Region, Chile. Fisheries and Aquaculture Journal, 5(3):1–5, ???? 2014. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
prefeasibility-study-for-the-installation-of-a-chilean-mussel-
mytilus-chilensis-hup-seed-hatchery-in-the-lakes-region-chiles-
2150-3508.1000102.pdf.

Rajakumaran:2014:SPS

[60] Perumal Rajakumaran and Baskralingam Vaseeharan. Survey on Penaeidae shrimp diversity and exploitation in South East Coast of India. Fisheries and Aquaculture Journal, 5(3):1–8, ???? 2014. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
survey-on-penaeidae-shrimp-diversity-and-exploitation-in-south-
east-coast-of-india-2150-3508.1000103.pdf.

Liu:2014:AGV

[61] Ling-Xiao Liu, Yun-Guo Liu, and Shi-Chao Xing. An analysis of genetic variability in wild and hatchery populations of swimming crab (Portunus trituberculatus) using AFLP markers. Fisheries and Aquaculture Journal, 5(3):1–5, ???? 2014. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/an-analysis-of-
genetic-variability-in-wild-and-hatchery-populations-of-swimming-
crab-portunus-trituberculatus-using-aflp-markers-2150-3508.1000104.
pdf.

Gallardi:2014:EBA

[62] Daria Gallardi. Effects of bivalve aquaculture on the environment and their possible mitigation: a review. Fisheries and Aquaculture Journal, 5 (3):1–8, ???? 2014. CODEN FAJIAO. ISSN 2150-3508. URL https:/

/www.longdom.org/open-access/effects-of-bivalve-aquaculture-
on-the-environment-and-their-possible-mitigation-a-review-2150-
3508.1000105.pdf.

Dorner:2014:VFA

[63] Jéssica Dörner, Pamela Carbonell, Soledad Pino, and Ana Farías. Variation of fatty acids in Isochrysis galbana (T-Iso) and Tetraselmis suecica, cultured under different nitrate availabilities. Fisheries and Aquaculture Journal, 5(3):1–3, ???? 2014. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/variation-of-fatty-
acids-in-isochrysis-galbana-tiso-and-tetraselmis-suecica-cultured-
under-different-nitrate-availabilities-2150-3508.1000106.pdf.

Ragusa:2014:OFS

[64] Gianluca Ragusa. Overview of the fisheries sector in The Gambia. Fisheries and Aquaculture Journal, 5(3):1–4, ???? 2014. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
overview-of-the-fisheries-sector-in-the-gambia-2150-3508.1000107.
pdf.

Riisgaard:2014:MMN

[65] Hans Ulrik Riisgård. ‘Mini-mussels’ — new opportunities and environmentally friendly production. Fisheries and Aquaculture Journal, 5 (3):1, ???? 2014. CODEN FAJIAO. ISSN 2150-3508. URL https:

//www.longdom.org/open-access/minimussels-new-opportunities-
and-environmentally-friendly-production-2150-3508.1000e109.pdf.

Andrada:2014:CSG

[66] Mitchel Andrada. Circles of sustainability: a good market value. Fisheries and Aquaculture Journal, 5(4):1, ???? 2014. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/circles-of-
sustainability-a-good-market-value-2150-3508.10000e110.pdf.

Karapanagiotidis:2014:RAN

[67] Ioannis T. Karapanagiotidis. The re-authorization of non-ruminant processed animal proteins in European aqua feeds. Fisheries and Aquaculture Journal, 5(4):1–3, ???? 2014. CODEN FAJIAO.

ISSN 2150-3508. URL https://www.longdom.org/open-access/the-
reauthorization-of-nonruminant-processed-animal-proteins-in-
european-aqua-feeds-2150-3508.10000e111.pdf.

Berg:2014:LF

[68] Ole Kristian Berg. Learning by failing? Fisheries and Aquaculture Journal, 5(4):1, ???? 2014. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/learning-by-failing-
2150-3508.10000e112.pdf.

Tsikliras:2014:FMM

[69] Athanassios C. Tsikliras. Fisheries mismanagement in the Mediterranean: a Greek tragedy. Fisheries and Aquaculture Journal, 5(4):

1, ???? 2014. CODEN FAJIAO. ISSN 2150-3508. URL https:/

/www.longdom.org/open-access/fisheries-mismanagement-in-the-
mediterranean-a-greek-tragedy-2150-3508.10000e113.pdf.

Rosas:2014:EDP

[70] Carlos Rosas, Maite Mascaró, Richard Mena, Claudia Caamal-Monsreal, and Pedro Domingues. Effects of different prey and rearing densities on growth and survival of Octopus maya hatchlings. Fisheries and Aquaculture Journal, 5(4):1–7, ???? 2014. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/effects-of-
different-prey-and-rearing-densities-on-growth-and-survival-
of-octopus-maya-hatchlings-2150-3508.10000108.pdf.

Asadujjaman:2014:DPL

[71] M. Asadujjaman, Shahangir Biswas, M. Manirujjaman, Matiar Rahman, M. A. Hossain, and M. A. Islam. Determination of protein, lipid and carbohydrate contents of conventional and non-conventional feed items used in carp polyculture pond. Fisheries and Aquaculture Journal, 5(4): 1–5, ???? 2014. CODEN FAJIAO. ISSN 2150-3508. URL https:/

/www.longdom.org/open-access/determination-of-protein-lipid-
and-carbohydrate-contents-of-conventional-and-nonconventional-
feed-items-used-in-carp-polyculture-pond-2150-3508.10000109.
pdf.

Rashid:2014:IBG

[72] Mudasir Rashid, Masood ul Hassan Balkhi, Gulzar Ah.Naiko, and Tamim Ahamad. Induced breeding of grass carp (Ctenopharyngodon idella) and silver carp (Hypophthalmichthys molitrix) using ovatide as synthetic hormone at National Fish Seed Farm (NFSF) Manasbal, Kashmir, J&K. Fisheries and Aquaculture Journal, 5(4):1–4, ???? 2014. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/induced-breeding-of-grass-carp-ctenopharyngodon-
idella-and-silver-carp-hypophthalmichthys-molitrix-using-ovatide-
2150-3508.10000110.pdf.

Michael:2014:RPA

[73] Ekundayo Taiye Michael, Sogbesan Olukayode Amos, and Lauratu Tahir Hussaini. A review on probiotics application in aquaculture. Fisheries and Aquaculture Journal, 5(4):1–3, ???? 2014. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/a-review-on-
probiotics-application-in-aquaculture-2150-3508.10000111.pdf.

Singh:2015:CBS

[74] Atul K. Singh and P. P. Srivastava. A CYP19 based sex determination and monosex production in aquaculture species Oreochromis niloticus L. and a cyprinid Cyprinus carpio L. Fisheries and Aquaculture Journal, 6(1):1–6, ???? 2015. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/a-cyp-based-sex-determination-and-monosex-production-
in-aquaculture-species-oreochromis-niloticus-l-and-a-cyprinid-
cyprinus-carpio-l2150-3508.10000112.pdf.

Mortuza:2015:LWR

[75] Golam Mortuza M and Fahad A. Al-Misned. Length–weight relationships of twelve fishes from the River Padma near Rajshahi City, Bangladesh. Fisheries and Aquaculture Journal, 6(1):1–2, ???? 2015. CODEN

FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/lengthweight-relationships-of-twelve-fishes-from-the-
river-padma-nearrajshahi-city-bangladesh-2150-3508-10000113.
pdf.

Gao:2015:IDB

[76] Yongwen Gao, Joseph Petersen, Robert Conrad, and David L. Dettman. Isotopic differences between the left and right side otoliths of flatfish indicating growth rather than environment. Fisheries and Aquaculture Journal, 6(1):1–5, ???? 2015. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/isotopic-differences-between-the-left-and-right-side-
otoliths-of-flatfishindicating-growth-rather-than-environment-
2150-3508-10000114.pdf.

Rodrigues:2015:RAP

[77] Marianna Vaz Rodrigues, Agar Costa Alexandrino de Pérez, Thaís Moron Machado, Fátima Maria Orisaka, Jacqueline Kazue Kurissio, and Andrea Lafisca. Research of Ascocotyle (Phagicola) longa in heat treated fillets of mullet (Mugil platanus). Fisheries and Aquaculture Journal, 6(1):1–3, ???? 2015. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/research-of-
ascocotyle-phagicola-longa-in-heat-treated-fillets-of-mulletmugil-
platanus-2150-3508-10000115.pdf.

Nwipie:2015:ISD

[78] G. N. Nwipie, E. S. Erondu, and N. Zabbey. Influence of stocking density on growth and survival of post fry of the African mud catfish, Clarias gariepinus. Fisheries and Aquaculture Journal, 6(1):1–4, ???? 2015. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/influence-of-stocking-density-on-growth-and-survival-
of-post-fry-of-theafrican-mud-catfish-clarias-gariepinus-2150-
3508-10000116.pdf.

Benzaken:2015:QAP

[79] Zehev Schwartz Benzaken, Clycia Pereira de Araújo Lima, and Vera Maria Fonseca de Almeida Val. Quantitative analyses of parasite load in the cardinal tetra (Paracherodon axelrodii), along the period of capture and storage. Fisheries and Aquaculture Journal, 6(1):1–3, ???? 2015. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/quantitative-
analyses-of-parasite-load-in-the-cardinal-tetra-paracherodonaxelrodii-
along-the-period-of-capture-and-storage-2150-3508-10000117.pdf.

Sarmah:2015:MPD

[80] Sushil Kumar Sarmah, Sharmistha Paul, and Subhas Chandra Dey. Mass propagation of Danio aequippinatus, an ornamental fish of North Eastern India. Fisheries and Aquaculture Journal, 6(1):1–2, ???? 2015. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/mass-propagation-of-danio-aequippinatus-an-ornamental-
fish-of-northeastern-india-2150-3508-10000118.pdf.

Daneshfozouna:2015:REI

[81] Hamed Daneshfozouna, Faraz Panjvinib, Fatemeh Ghorbanic, and Hamid Farahmandd. A review of epigenetic imprints in aquatic animals. Fisheries and Aquaculture Journal, 6(1):1–3, ???? 2015. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/a-review-
of-epigenetic-imprints-in-aquatic-animals-2150-3508-10000119.
pdf.

Andrada:2015:WA

[82] Mitchel Andrada. Women in aquaculture. Fisheries and Aquaculture Journal, 6(1):1, ???? 2015. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/women-in-
aquaculture-2150-3508.10000e114.pdf.

Boziaris:2015:CTS

[83] Ioannis S. Boziaris. Current trends on the study of microbiological spoilage of fresh fish. Fisheries and Aquaculture Journal, 6(1):1– 2, ???? 2015. CODEN FAJIAO. ISSN 2150-3508. URL https:/

/www.longdom.org/open-access/current-trends-on-the-study-of-
microbiological-spoilage-of-fresh-fish-2150-3508.10000e115.pdf.

Sharma:2015:CBM

[84] Chhatra M. Sharma. Can bio-manipulation be related to fisheries and aquaculture through environmental pollution perspective? Fisheries and Aquaculture Journal, 6(1):1, ???? 2015. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
can-biomanipulation-be-related-to-fisheries-and-aquaculture-
through-environmental-pollution-perspective.2150-3508.10000e116.
pdf.

Mancuso:2015:FME

[85] Monique Mancuso. Fish mycobateriosis: an emerging disease in Italy. Fisheries and Aquaculture Journal, 6(2):1, ???? 2015. CODEN FAJIAO.

ISSN 2150-3508. URL https://www.longdom.org/open-access/fish-
mycobateriosis-an-emerging-disease-in-italy-2150-3508-1000e120.
pdf.

Andrada:2015:DPA

[86] Mitchel Andrada. The dynamics of Philippine aquaculture. Fisheries and Aquaculture Journal, 6(2):1, ???? 2015. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/the-dynamics-
of-philippine-aquaculture-2150-3508-1000e121.pdf.

G:2015:UPP

[87] Caruso G. Use of plant products as candidate fish meal substitutes: an emerging issue in aquaculture productions. Fisheries and Aquaculture Journal, 6(2):1–3, ???? 2015. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
use-of-plant-products-as-candidate-fish-meal-substitutes-an-
emerging-issue-inaquaculture-productions-2150-3508-1000e123.
pdf.

Anderson:2015:LWR

[88] Wendy E. Anderson and Thomas P. Simon. Length–weight relationship, body morphometrics, and condition based on sexual stage in the rusty crayfish, Orconectes rusticus Girard, 1852 (Decapoda, Cambaridae) with emphasis on management implications. Fisheries and Aquaculture Journal, 6(2):1–7, ???? 2015. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/lengthweight-
relationship-body-morphometrics-and-condition-basedon-sexual-

stage-in-the-rusty-crayfish-orconectes-rusticus-girard-1852decapoda-
2150-3508-1000129.pdf.

Moussa:2015:IA

[89] Ragia Moussa Moussa. Invertebrate aquaculture. Fisheries and Aquaculture Journal, 6(2):1, ???? 2015. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/invertebrate-
aquaculture-2150-3508-1000130.pdf.

Goswami:2015:EDC

[90] Chandasudha Goswami and V. S. Zade. Effect of Daucus carota and Beta vulgaris on color of Anabus testudineus. Fisheries and Aquaculture Journal, 6(3):1–3, ???? 2015. CODEN FAJIAO. ISSN 2150-3508. URL

https://www.longdom.org/open-access/effect-of-daucus-carota-
and-beta-vulgaris-on-color-of-anabus-testudineus-2150-3508-1000132.
pdf.

Michael:2015:EMM

[91] K. G. Michael and O. A. Sogbesan. Evaluation of maggot meal (Muscadomestica) and single cell protein (mushroom) in the diet of Clarias gariepinus fingerlings (Burchell, 1822). Fisheries and Aquaculture Journal, 6(3):1–5, ???? 2015. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/evaluation-of-
maggot-meal-muscadomestica-and-single-cell-proteinmushroom-in-
the-diet-of-clarias-gariepinus-fingerlings-burchell-1822-2150-
3508-1000133.pdf.

Ying:2015:RAS

[92] Liu Ying, Liu Baoliang, Shi Ce, and Sun Guoxiang. Recirculating aquaculture systems in China — current application and prospects. Fisheries and Aquaculture Journal, 6(3):1–3, ???? 2015. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/recirculating-
aquaculture-systems-in-chinacurrent-application-andprospects-
2150-3508-1000134.pdf.

Marudhupandi:2015:PAD

[93] Thangapandi Marudhupandi and Dhinakarasamy Inbakandan. Polysaccharides in aquatic disease management. Fisheries and Aquaculture Journal, 6(3):1–3, ???? 2015. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/polysaccharides-in-
aquatic-disease-management-2150-3508-1000135.pdf.

Ganjoor:2015:SRI

[94] Mohammedsaeed Ganjoor. A short review on infectious viruses in cultural shrimps (Penaeidae family). Fisheries and Aquaculture Journal, 6(3):1–11, ???? 2015. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/a-short-review-on-
infectious-viruses-in-cultural-shrimps-penaeidaefamily-2150-
3508-1000136.pdf.

Riisgaard:2015:JCE

[95] Hans Ulrik Riisgård, Josephine Goldstein, Kim Lundgreen, and Florian Lüskow. Jellyfish and ctenophores in the environmentally degraded Limfjorden (Denmark) during 2014 — species composition, population densities and predation impact. Fisheries and Aquaculture Journal, 6(3):1–10, ???? 2015. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/jellyfish-and-
ctenophores-in-the-environmentally-degraded-limfjordendenmark-
during-2014--species-composition-population-densities-andpredation-
impact-2150-3508-1000137.pdf.

Rostamian:2015:STV

[96] Narges Rostamian, Ebrahim Masoudi, Mohammad Hasan Gerami, Sirvan Azizpour, and Sana Ullah. Spatial and temporal variability of phytoplankton assemblages and physico-chemical characterization in three similar dams. Fisheries and Aquaculture Journal, 6(3):1–4, ???? 2015. CODEN FAJIAO. ISSN 2150-3508. URL https://www.

longdom.org/open-access/spatial-and-temporal-variability-of-
phytoplankton-assemblages-andphysicochemical-characterization-
in-three-similar-dams-2150-3508-1000138.pdf.

Olaniyan:2015:FMT

[97] R. F. Olaniyan. Fishing methods and their implications for a sustainable environment. Fisheries and Aquaculture Journal, 6(3):1–3, ???? 2015. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/fishing-methods-and-their-implications-for-a-sustainable-
environment-2150-3508-1000139.pdf.

K:2015:MSF

[98] Roy K. A model strategic framework for prioritization and development of inland water bodies under fisheries and aquaculture. Fisheries and Aquaculture Journal, 6(3):1–6, ???? 2015. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/a-model-strategic-
framework-for-prioritization-and-development-ofinland-water-
bodies-under-fisheries-and-aquaculture-2150-3508-1000140.pdf.

Wakchaure:2015:ITF

[99] Rajesh Wakchaure, Subha Ganguly, Kausar Qadri, Praveen Kumar Praveen, and Tanvi Mahajan. Importance of transgenic fish to global aquaculture: a review. Fisheries and Aquaculture Journal, 6(4):1– 3, ???? 2015. CODEN FAJIAO. ISSN 2150-3508. URL https:

//www.longdom.org/open-access/importance-of-transgenic-fish-
to-global-aquaculture-a-review-2150-3508-1000e124.pdf.

Sogbesan:2015:UTD

[100] O. A. Sogbesan, C. F. Onoja, H. A. Adedeji, and T. A. Idowu. Utilization of treated duckweed meal (Lemna pausicostata) as plant protein supplement in African mud catfish (Clarias gariepinus) juvenile diets. Fisheries and Aquaculture Journal, 6(4):1–5, ???? 2015. CODEN FAJIAO. ISSN 2150-3508. URL https://

www.longdom.org/open-access/utilization-of-treated-duckweed-
meal-lemna-pausicostata-as-plantprotein-supplement-in-african-
mud-catfish-clarias-gariepinus-juven-2150-3508-1000141.pdf.

Khan:2015:CCF

[101] Mohammad Ali Khan, Joydev Ghosh, and Banshidhar Sahoo. Controlling chaos in a food chain model through threshold harvesting. Fisheries and Aquaculture Journal, 6(4):1–4, ???? 2015. CO-DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/controlling-chaos-in-a-food-chain-model-through-
threshold-harvesting-2150-3508-1000142.pdf.

Zehev:2015:OFR

[102] Benzaken S. Zehev, Almeida Vera, Benzaken Asher, and Ribeiro Raimundo. Ornamental fishery in Rio Negro (Amazon Region), Brazil: Combining social, economic and fishery analyses. Fisheries and Aquaculture Journal, 6(4):1–4, ???? 2015. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/ornamental-fishery-in-rio-negro-amazon-region-brazil-
combiningsocial-economic-and-fishery-analyses-2150-3508-1000143.
pdf.

Md:2015:PYS

[103] Monjurul Hasan Md, Bhakta Supratim Sarker, Mahabubur Rahman, Shamsul Alam Patwary Md, Jahangir Sarker Md, K. M. Shahriar Nazrul, and Mohammed Rashed Parvej. Present yield status, percentage composition and seasonal abundance of shark in two geographically important zones of Bangladesh. Fisheries and Aquaculture Journal, 6(4):1–6, ????

2015. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/present-yield-status-percentage-composition-
and-seasonal-abundance-ofshark-in-two-geographically-important-
zones-of-bangladesh-2150-3508-1000144.pdf.

Oketoki:2015:SPB

[104] T. O. Oketoki. Survey on phytoplankton biomass and water parameters in the habitats of invasive tigers shrimps (Penaeus monodon) in Nigeria. Fisheries and Aquaculture Journal, 6(4):1–12, ???? 2015. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
survey-on-phytoplankton-biomass-and-water-parameters-in-the-
habitatsof-invasive-tigers-shrimps-penaeus-monodon-in-nigeria-
2150-3508-1000145.pdf.

Kannan:2015:PMS

[105] Kannan D, Thirunavukkarasu P, Jagadeesan K, Shettu N, and Aswini Kumar. Procedure for maturation and spawning of imported shrimp Litopenaeus vannamei in commercial hatchery, South East Coast of India. Fisheries and Aquaculture Journal, 6(4):1–5, ???? 2015. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
procedure-for-maturation-and-spawning-of-imported-shrimp-litopenaeusvannamei-
in-commercial-hatchery-south-east-coast-of-india-2150-3508-1000146.
pdf.

Ahmad:2015:MMC

[106] Fayaz Ahmad, Khalid M. Fazili, Tanveer A. Sofi, Bashir A. Sheikh, Ajaz A. Waza, Rabiya Rashid, and Tantry Tariq Gani. Morphological and molecular characterization of Diplozoon kashmirensis; D. aegyptensis and D. guptai collected from fishes of Kashmir Valley — India. Fisheries and Aquaculture Journal, 6(4):1–9, ???? 2015. CO-DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/morphological-and-molecular-characterization-of-
diplozoon-kashmirensisd-aegyptensis-and-d-guptai-collected-from-
fishes-of-kashmir-2150-3508-1000147.pdf.

Naik:2015:CPC

[107] Gulzar Naik, Mudasir Rashid, and M. H. Balkhi. Changes in physicochemical parameters at different sites of Manasbal Lake of Kashmir, India. Fisheries and Aquaculture Journal, 6(4):1–4, ???? 2015. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
changes-in-physicochemical-parameters-at-different-sites-of-
manasballake-of-kashmir-india-2150-3508-1000148.pdf.

Hadied:2015:STA

[108] Marwa Abou Hadied, Abdel Aziz Khidr, Ola Abu Samak, and Ashraf Said. Surface topography of the anterior adhesive apparatus of the gill monogenean parasite Diplectanum sp. diesing, 1858, with some surface criteria. Fisheries and Aquaculture Journal, 6(4):1–5, ???? 2015. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/surface-topography-of-the-anterior-adhesive-
apparatus-of-the-gillmonogenean-parasite-diplectanum-sp-diesing-
1858-with-some-surface-2150-3508-1000150.pdf.

Raja:2015:LSS

[109] Kuzhanthaivel Raja, A. Gopalakrishnan, Rajkumar Singh, and R. Vijayakumar. Loose shell syndrome (LSS) in Litopenaeus vannamei growout ponds and its effect on growth and production. Fisheries and Aquaculture Journal, 6(4):1–4, ???? 2015. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/loose-shell-
syndrome-lss-in-litopenaeus-vannamei-growout-ponds-andits-effect-
on-growth-and-production-2150-3508-1000151.pdf.

Al-Beak:2015:PDS

[110] Ahmed M. Al-Beak, S. I. Ghoneim, A. Y. El-Dakar, and M. Salem. Population dynamic and stock assesment of white seabream Diplodus sargus (Linnaeus, 1758) in the Coast of North Siani. Fisheries and Aquaculture Journal, 6(4):1–6, ???? 2015. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/population-dynamic-
and-stock-assesment-of-white-seabream-diplodussargus-linnaeus-
1758-in-the-coast-of-north-siani-2150-3508-1000152.pdf.

El-Khayat:2015:SFP

[111] Hanaa M. M. El-Khayat, Hoda Abdel-Hamid, Hanan S. Gaber, Kadria M. A. Mahmoud, and Hassan E. Flefel. Snails and fish as pollution biomarkers in Lake Manzala and Laboratory A: Lake Manzala snails. Fisheries and Aquaculture Journal, 6(4):1–9, ???? 2015. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
snails-and-fish-as-pollution-biomarkers-in-lake-manzala-and-
laboratorya-lake-manzala-snails-2150-3508-1000153.pdf.

Getu:2015:PHM

[112] Addis Getu, Kidanie Misganaw, and Meseret Bazezew. Post-harvesting and major related problems of fish production. Fisheries and Aquaculture Journal, 6(4):1–6, ???? 2015. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/postharvesting-and-
major-related-problems-of-fish-production-2150-3508-1000154.
pdf.

Naik:2015:FFH

[113] Gulzar Naik, Mudasir Rashid, M. H. Balkhi, and F. A. Bhat. Food and feeding habits of Cyprinus carpio Var. communis: a reason that decline schizothoracine fish production from Dal Lake of Kashmir Valley. Fisheries and Aquaculture Journal, 6(4):1–5, ???? 2015. CODEN FAJIAO. ISSN 2150-3508. URL https://www.

longdom.org/open-access/food-and-feeding-habits-of-cyprinus-
carpio-var-communis-a-reasonthat-decline-schizothoracine-fish-
production-from-dal-lake-of-kash-2150-3508-1000155.pdf.

Dutta:2016:FCD

[114] Binku Dutta, Nabin Ch Das, and Devashish Kar. The fish catching devices with their efficacy and cost-benefit analysis in the Towkak River in Assam and Nagaland, India. Fisheries and Aquaculture Journal, 7(1): 1–5, ???? 2016. CODEN FAJIAO. ISSN 2150-3508. URL https:/

/www.longdom.org/open-access/the-fish-catching-devices-with-
their-efficacy-and-costbenefit-analysis-inthe-towkak-river-in-
assam-and-nagaland-india-2150-3508-1000156.pdf.

Md:2016:MCS

[115] Jahangir Sarker Md, Shamsul Alam Patwary Md, A. M. M. Borhan Uddin, Monjurul Hasan Md, Mehedi Hasan Tanmay, Indrani Kanungo, and Mohammed Rashed Parvej. Macrobenthic community structure — an approach to assess coastal water pollution in Bangladesh. Fisheries and Aquaculture Journal, 7(1):1–10, ???? 2016. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/macrobenthic-
community-structure--an-approach-to-assess-coastal-water-pollution-
in-bangladesh-2150-3508-1000157.pdf.

Dar:2016:IAG

[116] Sabzar Ahmad Dar, Abdul Rehman Yousuf, and Masood ul Hassan Balkhi. An introduction about genotoxicology methods as tools for monitoring aquatic ecosystem: Present status and future perspectives. Fisheries and Aquaculture Journal, 7(1):1–11, ???? 2016. CO-DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/an-introduction-about-genotoxicology-methods-as-
tools-for-monitoring-aquatic-ecosystem-present-status-and-future-
perspectives-2150-3508-1000158.pdf.

Md:2016:SDH

[117] Jahangir Sarker Md, Indrani Kanungo, Mehedi Hasan Tanmay, and Shamsul Alam Patwary Md. A study on the determination of heavy metals in sediment of fish farms in Bangladesh. Fisheries and Aquaculture Journal, 7(1):1–5, ???? 2016. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/a-study-on-
the-determination-of-heavy-metals-in-sediment-of-fish-farms-
in-bangladesh-2150-3508-1000159.pdf.

Kumar:2016:PFE

[118] Rohit Kumar, Veena Pande, Lalit Singh, Lata Sharma, Neha Saxena, Dimpal Thakuria, Atul K. Singh, and Prabhati K. Sahoo. Pathological findings of experimental Aeromonas hydrophila infection in golden mahseer (Tor putitora). Fisheries and Aquaculture Journal, 7(1):1–6, ???? 2016. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/pathological-
findings-of-experimental-aeromonas-hydrophila-infection-in-golden-
mahseer-tor-putitora-2150-3508-1000160.pdf.

Wang:2016:PUM

[119] Guoqiang Wang and Iain J. McGaw. Potential use of mussel farms as multitrophic on-growth sites for American lobster, Homarus americanus (Milne Edwards). Fisheries and Aquaculture Journal, 7(1):1– 11, ???? 2016. CODEN FAJIAO. ISSN 2150-3508. URL https:

//www.longdom.org/open-access/potential-use-of-mussel-farms-
as-multitrophic-ongrowth-sites-for-american-lobster-homarus-
americanus-milne-edwards-2150-3508-1000161.pdf.

Yones:2016:IDS

[120] Abdel Moneim Yones M and Atallah Metwalli A. Influence of dietary Sorghum starch on growth performance, digestibility coefficient and some hepatic enzyme activities in hybrid red tilapia (Oreochromis mossambicus × Oreochromis niloticus) fingerlings. Fisheries and Aquaculture Journal, 7(1):1–8, ???? 2016. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/influence-of-
dietary-sorghum-starch-on-growth-performance-digestibility-coefficient-
and-some-hepatic-enzyme-activities-in-hybrid-r-2150-3508-1000162.
pdf.

Brown:2016:RFC

[121] Katherine R. Brown, Michael E. Barnes, Timothy M. Parker, and Brian Fletcher. Retention of fillet coloration in rainbow trout after dietary astaxanthin cessation. Fisheries and Aquaculture Journal, 7(1):1–3, ???? 2016.

CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/retention-of-fillet-coloration-in-rainbow-trout-
after-dietary-astaxanthin-cessation-2150-3508-1000163.pdf.

Cappello:2016:EN

[122] Tiziana Cappello. Editor note. Fisheries and Aquaculture Journal, 7(2): 1, ???? 2016. CODEN FAJIAO. ISSN 2150-3508. URL https://www.

longdom.org/open-access/editor-note-2150-3508-1000e126.pdf.

Bhouiyan:2016:IID

[123] Naser Ahmed Bhouiyan, Mohammad Abdul Baki, Anirban Sarker, and Md. Muzammel Hossain. Inventory of ichyofaunal diversity, fishing gear and craft in Turag River, Dhaka, Bangladesh. Fisheries and Aquaculture Journal, 7(2):1–6, ???? 2016. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/inventory-of-
ichyofaunal-diversity-fishing-gear-and-craft-in-turag-river-
dhaka-bangladesh-2150-3508-1000165.pdf.

Peng:2016:PPR

[124] Jianpeng Peng, Anji Lian, and Quan Jiang. Production and purification of recombinant somatolactin and its effects on insulin-like growth factors gene expression in tilapia hepatocytes. Fisheries and Aquaculture Journal, 7(2):1–7, ???? 2016. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/production-and-
purification-of-recombinant-somatolactin-and-its-effects-on-
insulinlike-growth-factors-gene-expression-in-tilapia-h-2150-
3508-1000167.pdf.

Thiaw:2016:SSS

[125] Modou Thiaw, Mbaye Tine, Hamet Diaw Diadhiou, Oualbadet Magoma, and Patrice Brehmer. Shift of small scale fishing impacts on fish trophic levels in Lake Iro revealed by species-based indicators. Fisheries and Aquaculture Journal, 7(2):1–8, ???? 2016. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/shift-of-small-scale-fishing-impacts-on-fish-trophic-
levels-in-lake-iro-revealed-by-speciesbased-indicators-2150-
3508-1000166.pdf.

Md:2016:LSH

[126] Jahangir Sarker Md, A. M. M. Borhan Uddin, Shamsul Alam Patwary Md, Mehedi Hasan Tanmay, Farhana Rahman, and Moshiur Rahman. Livelihood status of hilsa (Tenualosa ilisha) fishermen of Greater Noakhali Regions of Bangladesh. Fisheries and Aquaculture Journal, 7(2):1–6, ????

2016. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/livelihood-status-of-hilsa-tenualosa-ilisha-
fishermen-of-greater-noakhali-regions-of-bangladesh-2150-3508-
1000168.pdf.

Bergmann:2016:TSS

[127] Sven M. Bergmann, Michael Cieslak, Dieter Fichtner, Juliane Dabels, Sean J. Monaghan, Qing Wang, Weiwei Zeng, and Jolanta Kempter. Is there any species specificity in infections with aquatic animal herpesviruses? The koi herpesvirus (KHV): an alloherpesvirus model. Fisheries and Aquaculture Journal, 7(2):1–5, ???? 2016. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
is-there-any-species-specificity-in-infections-with-aquatic-
animalherpesvirusesthe-koi-herpesvirus-khv-an-alloherpesvirus-
model-2150-3508-1000169.pdf.

Gupta:2016:RFR

[128] Sandipan Gupta and Samir Banerjee. A review on feeding and reproductive biology of Cirrhinus reba (Hamilton, 1822), a threatened freshwater fish of Indian subcontinent with an emphasis on its conservation. Fisheries and Aquaculture Journal, 7(2):1–4, ???? 2016. CO-DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/a-review-on-feeding-and-reproductive-biology-of-
cirrhinus-reba-hamilton1822-a-threatened-freshwater-fish-of-
indian-subcontinent-wi-2150-3508-1000170.pdf.

Latuihamallo:2016:PNF

[129] Magdalena Latuihamallo, Joice W. Loupatty, and Gratia Dolores Manuputty. The proximate of natural foods Gracilaria lichenoides and Ulva fasciata for abalone Haliotis squamata culture. Fisheries and Aquaculture Journal, 7(2):1–5, ???? 2016. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/the-proximate-
of-natural-foods-gracilaria-lichenoides-and-ulva-fasciata-for-
abalone-haliotis-squamata-culture-2150-3508-1000171.pdf.

Jeannette:2016:LET

[130] Tombi Jeannette, Akoumba John Francis, Mieguim Ngninpogni Dominique, and Bilong Bilong Charles Felix. Longitudinal exploitation of the transversal gradient of Oreochromis niloticus gill system by four monogeneans species at Melen Fish Station (Yaounde, Cameroon). Fisheries and Aquaculture Journal, 7(2):1–4, ???? 2016. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/longitudinal-
exploitation-of-the-transversal-gradient-of-oreochromisniloticus-

gill-system-by-four-monogeneans-species-at-melen-fis-2150-3508-
1000172.pdf.

Watanabe:2016:SFC

[131] Kyuji Watanabe. In-season forecast of chum salmon return using smoothing spline. Fisheries and Aquaculture Journal, 7(2):1–4, ???? 2016. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/inseason-forecast-of-chum-salmon-return-using-smoothing-
spline-2150-3508-1000173.pdf.

Owen:2016:BRH

[132] Kevin C. Owen. The beneficial reuse of hypersaline waste water from desalination plants to treat harmful algal blooms. Fisheries and Aquaculture Journal, 7(3):1–3, ???? 2016. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/the-beneficial-reuse-of-hypersaline-waste-water-from-
desalination-plants-to-treat-harmful-algal-blooms-2150-3508-
1000e125.pdf.

Misganaw:2016:MLC

[133] Kidanie Misganaw and Addis Getu. Marketing and livelihood contribution of fishermen in Lake Tana, North Western part of Ethiopia. Fisheries and Aquaculture Journal, 7(3):1–5, ???? 2016. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/marketing-
and-livelihood-contribution-of-fishermen-in-lake-tana-north-
western-part-of-ethiopia-2150-3508-1000174.pdf.

Misganaw:2016:RMP

[134] Kidanie Misganaw and Addis Getu. Review on major parasitic crustacean in fish. Fisheries and Aquaculture Journal, 7(3):1–5, ???? 2016. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/review-on-major-parasitic-crustacean-in-fish-2150-3508-
1000175.pdf.

Ganjoor:2016:RMF

[135] Saeed Ganjoor M. To reduce mortality of fry fish (Oncorhynchus mykiss) caused with viral infection (IPNV and VHSV) by water treatment with chloramin-T as disinfectant. Fisheries and Aquaculture Journal, 7(3):1–5, ???? 2016. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/to-reduce-mortality-
of-fry-fish-oncorhynchus-mykiss-caused-with-viral-infection-
ipnv-and-vhsv-by-water-treatment-with-chloramint-a-2150-3508-
1000176.pdf.

Samanta:2016:HSS

[136] Palas Samanta, Sandipan Pal, Aloke Kumar Mukherjee, Debraj Kole, and Apurba Ratan Ghosh. Histopathological study in stomach and intestine of Anabas testudineus (Bloch, 1792) under Almix exposure. Fisheries and Aquaculture Journal, 7(3):1–6, ???? 2016. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
histopathological-study-in-stomach-and-intestine-of-anabas-testudineus-
bloch-1792-under-almix-exposure-2150-3508-1000177.pdf.

Kareem:2016:GPC

[137] O. K. Kareem, A. N. Olanrewaju, E. F. Osho, O. Orisasona, and M. A. Akintunde. Growth patterns and condition factor of Hepsetus odoe (Bloch, 1794) captured in Eleyele Lake, Southwest Nigeria. Fisheries and Aquaculture Journal, 7(3):1–4, ???? 2016. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
growth-patterns-and-condition-factor-of-hepsetus-odoe-bloch-
1794-captured-in-eleyele-lake-southwest-nigeria-2150-3508-1000178.
pdf.

Raja:2016:ZWM

[138] Manickam Raja, Ramalingam Karthik Raja, and Pachiappan Perumal. Zebrafish in the wild: Microhabitat use by zebrafish Danio rerio (Hamilton, 1822) from Karala River of Jalpaiguri District, Northern Bengal, India. Fisheries and Aquaculture Journal, 7(4):1–6, ???? 2016. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
zebrafish-in-the-wild-microhabitat-use-by-zebrafish-danio-rerio-
hamilton-1822-from-karala-river-of-jalpaiguri-district-northern-
be-2150-3508-1000179.pdf.

Espino-Barr:2016:RMC

[139] Elaine Espino-Barr, Manuel Gallardo-Cabello, Marcos Puente-Gómez, and Arturo Garcia-Boa. Reproduction of Mugil cephalus (Percoidei: Mugilidae) off the Central Mexican Pacific Coast. Fisheries and Aquaculture Journal, 7(4):1–9, ???? 2016. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/reproduction-
of-mugil-cephalus-percoidei-mugilidae-off-the-centralmexican-
pacific-coast-2150-3508-1000180.pdf.

Mamat:2016:PCV

[140] Nawwar Zawani Mamat, Mohd Idrus Shaari, and Nur Amirul Anas Abdul Wahab. The production of catfish and vegetables in an aquaponic system. Fisheries and Aquaculture Journal, 7(4):1–3, ???? 2016.

CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/the-production-of-catfish-and-vegetables-in-
an-aquaponic-system-2150-3508-1000181.pdf.

Asmare:2016:FJW

[141] Erkie Asmare, Sewmehon Demissie, and Dereje Tewabe. Fisheries of Jemma and Wonchit Rivers: As a means of livelihood diversification and its challenges in North Shewa Zone, Ethiopia. Fisheries and Aquaculture Journal, 7(4):1–6, ???? 2016. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/fisheries-of-jemma-
and-wonchit-rivers-as-a-means-of-livelihooddiversification-and-
its-challenges-in-north-shewa-zone-ethiopia-2150-3508-1000182.
pdf.

Oke:2016:GPF

[142] Vincent Oké, Youssouf Abou, Alphonse Adité, and Jean-André T. Kabré. Growth performance, feed utilization and body composition of Clarias gariepinus (Burchell 1822) fed marine fish viscera-based-diet in earthen ponds. Fisheries and Aquaculture Journal, 7(4):1–7, ???? 2016. CODEN FAJIAO. ISSN 2150-3508. URL https://www.

longdom.org/open-access/growth-performance-feed-utilization-
and-body-composition-of-clariasgariepinus-burchell-1822-fed-
marine-fish-viscerabaseddiet-in-ea-2150-3508-1000183.pdf.

Kebede:2016:IIE

[143] B. Kebede and T. Habtamu. Isolation and identification of Edwardsiella tarda from Lake Zeway and Langano, Southern Oromia, Ethiopia. Fisheries and Aquaculture Journal, 7(4):1–6, ???? 2016. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
isolation-and-identification-of-edwardsiella-tarda-from-lake-
zeway-andlangano-southern-oromia-ethiopia-2150-3508-1000184.
pdf.

Oyediran:2016:CLE

[144] W. O. Oyediran, A. M. Omoare, O. B. Oladoyinbo, B. O. Ajagbe, and T. T. Dick. Constraints limiting the effective utilization of low-cost fish processing technologies among women in selected fishing communities of Lagos State, Nigeria. Fisheries and Aquaculture Journal, 7(4):1–5, ???? 2016. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/constraints-
limiting-the-effective-utilization-of-lowcost-fish-processingtechnologies-
among-women-in-selected-fishing-communities-2150-3508-1000185.
pdf.

Zadlock:2017:NAA

[145] Frank J. Zadlock IV, Satshil B. Rana, Zain A. Alvi, Ziping Zhang, Wyatt Murphy, and Carolyn S. Bentivegna. De Novo assembly and analysis of the testes transcriptome from the menhaden, Bervoortia tyrannus. Fisheries and Aquaculture Journal, 8(1):1–8, ???? 2017. CODEN FAJIAO.

ISSN 2150-3508. URL https://www.longdom.org/open-access/de-
novo-assembly-and-analysis-of-the-testes-transcriptome-from-
the-menhaden-bervoortia-tyrannus-2150-3508-1000186.pdf.

Khan:2017:EDP

[146] Imtiaz Ahmed Khan and Amir Maqbool. Effects of dietary protein levels on the growth, feed utilization and haemato-biochemical parameters of freshwater fish, Cyprinus carpio Var. specularis. Fisheries and Aquaculture Journal, 8(1):1–12, ???? 2017. CODEN

FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/effects-of-dietary-protein-levels-on-the-growth-feed-
utilization-and-haematobiochemical-parameters-of-freshwater-
fish-cyprinus-car-2150-3508-1000187.pdf.

Ayeloja:2017:PMP

[147] A. A. Ayeloja, F. George, E. Sodeeq, and G. L. Adebisi. Price modulation policy of Federal Government of Nigeria: Effects on fish production. Fisheries and Aquaculture Journal, 8(1):1–4, ???? 2017. CODEN FAJIAO.

ISSN 2150-3508. URL https://www.longdom.org/open-access/price-
modulation-policy-of-federal-government-of-nigeria-effects-on-
fish-production-2150-3508-1000188.pdf.

Wehye:2017:GME

[148] A. S. Wehye, S. K. K. Amponsah, and A. S. Jueseah. Growth, mortality and exploitation of Sardinella maderensis (Lowe, 1838) in the Liberian coastal waters. Fisheries and Aquaculture Journal, 8(1):1–5, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/growth-mortality-and-exploitation-of-sardinella-
maderensis-lowe-1838-in-the-liberian-coastal-waters-2150-3508-
1000189.pdf.

Li:2017:MIA

[149] Yajuan Li, D. L. Ren, M. Chen, S. C. Ge, and Bing Hu. Myeloperoxidase inactivation affects neutrophil recruitment in zebrafish injuryinduced model. Fisheries and Aquaculture Journal, 8(1):1–7, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/myeloperoxidase-inactivation-affects-neutrophil-

recruitment-in-zebrafishinjuryinduced-model-2150-3508-1000190.
pdf.

Abedalhammed:2017:ENH

[150] Hazem S. Abedalhammed, Nasreen M. Abdulrahman, and Haitham L. Sadik. Effect of natural and hydroponic barley plant and sprout on the common carp (Cyprinus carpio) growth performances. Fisheries and Aquaculture Journal, 8(1):1–4, ???? 2017. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
effect-of-natural-and-hydroponic-barley-plant-and-sprout-on-
the-common-carp-cyprinus-carpio-growth-performances-2150-3508-
1000191.pdf.

Harlioglu:2017:IFS

[151] M. M. Harlioglu and A. Farhadi. Iranian fisheries status: an update (2004–2014). Fisheries and Aquaculture Journal, 8(1):1–8, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/iranian-fisheries-status-an-update-20042014-
2150-3508-1000192.pdf.

Patra:2017:ENC

[152] B. C. Patra, S. Patra, and M. Bhattacharya. Evaluating the nutritional condition of an Indian climbing perch, Anabas testudineus fingerlings by the RNA/DNA, Ca/P ratio and protein bio-synthesis in liver and muscle. Fisheries and Aquaculture Journal, 8(1):1–8, ???? 2017. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
evaluating-the-nutritional-condition-of-an-indian-climbing-perch-
anabastestudineus-fingerlings-by-the-rnadna-cap-ratio-and-protein-
2150-3508-1000193.pdf.

Yimer:2017:MES

[153] Adamu Yimer, Minwyelet Mingist, and Behailu Bekele. Microbial evaluation of selected post harvest processing techniques for quality fish product at Bahir Dar Town, Ethiopia. Fisheries and Aquaculture Journal, 8(2):1–5, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL

https://www.longdom.org/open-access/microbial-evaluation-of-
selected-post-harvest-processing-techniques-for-quality-fish-
product-at-bahir-dar-town-ethiopia-2150-3508-1000194.pdf.

Chow:2017:DEB

[154] Edwin Pei Yong Chow, Kah Heng Liong, and Elke Schoeters. Dietary encapsulated butyric acid (Butipearl^{T M}) and microemulsified carotenoids (quantum GLO^{T M} Y) on the growth, immune parameters and their synergistic effect on pigmentation of hybrid catfish (Clarias macrocephalus × Clarias gariepinus). Fisheries and Aquaculture Journal, 8(2):1–6, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/dietary-encapsulated-butyric-acid-butipearl-
and-microemulsifiedcarotenoids-quantum-glo-y-on-the-growth-immune-
parameters-and-their-2150-3508-1000195.pdf.

Ozdemir:2017:DHS

[155] Rahmi Can Ozdemir and Ekici Aygül. Different heat shock application effect on gynogenetic production of zebrafish (Danio rerio). Fisheries and Aquaculture Journal, 8(2):1–6, ???? 2017. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/different-heat-
shock-application-effect-on-gynogenetic-production-of-zebrafish-
danio-rerio-2150-3508-1000196.pdf.

Huicab-Pech:2017:PBO

[156] Z. G. Huicab-Pech, M. R. Castaneda-Chavez, and F. Lango-Reynoso. Pathogenic bacteria in Oreochromis niloticus var. Styrling tilapia culture. Fisheries and Aquaculture Journal, 8(2):1–7, ???? 2017. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
pathogenic-bacteria-in-oreochromis-niloticus-var-stirling-tilapia-
culture-2150-3508-1000197.pdf.

Long:2017:CIF

[157] Stephen Long, Richard Ffrench-Constant, Kristian Metcalfe, and Matthew J. Witt. Have centuries of inefficient fishing sustained a wild oyster fishery: a case study. Fisheries and Aquaculture Journal, 8(2): 1–7, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL https:

//www.longdom.org/open-access/have-centuries-of-inefficient-
fishing-sustained-a-wild-oyster-fishery-a-case-study-2150-3508-
1000198.pdf.

Ikenna:2017:BFF

[158] Obiyor Kelvin Ikenna, Nwani Chris Didigwu, Odo Gregory Ejikeme, Madu Josephine Chinenye, Ndudim Doris Ulumma Aguzie, and Ifeanyi Oscar Ndimkaoha. Benthic fish fauna and physicochemical parameters of Otamiri River, Imo State, Nigeria. Fisheries and Aquaculture Journal, 8(2):1–8, ???? 2017. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/benthic-fish-fauna-and-physicochemical-parameters-of-
otamiri-river-imo-state-nigeria-2150-3508-1000199.pdf.

Kalhoro:2017:SAI

[159] Muhammad Talib Kalhoro, Mu Yongtong, Muhsan Ali Kalhoro, Memon Aamir Mahmood, Shah Syed Babar Hussain, Mohsin Muhammad, and Pavase Tushar Ramesh. Stock assessment of Indian scad, Decapterus russelli in Pakistani marine waters and its impact on the national economy. Fisheries and Aquaculture Journal, 8(2):1–10, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL https://

www.longdom.org/open-access/stock-assessment-of-indian-scad-
decapterus-russelli-in-pakistani-marine-waters-and-its-impact-
on-the-national-economy-2150-3508-1000200.pdf.

Wehye:2017:PDP

[160] Austin Saye Wehye, Patrick K. Ofori-Danson, and Angela Manekuor Lamptey. Population dynamics of Pseudotolithus senegalensis and Pseudotolithus typus and their implications for management and conservation within the coastal waters of Liberia. Fisheries and Aquaculture Journal, 8(2):1–9, ???? 2017. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/population-dynamics-
of-pseudotolithus-senegalensis-and-pseudotolithus-typus-and-
their-implications-for-management-and-conservation-2150-3508-
1000201.pdf.

Ahmed:2017:ESV

[161] Ahmed Mohammed Musa Ahmed. Effects of seasonal variation on fish catching in Jebel Aulia Reservoir on the White Nile, Sudan. Fisheries and Aquaculture Journal, 8(2):1–5, ???? 2017. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/effects-of-
seasonal-variation-on-fish-catching-in-jebel-aulia-reservoir-
on-the-white-nile-sudan-2150-3508-1000202.pdf.

Pandey:2017:FSM

[162] Dipak Pandey, Yong-Woon Ryu, and Takahiro Matsubara. Features of sperm motility and circadian rhythm in Japanese anchovy (Engraulis japonicus). Fisheries and Aquaculture Journal, 8(2):1–5, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/features-of-sperm-motility-and-circadian-rhythm-
in-japanese-anchovyengraulis-japonicus-2150-3508-1000203.pdf.

Jakobsen:2017:FSF

[163] Jette Jakobsen and Cat Smith. Farmed salmon and farmed rainbow trout — excellent sources of vitamin D? Fisheries and Aquaculture Journal, 8(2):1–5, ???? 2017. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/farmed-salmon-and-
farmed-rainbow-trout--excellent-sources-of-vitamin-d-2150-3508-
1000204.pdf.

Dwivedi:2017:SSA

[164] Amitabh Chandra Dwivedi, Priyanka Mayank, and Ashish Tiwari. Size selectivity of active fishing gear: Changes in size, age and growth of Cirrhinus mrigala from the Ganga River, India. Fisheries and Aquaculture Journal, 8(3):1–5, ???? 2017. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
size-selectivity-of-active-fishing-gear-changes-in-size-age-
and-growth-of-cirrhinus-mrigala-from-the-ganga-river-india-2150-
3508-1000205.pdf.

Benoit:2017:SDM

[165] L. F. G. Benoit and F. Abraham. Spatial distribution of Myxobolus pethericii and Henneguya pethericii on the gills of an African Anabantid Ctenopoma petherici from the Sange River, Cameroon. Fisheries and Aquaculture Journal, 8(3):1–6, ???? 2017. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/spatial-
distribution-of-myxobolus-pethericii-and-henneguya-pethericii-
on-the-gills-of-an-african-anabantid-ctenopoma-petherici-fro-
2150-3508-1000206.pdf.

Shovon:2017:HML

[166] M. N. H. Shovon, B. C. Majumdar, and Z. Rahman. Heavy metals (lead, cadmium and nickel) concentration in different organs of three commonly consumed fishes in Bangladesh. Fisheries and Aquaculture Journal, 8 (3):1–5, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL https:

//www.longdom.org/open-access/heavy-metals-lead-cadmium-and-
nickel-concentration-in-different-organs-of-three-commonly-consumed-
fishes-in-bangladesh-2150-3508-1000207.pdf.

Azua:2017:VMM

[167] E. T. Azua, T. J. Akaahan, and S. A. Akogwu. Variation in the morphometry measurements of two tilapia fish species in relation to their body weight obtained from Lower Benue River at Makurdi, Benue State Nigeria. Fisheries and Aquaculture Journal, 8(3):1–4, ???? 2017. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
variation-in-the-morphometry-measurements-of-two-tilapia-fish-
species-inrelation-to-their-body-weight-obtained-from-lower-
benue-ri-2150-3508-1000208.pdf.

Cherif:2017:NWF

[168] Nadia Cherif and Amdouni Fatma. Nodaviruses in wild fish population collected around aquaculture cage sites from coastal areas of Tunisia. Fisheries and Aquaculture Journal, 8(3):1–6, ???? 2017. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/nodaviruses-in-wild-fish-population-collected-around-
aquaculture-cage-sites-from-coastal-areas-of-tunisia-2150-3508-
1000209.pdf.

Gandotra:2017:GRJ

[169] Roopma Gandotra, Ritu Kumari, Monika Sharma, and Dalbir Singh. Growth response of juveniles of rohu, Labeo rohita to different levels of lipid in the diet. Fisheries and Aquaculture Journal, 8(3):1– 4, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL https://

www.longdom.org/open-access/growth-response-of-juveniles-of-
rohu-labeo-rohita-to-different-levels-oflipid-in-the-diet-2150-
3508-1000210.pdf.

Cheal:2017:MSS

[170] Jason Cheal, Anup Chamrajnagar, Xander Fong, and John Glance. A model for self-sustaining Litopenaeus vannamei farm alternatives. Fisheries and Aquaculture Journal, 8(3):1–8, ???? 2017. CODEN

FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/a-model-for-selfsustaining-litopenaeus-vannamei-farm-
alternatives-2150-3508-1000211.pdf.

Al-Zaidan:2017:AEI

[171] Abdullah Salim Al-Zaidan. The acute effects of un-ionized ammonia on zebrafish (Danio rerio). Fisheries and Aquaculture Journal, 8(3): 1–10, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL https:/

/www.longdom.org/open-access/the-acute-effects-of-unionized-
ammonia-on-zebrafish-danio-rerio-2150-3508-1000212.pdf.

Melaku:2017:IIP

[172] Hamere Melaku, Matios Lakew, Esayas Alemayehu, Alemayehu Wubie, and Mersha Chane. Isolation and identification of pathogenic fungus from African catfish (Clarias gariepinus) eggs and adults in National Fishery and Aquatic Life Research Center Hatchery, Ethiopia. Fisheries and Aquaculture Journal, 8(3):1–5, ???? 2017. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/isolation-and-
identification-of-pathogenic-fungus-from-african-catfish-clarias-
gariepinus-eggs-and-adults-in-national-fishery-and-2150-3508-
1000213.pdf.

Stutzman:2017:URF

[173] Emily Stutzman, Joseph Molnar, Gertrude Atukunda, and John Walakira. Understanding the role of fish farmer associations as intermediaries for the commercialization of aquaculture in Uganda. Fisheries and Aquaculture Journal, 8(3):1–12, ???? 2017. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/understanding-the-
role-of-fish-farmer-associations-as-intermediaries-for-the-commercialization-
of-aquaculture-in-uganda-2150-3508-1000214.pdf.

Laimeheriwa:2017:PRS

[174] Bruri Melky Laimeheriwa. Phenetic relationship study of gold ring cowry, Cypraea annulus (Gastropods: Cypraeidae) in Mollucas Islands based on shell morphological. Fisheries and Aquaculture Journal, 8(3):1–15, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL https://www.

longdom.org/open-access/phenetic-relationship-study-of-gold-
ring-cowry-cypraea-annulus-gastropods-cypraeidae-in-mollucas-
islands-based-on-shell-morphologi-2150-3508-1000215.pdf.

Akinloye:2017:MUB

[175] Ogundiran Mathew Akinloye and Fawole Olatunde Olubanjo. Metal uptake and bioaccumulation potentials of Clarias buthupogon and Heterobranchus longifilis collected from Asa River, Ilorin, Nigeria. Fisheries and Aquaculture Journal, 8(3):1–7, ???? 2017. CODEN FAJIAO.

ISSN 2150-3508. URL https://www.longdom.org/open-access/metal-
uptake-and-bioaccumulation-potentials-of-clarias-buthupogon-
andheterobranchus-longifilis-collected-from-asa-river-ilorin-
nig-2150-3508-1000216.pdf.

Alam:2017:SET

[176] Maksud Alam, Shyamal Kumar Paul, and Kongchain Marma. Study on existing technology and knowledge on aquaculture by fish farmers in Gomastapur Upazila of Chapai Nawabgonj District, Bangladesh. Fisheries and Aquaculture Journal, 8(3):1–8, ???? 2017. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/study-on-
existing-technology-and-knowledge-on-aquaculture-by-fishfarmers-
in-gomastapur-upazila-of-chapai-nawabgonj-district-bangla-2150-
3508-1000217.pdf.

Das:2017:HMF

[177] Pritika Rani Das, Kamal Hossain, Bhakta Supratim Sarker, Afroza Parvin, Saborni Swarna Das, Mohammad Moniruzzaman, and Badhan Saha. Heavy metals in farm sediments, feeds and bioaccumulation of some selected heavy metals in various tissues of farmed Pangasius hypophthalmus in Bangladesh. Fisheries and Aquaculture Journal, 8(3):1–8, ???? 2017. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/heavy-metals-
in-farm-sediments-feeds-and-bioaccumulation-of-someselected-
heavy-metals-in-various-tissues-of-farmed-pangasius-hypop-2150-
3508-1000218.pdf.

Ajibade:2017:PBN

[178] A. O. Ajibade, E. K. Ajani, and B. O. Omitoyin. The predatory behaviour of nymphs of dragonfly (Africocypha varicolor) on fry of African mud catfish (Clarias gariepinus) and control by skunk weed (Petivera alliacea) root-extract in aquaculture. Fisheries and Aquaculture Journal, 8(3):1–5, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL

https://www.longdom.org/open-access/the-predatory-behaviour-
of-nymphs-of-dragonfly-africocypha-varicoloron-fry-of-african-
mud-catfish-clarias-gariepinus-and-control-b-2150-3508-1000219.
pdf.

Pan:2017:PCR

[179] Zhonghua Pan, Xuehong Song, Xiaolong Hu, Renyu Xue, Guangli Cao, Mian Sahib Zar, Dhiraj Kumar, Yongjie Feng, Yuhong Wei, Weiye Zhang, Wei Zhang, and Chengliang Gong. Pathological changes and risk factors of hepatopancreas necrosis disease of mitten crab, Eriocheir sinensis. Fisheries and Aquaculture Journal, 8(3):1– 5, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL https:

//www.longdom.org/open-access/pathological-changes-and-risk-
factors-of-hepatopancreas-necrosisdisease-of-mitten-crab-eriocheir-
sinensis-2150-3508-1000220.pdf.

Olanrewaju:2017:LWR

[180] A. N. Olanrewaju, E. K. Ajani, O. K. K. Kareem, and O. Orisasona. Length–weight relationships and state of well-being of Parachanna obscura Günther 1861, in Eleyele Reservoir, Southwestern Nigeria. Fisheries and Aquaculture Journal, 8(3):1–8, ???? 2017. CODEN

FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/lengthweight-relationships-and-state-of-wellbeing-of-
parachannaobscura-gunther1861-in-eleyele-reservoir-southwestern-
nigeria-2150-3508-1000221.pdf.

Sheikh:2017:CST

[181] M. Sheikh, M. Y. Laghari, P. K. Lashari, A. R. Khooharo, and N. T. Narejo. Current status of three major carps (Labeo rohita, Cirrhinus mrigala and Catla catla) in the Downstream Indus River, Sindh. Fisheries and Aquaculture Journal, 8(3):1–3, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL https://www.

longdom.org/open-access/current-status-of-three-major-carps-
labeo-rohita-cirrhinus-mrigala-andcatla-catla-in-the-downstream-
indus-river-sindh-2150-3508-1000222.pdf.

Ramasamy:2017:MRB

[182] Mathialagan Ramasamy and Sivakumar Rajangam. Maturation and reproductive biology of reba carp Cirrhinus reba (Hamilton) in Lower Anicut Reservoir, Tamil Nadu, India. Fisheries and Aquaculture Journal, 8(3):1–11, ???? 2017. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/maturation-and-
reproductive-biology-of-reba-carp-cirrhinus-rebahamilton-in-
lower-anicut-reservoir-tamil-nadu-india-2150-3508-1000223.pdf.

Ramasamy:2017:AGD

[183] Mathialagan Ramasamy and Sivakumar Rajangam. Age, growth and diet of reba carp Cirrhinus reba (Hamilton 1822) in Lower Anicut reservoir, Tamil Nadu, India. Fisheries and Aquaculture Journal, 8(3):1–9, ???? 2017. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/age-growth-and-diet-
of-reba-carp-cirrhinus-reba-hamilton-1822-inlower-anicut-reservoir-
tamil-nadu-india-2150-3508-1000224.pdf.

Sandre:2017:GEM

[184] L. C. G. Sandre, H. Buzollo, L. M. Neira, T. M. T. Nascimento, R. K. Jomori, and D. J. Carneiro. Growth and energy metabolism of tambaqui (Colossoma macropomum) fed diets with different levels of carbohydrates and lipids. Fisheries and Aquaculture Journal, 8(3):1–7, ???? 2017. CODEN FAJIAO. ISSN 2150-

3508.  URL https://www.longdom.org/open-access/growth-and-
energy-metabolism-of-tambaqui-colossoma-macropomumfed-diets-
with-different-levels-of-carbohydrates-and-lipids-2150-3508-
1000225.pdf.

Kwikiriza:2017:CGP

[185] G. Kwikiriza, A. Barekye, A. R. Aheisibwe, E. Byakora, and P. D. Tibihika. Comparative growth performance and proximate nutrient composition of three local strains of Nile tilapia (Oreochromis niloticus L.) collected from different locations in Uganda. Fisheries and Aquaculture Journal, 8(3):1–5, ???? 2017. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/comparative-growth-
performance-and-proximate-nutrient-compositionof-three-local-

strains-of-nile-tilapia-oreochromis-niloticus-l-co-2150-3508-
1000226.pdf.

Qasimov:2017:CSS

[186] R. Y. U. Qasimov, U. F. Hashimova, C. H. A. Mamedov, A. A. Gaisina, G. R. Vagabova, and Arif Mekhtiev. Comparative studies of several physiologic and biochemical indexes of wild type and hatchery-bred sturgeons in the early ontogenesis. Fisheries and Aquaculture Journal, 8(3): 1–9, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL https:/

/www.longdom.org/open-access/comparative-studies-of-several-
physiologic-and-biochemical-indexes-ofwild-type-and-hatcherybred-
sturgeons-in-the-early-ontogenesis-2150-3508-1000227.pdf.

Fenchel:2017:FNM

[187] Tom Fenchel, Bo Barker Jørgensen, and Hans Ulrik Riisgård. Fake news mussel farming a “New Climate Bomb”. Fisheries and Aquaculture Journal, 8(4):1–2, ???? 2017. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/fake-news-
mussel-farming-a-new-climate-bomb-2150-3508-1000e127.pdf.

Okey:2017:LWR

[188] Irom Okey, B. O. Offem, and R. I. Keremah. Length–weight relationship, condition factor and gut content of Chrysichthys furcatus Günther, 1864 (Bagridae) from Cross River at Ahaha. Fisheries and Aquaculture Journal, 8(4):1–8, ???? 2017. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/lengthweight-
relationship-condition-factor-and-gut-content-ofchrysichthys-
furcatus-gunther-1864-bagridae-from-cross-river-atahaha-2150-
3508-1000228.pdf.

Eke:2017:FCF

[189] F. N. Eke, G. E. Odo, J. E. Agwu, and C. B. Anya. Fecundity and condition factor of Clarias anguillaris of Oguta Lake, Nigeria. Fisheries and Aquaculture Journal, 8(4):1–6, ???? 2017. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/fecundity-and-
condition-factor-of-clarias-anguillaris-of-oguta-lake-nigeria-
2150-3508-1000229.pdf.

Huong:2017:EDI

[190] Nguyen Van Huong, Tran Huu Cuong, Tran Thi Nang Thu, and Philippe Lebailly. Efficiency of different integrated agriculture aquaculture systems in the Red River Delta of Vietnam. Fisheries and Aquaculture Journal, 8(4):1–10, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL

https://www.longdom.org/open-access/efficiency-of-different-
integrated-agriculture-aquaculture-systems-in-thered-river-delta-
of-vietnam-2150-3508-1000230.pdf.

Nguyen:2017:CSS

[191] Khanh Q. Nguyen and Vang Y. Nguyena. Changing of sea surface temperature affects catch of Spanish mackerel Scomberomorus commerson in the set-net fishery. Fisheries and Aquaculture Journal, 8 (4):1–7, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL

https://www.longdom.org/open-access/changing-of-sea-surface-
temperature-affects-catch-of-spanish-mackerel-scomberomorus-
commerson-in-the-setnet-fishery-2150-3508-1000231.pdf.

Das:2017:PPW

[192] Saborni Swarna Das, Kamal Hossain, Golam Mustafa M, Afroza Parvin, Badhan Saha, Pritika Rani Das, and Mohammad Moniruzzaman. Physicochemical properties of water and heavy metals concentration of sediments, feeds and various farmed tilapia (Oreochoromis niloticus) in Bangladesh. Fisheries and Aquaculture Journal, 8(4):1–8, ???? 2017. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/physicochemical-properties-of-water-and-heavy-metals-
concentration-ofsediments-feeds-and-various-farmed-tilapia-oreochoromis-
nilot-2150-3508-1000232.pdf.

Amponsah:2017:PDB

[193] Samuel Kweku Konney Amponsah, Ali Abdulhakim, Patrick Ofori-Danson, and Kofi Ferni Anyan. Population dynamics of bigeye grunt, Brachydeuterus auritus (Valenciennes, 1831) in Ghana and management implications. Fisheries and Aquaculture Journal, 8(4):1–6, ???? 2017. CODEN FAJIAO. ISSN 2150-3508. URL https:

//www.longdom.org/open-access/population-dynamics-of-bigeye-
grunt-brachydeuterus-auritus-valenciennes-1831-in-ghana-and-
management-implications-2150-3508-1000233.pdf.

Diadhiou:2017:SBM

[194] Hamet Diaw Diadhiou, Mbaye Tine, Anis Diallo, and et Mor Sylla. A survey of the benthic microfauna of the marine ecosystem on IRIS-1 oil platform at the Cape Skirring in Casamance, Senegal. Fisheries and Aquaculture Journal, 8(4):1–7, ???? 2017. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/a-survey-of-
the-benthic-microfauna-of-the-marine-ecosystem-on-iris1-oil-
platform-at-the-cape-skirring-in-casamance-senegal-2150-3508-
1000234.pdf.

Arnaud:2017:DMT

[195] Deli Arnaud, Lekeufack Folefack Guy Benoit, and Fomena Abraham. Description of Myxidium tetraodoni sp. nov., Myxidium anisocapsularis sp. nov. and Myxobolus magai sp. nov. (Myxosporea: Bivalvulida) infecting some freshwater fishes in Cameroon (Central Africa). Fisheries and Aquaculture Journal, 8(4):1–10, ???? 2017. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/description-of-
myxidium-tetraodoni-sp-nov-myxidium-anisocapsularissp-nov-and-
myxobolus-magai-sp-nov-myxosporea-bivalvulida-infecti-2150-3508-
1000236.pdf.

Pramanik:2017:DRT

[196] M. M. H. Pramanik and M. M. Hasan. Dhonagoda River: Threats investigation of river and biodiversity for policy implementation. Fisheries and Aquaculture Journal, 8(4):1–11, ???? 2017. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/dhonagoda-
river-threats-investigation-of-river-and-biodiversity-forpolicy-
implementation-2150-3508-1000237.pdf.

Dauda:2018:ACA

[197] Akeem Babatunde Dauda, Ikhsan Natrah, Murni Karim, Mohd Salleh Kamarudin, and Armaya’u Hamisu Bichi. African catfish aquaculture in Malaysia and Nigeria: Status, trends and prospects. Fisheries and Aquaculture Journal, 9(1):1–5, ???? 2018. CO-DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/african-catfish-aquaculture-in-malaysia-and-
nigeria-status-trends-and-prospects-2150-3508-1000237.pdf.

Tesfaye:2018:PSG

[198] Shimels Tesfaye, Misaw Kasye, Mersha Chane, Baseazinew Bogale, and Zewudie Abebe agere. Preliminary survey of Gram-negative bacterial pathogens from commonly caught fish species (Oreochromis niloticus, Cyprinus carpio and Clarias gariepinus) in Lake Hayiq, Ethiopia. Fisheries and Aquaculture Journal, 9(1):1–7, ???? 2018. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
preliminary-survey-of-gramnegative-bacterial-pathogens-from-
commonly-caught-fish-species-oreochromis-niloticus-cyprinus-
carpio-and-2150-3508-1000238.pdf.

Asnake:2018:FFR

[199] Wubshet Asnake and Minwyelet Mingist. Freshwater fisheries resource potential estimation: The case of Lake Ardibo, Northern Ethiopia. Fisheries and Aquaculture Journal, 9(1):1–6, ???? 2018. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
freshwater-fisheries-resource-potential-estimation-the-case-
of-lake-ardibo-northern-ethiopia-2150-3508-1000239.pdf.

Zeng:2018:ATZ

[200] Ling Zeng, Lan Huang, Ming Zhao, Sheng Liu, Zhengjian He, Jupan Feng, Chuanjie Qin, and Dengyue Yuan. Acute toxicity of zinc sulfate heptahydrate (ZnSO_{4} * 7 H_{2}O) and copper (II) sulfate pentahydrate (CuSO_{4} * 5 H_{2}O) on freshwater fish, Percocypris pingi. Fisheries and Aquaculture Journal, 9(1):1–5, ???? 2018. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/acute-toxicity-
of-zinc-sulfate-heptahydrate-znso47h2o-and-copperii-sulfate-
pentahydrate-cuso45h2o-on-freshwater-fish-percocyprispi-2150-
3508-1000240.pdf.

Aynalem:2018:DRA

[201] Yibeletal Aynalem, Minwyelet Mingist, and Addis Getu. Diversity, relative abundance, species composition and some biological aspects of fishes in Gilgel Abay and Andassa Rivers, Blue Nile Basin, Ethiopia. Fisheries and Aquaculture Journal, 9(1):1–13, ???? 2018. CODEN

FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/diversity-relative-abundance-species-composition-and-
some-biologicalaspects-of-fishes-in-gilgel-abay-and-andassa-
rivers-blue-nile-2150-3508-1000241.pdf.

Hiko:2018:HGN

[202] Adem Hiko, Kanani Tasisa, and Getahun E. Agga. Helminthiasis and Gram negative enteric bacteria in freshwater fish from selected lakes of Haramaya District, Ethiopia. Fisheries and Aquaculture Journal, 9(2): 1–7, ???? 2018. CODEN FAJIAO. ISSN 2150-3508. URL https://

www.longdom.org/open-access/helminthiasis-and-gram-negative-
enteric-bacteria-in-freshwater-fishfrom-selected-lakes-of-haramaya-
district-ethiopia-2150-3508-1000242.pdf.

Wang:2018:PSC

[203] He Wang, Sebsibe Amesa, Megerssa Endebu, Girma Tirfessa, Zou Zhong-Yi, and Wu Zhi-Gang. Preliminary study on construction of three-in-one fishpond and its effect on aquaculture in Ethiopia. Fisheries and Aquaculture Journal, 9(2):1–6, ???? 2018. CO-DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/preliminary-study-on-construction-of-threeinone-

fishpond-and-itseffect-on-aquaculture-in-ethiopia-2150-3508-
1000243.pdf.

Yongo:2018:SBS

[204] Edwine Yongo and Alex Wairimu. Studies on the biology of Synodontis victoriae in the Nyanza Gulf of Lake Victoria, Kenya. Fisheries and Aquaculture Journal, 9(2):1–3, ???? 2018. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/studies-on-the-
biology-of-synodontis-victoriae-in-the-nyanza-gulf-of-lakevictoria-
kenya-2150-3508-1000244.pdf.

P:2018:GSB

[205] Karthiga Priya P, Sundaramoorthy B, Neethiselvan N, Sukumar D, and Masilan K. Gillnet selectivity for big eye barracuda, Sphyraena forsteri (Cuvier, 1829) in Thoothukudi Waters, Southeast Coast of India. Fisheries and Aquaculture Journal, 9(2):1–5, ???? 2018. CODEN FAJIAO. ISSN 2150-3508. URL https://

www.longdom.org/open-access/gillnet-selectivity-for-big-eye-
barracuda-sphyraena-forsteri-cuvier-1829-in-thoothukudi-waters-
southeast-coast-of-india-2150-3508-1000245.pdf.

Mohammed:2018:EMO

[206] Suleiman Adamu Mohammed, Orire Abdullahi Muhammad, and Sadiku Suleiman Omeiza Eku. Effects of Moringa oleifera leaves, bark stem of Lannea barteri and antibiotic (oxytetracycline) on haematological parameters of Clarias gariepinus fingerlings. Fisheries and Aquaculture Journal, 9(2):1–5, ???? 2018. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/effects-of-moringa-
oleifera-leaves-bark-stem-of-lannea-barteri-and-antibiotic-oxytetracycline-
on-haematological-parameters-of-clar-2150-3508-1000246.pdf.

Awodiran:2018:GDC

[207] Michael O. Awodiran and Olumide Afolabi. Genetic diversity in cultured and wild population of Clarias gariepinus (Burchell, 1822) in Nigeria using Random Amplified Polymorphic DNA (RAPD) and microsatellite DNA. Fisheries and Aquaculture Journal, 9(2):1–6, ???? 2018. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/genetic-diversity-in-cultured-and-wild-population-of-
clarias-gariepinusburchell-1822-in-nigeria-using-random-amplified-
polymorphic-2150-3508-1000247.pdf.

Dyachuk:2018:EMC

[208] Vyacheslav Dyachuk. Extracellular matrix components in bivalvia: Shell and ECM components in developmental and adult tissues. Fisheries and Aquaculture Journal, 9(2):1–6, ???? 2018. CO-DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/extracellular-matrix-components-in-bivalvia-
shell-and-ecm-componentsin-developmental-and-adult-tissues-2150-
3508-1000248.pdf.

Alemayehu:2018:RFF

[209] Tewodros Abate Alemayehu, Akwake Geremew, and Abebe Getahun. The role of functional feed additives in tilapia nutrition. Fisheries and Aquaculture Journal, 9(2):1–6, ???? 2018. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/the-role-
of-functional-feed-additives-in-tilapia-nutrition-2150-3508-
1000249.pdf.

Keyombe:2018:LFI

[210] James Last A. Keyombe, Ruth M. Lewo, Edna Waithaka, Alice Mutie, and Priscilla Boera. Has the latest fish introduction in Lake Naivasha improved income of fishermen? The economics of Nile tilapia (Oreochromis niloticus) in Lake Naivasha. Fisheries and Aquaculture Journal, 9(3):1–3, ???? 2018. CODEN FAJIAO. ISSN 2150-3508. URL https://www.

longdom.org/open-access/has-the-latest-fish-introduction-in-
lake-naivasha-improved-income-offishermen-the-economics-of-nile-
tilapia-oreochromis-niloticus-2150-3508-1000250.pdf.

Abbassy:2018:WFQ

[211] Moustafa Mohamed Saleh Abbassy. Water and fish quality of aquaculture pond adjacent to intensive pesticides application agro-system. Fisheries and Aquaculture Journal, 9(3):1–9, ???? 2018. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/water-and-
fish-quality-of-aquaculture-pond-adjacent-to-intensivepesticides-
application-agrosystem-2150-3508-1000251.pdf.

Chen:2018:SRD

[212] Wang Ching Chen, Yung Song Chen, and Hong Nong Chou. Strategic review on developing salmon aquaculture with deep ocean water: a case study of Nan’ao Deep Ocean Water Park in Yilan County, Taiwan. Fisheries and Aquaculture Journal, 9(3):1–7, ???? 2018. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
strategic-review-on-developing-salmon-aquaculture-with-deep-

ocean-water-a-case-study-of-nan8217ao-deep-ocean-water-park-
in-yilan-c-2150-3508-1000252.pdf.

Fouda:2018:WMS

[213] Tarek Fouda. Waste management for smoking salmon by-products to extract omega-3 fish oil. Fisheries and Aquaculture Journal, 9(3):1–3, ???? 2018. CODEN FAJIAO. ISSN 2150-3508. URL https://www.

longdom.org/open-access/waste-management-for-smoking-salmon-
byproducts-to-extract-omega3-fish-oil-2150-3508-1000253.pdf.

Masud:2018:FFH

[214] Shadab Masud and K. P. Singh. Food and feeding habit of Gonialosa manmina (Ham.) from the River Yamuna, Allahabad, India. Fisheries and Aquaculture Journal, 9(3):1–5, ???? 2018. CODEN

FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/food-and-feeding-habit-of-gonialosa-manmina-ham-from-
the-river-yamuna-allahabad-india-2150-3508-1000254.pdf.

Tiwo:2018:EIS

[215] Cristelle Tsapla Tiwo, M. V. Chandra, H. M. Womeni, N. F. Zambou, S. Ndomou, F. Tchoumbougnang, D. A. Dzoukoua, B. B. Nayak, Anandan R, and Pankaj K. Effect of ice storage on the textural and rheological properties of proteins from freshwater fish, Cyprinus carpio (common carp). Fisheries and Aquaculture Journal, 9(3):1–10, ???? 2018. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/effect-of-ice-
storage-on-the-textural-and-rheological-properties-of-proteins-
from-freshwater-fish-cyprinus-carpio-common-carp-2150-3508-1000255.
pdf.

Pall:2018:EMO

[216] Jewel Chandra Pall, Shuvagato Mondal, Priyanka Rani Majumdar, and Md. Abul Hossain. Effect of multi-ownership on pond aquaculture production in Bhola District, Bangladesh. Fisheries and Aquaculture Journal, 9(3):1–7, ???? 2018. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/effect-of-
multiownership-on-pond-aquaculture-production-in-bhola-district-
bangladesh-2150-3508-1000256.pdf.

Samadan:2018:UMS

[217] Gamal M. Samadan, Rustadi, Djumanto, and Murwantoko. Utilization of marginal sand land for culture of white leg shrimp (Litopenaeus vannamei) with different stocking density in Coastal Purworejo Regency, Central Java, Indonesia. Fisheries and Aquaculture Journal, 9(3):1–8, ???? 2018. CODEN FAJIAO. ISSN 2150-3508. URL

https://www.longdom.org/open-access/utilization-of-marginal-
sand-land-for-culture-of-white-leg-shrimp-litopenaeus-vannamei-
with-different-stocking-density-in-coastal-2150-3508-1000257.
pdf.

Atujona:2018:MRV

[218] Denicia Atujona, Shuanghu Cai, and Eric Amenyogbe. Mini review on Vibrio infection — a case study on Vibrio harveyi clade. Fisheries and Aquaculture Journal, 9(4):1–4, ???? 2018. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/mini-review-on-
vibrio-infectiona-case-study-on-emvibrio-harveyiem-clade.pdf.

Duodu:2018:GCM

[219] Collins Prah Duodu, Regina Edziyie, Nelson Winston Agbo, Daniel Adjei-Boateng, and Peter Vilhelm Skov. Groundnut and cottonseed meals in tilapia juvenile diets: Effects on apparent nutrient digestibility, short-term growth performance, feed utilization and carcass composition. Fisheries and Aquaculture Journal, 9(4):1–8, ???? 2018. CODEN FAJIAO. ISSN 2150-3508. URL https:/

/www.longdom.org/open-access/groundnut-and-cottonseed-meals-
in-tilapia-juvenile-diets-effects-on-apparent-nutrient-digestibility-
shortterm-growth-performance-f-2150-3508-1000259.pdf.

Gabal:2019:OSP

[220] Ashgan A. Abou Gabal, Ahemd E. M. Khaled, Heba Saad El-Sayed, Haiam M. Aboul-Ela, Ola Kh. Shalaby, Asmaa A. Khaled, and Laila A. Mohamed. Optimization of Spirulina platensis biomass and evaluation of its protective effect against chromosomal aberrations of bone marrow cells. Fisheries and Aquaculture Journal, 10(1):1–6, ???? 2019. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/optimization-of-emspirulina-platensisem-biomass-
and-evaluation-of-its-protective-effect-against-chromosomal-
aberrations-.pdf.

Sakyi:2019:SCA

[221] Rhoda Lims Sakyi, P. K. Ofori-Danson, Sam Addo, and E. Nyarko. Stomach content analysis and concentrations of chemical pollutants in the Clymene dolphin (Stenella clymene, Gray 1846) from the coastal waters of Ghana. Fisheries and Aquaculture Journal, 10(1):1–7, ???? 2019. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/stomach-content-analysis-and-concentrations-

of-chemical-pollutants-in-the-clymene-dolphin-emstenella-clymeneem-
gray-1846.pdf.

Henry:2019:ODP

[222] Chadwick Bironga Henry, Christopher Aura Mulanda, and James Njiru. Ovarian development of the penaeid shrimp Penaeus indicus (Decapoda): a case for the Indian Ocean coastal waters of Kilifi Creek, Kenya. Fisheries and Aquaculture Journal, 10(1):1–9, ???? 2019. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
ovarian-development-of-the-penaeid-shrimp-ipenaeus-indicusi-
decapoda-a-case-for-the-indian-ocean-coastal-waters-of-kilif.
pdf.

Mehanna:2019:STR

[223] Sahar F. Mehanna, Mohammed G. Desouky, and Ahmed F. Makkey. Some targeted reference points for thin lip grey mullet Liza ramada management in Bardawil Lagoon, North Sinai, Egypt. Fisheries and Aquaculture Journal, 10(1):1–6, ???? 2019. CODEN FAJIAO. ISSN 2150-3508. URL

https://www.longdom.org/open-access/some-targeted-reference-
points-for-thin-lip-grey-mullet-liza-ramada-management-in-bardawil-
lagoon-north-sinai-egypt.pdf.

Genc:2019:EDS

[224] Murat Genc, Gonca Alak, Muhammed Atamanalp, Nilufer Sabuncuoglu, Esat Mahmut Kocaman, Ziya Gokalp Ceylan, and Omer Coban. The effects of different stunning techniques on meat quality of brown trout (Salmo trutta fario). Fisheries and Aquaculture Journal, 10(1):1–7, ???? 2019. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/the-effects-of-
different-stunning-techniques-on-meat-quality-of-brown-trout-
emsalmo-trutta-farioem.pdf.

Mohamed:2019:LEA

[225] Amal S. Mohamed, Nahed S. Gad, and Mohamed A. El Desoky. Liver enzyme activity of Tilapia zillii and Mugil capito collected seasonally from Qarun Lake, Egypt. Fisheries and Aquaculture Journal, 10(2):1–4, ???? 2019. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/liver-enzyme-activity-of-emtilapia-zilliiem-
and-emmugil-capitoem-collected-seasonally-from-qarun-lake-egypt.
pdf.

Kerie:2019:ESI

[226] Yoseph Kerie, Anwar Nuru, and Takele Abayneh. Edwardsiella species infection in fish population and its status in Ethiopia. Fisheries and Aquaculture Journal, 10(2):1–6, ???? 2019. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/edwardsiella-
species-infection-in-fish-population-and-its-status-in-ethiopia.
pdf.

Agboola:2019:GPA

[227] E. O. Agboola, O. A. Owoeye, J. K. Balogun, J. Auta, and S. A. Abdullahi. Growth performance of the African catfish, Clarias gariepinus (Burchell), fed varying inclusion levels of castor seed (Ricinus communis L.). Fisheries and Aquaculture Journal, 10(2):1–7, ???? 2019. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/growth-performance-of-the-african-catfish-clarias-gariepinus-
burchell-fed-varying-inclusion-levels-of-castor-seed-ricinu.
pdf.

Hammed:2019:CMQ

[228] Ayofe M. Hammed, Fatai G. Owodeinde, Albert O. Amosu, Folalu A. Awe, and Hakeem A. Fashina-Bombata. Chemical and microbial quality assessments of some economic important artisanal fresh water fish species (Tilapia zilli, Clarias gariepinus, Chrysichthys nigrodigitatus and S. melanotheron) in selected coastal markets of Lagos State, Nigeria. Fisheries and Aquaculture Journal, 10(2):1–6, ???? 2019. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
chemical-and-microbial-quality-assessments-of-some-economic-
important-artisanal-fresh-water-fish-species-tilapia-zilli-c.
pdf.

Elijah:2019:SPP

[229] S. Okwuonu Elijah, E. Odo Gregory, S. O. Eme, and O. Okoye Charles. Studies on the physicochemical properties of Ebonyi River in Ebonyi State, Nigeria and on the abundance of benthic fauna during the rainy season. Fisheries and Aquaculture Journal, 10(3):1– 3, ???? 2019. CODEN FAJIAO. ISSN 2150-3508. URL https:/

/www.longdom.org/open-access/studies-on-the-physicochemical-
properties-of-ebonyi-river-in-ebonyi-state-nigeria-and-on-the-
abundance-of-benthic-fauna-.pdf.

Pejuan:2019:OHT

[230] Wolfgang Pejuán, Valeria Criollo, and Patricio Enrique Paz. Optimum harvest time of the Nile tilapia (Oreochromis niloticus) in Honduras: a two step OLS procedure to obtain the growth function. Fisheries and Aquaculture Journal, 10(3):1–15, ???? 2019. CODEN

FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/optimum-harvest-time-of-the-nile-tilapia-oreochromis-
niloticus-in-honduras-a-two-step-ols-procedure-to-obtain-the-
growth.pdf.

Zhao:2019:EDY

[231] Xing Zhao, Ying Sun, Haiyan Liu, and Zhencai Yang. Effects of dietary yeast extract levels on growth performance, digestibility and antioxidant capacity of the Taiwan loach. Fisheries and Aquaculture Journal, 10(3):1–6, ???? 2019. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/effects-of-dietary-
yeast-extract-levels-on-growth-performance-digestibility-and-
antioxidant-capacity-of-the-taiwan-loach.pdf.

Bjarnason:2019:SSF

[232] Arnar Bjarnason and Soffia Karen Magnsdottir. The salmon sea fish farming industry in Iceland: a review. Fisheries and Aquaculture Journal, 10(4):1–5, ???? 2019. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/the-salmon-sea-fish-
farming-industry-in-iceland-a-review.pdf.

Ramirez-Villalobos:2020:AES

[233] Hugo Gael Ramírez-Villalobos and Ernesto A. Chávez. Assessing the exploitation strategies of the shrimp fishery in the Gulf of California. Fisheries and Aquaculture Journal, 11(1):1–9, ???? 2020. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
assessing-the-exploitation-strategies-of-the-shrimp-fishery-
in-the-gulf-of-california.pdf.

Yu:2020:PGP

[234] Xinxiu Yu, Tormod Ådnøy, Zhenming Lv, Changwen Wu, and Hans Magnus Gjøen. Phenotypic and genetic parameter estimation for growth traits in juvenile large yellow croaker (Larimichthys crocea). Fisheries and Aquaculture Journal, 11(1):1–5, ???? 2020. CO-DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/phenotypic-and-genetic-parameter-estimation-for-
growth-traits-in-juvenile-large-yellow-croaker-larimichthys-
crocea.pdf.

Pandiri:2020:EFA

[235] Karunasree Pandiri. Editorial — Fisheries and Aquaculture Journal (2150-3508). Fisheries and Aquaculture Journal, 11(2):1–2, ???? 2020. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/editorial-fisheries-and-aquaculture-journal-21503508.
pdf.

VonEschen:2020:EAA

[236] Aaron J. Von Eschen, Michael L. Brown, and Kurt Rosentrater. Effects of amino acid supplements in plant-based yellow perch diets. Fisheries and Aquaculture Journal, 11(2):1–8, ???? 2020. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/effects-of-
amino-acid-supplements-in-plantbased-yellow-perch-diets.pdf.

Benoit:2020:DTN

[237] Lekeufack Folefack Guy Benoît, Fongang Tsekeng Charlotte Vanessa, and Fomena Abraham. Description of three new species of Myxosporidia (Cnidaria: Myxobolidae) parasites of Paramormyrops kingsleyae Günther, 1896 (Mormyridae) in the Nyong basin in Cameroon. Fisheries and Aquaculture Journal, 11(2):1–8, ???? 2020. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/description-of-
three-new-species-of-myxosporidia-cnidaria-myxobolidae-parasites-
of-paramormyrops-kingsleyae-gnther-1896-.pdf.

P:2020:MMF

[238] Karunasree P. Mangrove macro flora and faunal diversity: Editorial. Fisheries and Aquaculture Journal, 11(3):1, ???? 2020. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
mangrove-macro-flora-and-faunal-diversity-editorial.pdf.

Villarino:2020:FFG

[239] Resti Tito Villarino. Formulated feeds for genetically improved farmed tilapia (GIFT). Fisheries and Aquaculture Journal, 11(3):1–6, ???? 2020. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/formulated-feeds-for-genetically-improved-farmed-
tilapia-gift.pdf.

Georges:2020:EWP

[240] Fonkwa Georges, Kouam K. Marc, Tchuinkam Timoléon, Tomedi Eyango Minette, and Tchoumboue. Effect of water physico-chemical characteristics on Myxobolus tilapiae, a myxosporean parasite of the fish Oreochromis niloticus at MAPE Dam in Adamawa Region of Cameroon. Fisheries and Aquaculture Journal, 11(3):1–5, ???? 2020. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/effect-of-
water-physicochemical-characteristics-on-myxobolus-tilapiae-
a-myxosporean-parasite-of-the-fish-oreochromis-nil.pdf.

Uddin:2020:SEC

[241] Md Kamal Uddin, Md Robiul Hasan, Shyamal Kumar Paul, and Tasnim Sultana. Socio-economic condition and livelihood status of the fisherman community at Muradnagar Upazila in Cumilla. Fisheries and Aquaculture Journal, 11(3):1–5, ???? 2020. CODEN

FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/socioeconomic-condition-and-livelihood-status-of-the-
fisherman-community-at-muradnagar-upazila-in-cumilla.pdf.

Keyombe:2020:SAF

[242] James Last A. Keyombe, Obiero Kevin, Edna Waithaka, Outa Nicholas, Donde Oscar, and Domitila N. Kyule. Some aspects of fish growth in redbelly tilapia (Coptodon zilli) and largemouth bass (Micropterus salmoides) in Lake Naivasha, Kenya. Fisheries and Aquaculture Journal, 11(3):1–5, ???? 2020. CODEN FAJIAO. ISSN 2150-3508. URL https:/

/www.longdom.org/open-access/some-aspects-of-fish-growth-in-
redbelly-tilapia-coptodon-zilli-and-largemouth-bass-micropterus-
salmoides-in-lake-naivash.pdf.

Ehsan:2021:AMA

[243] Ali Carl Ehsan. Aquaculture market analysis. Fisheries and Aquaculture Journal, 12(1):1, ???? 2021. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/aquaculture-
market-analysis.pdf.

Ehsan:2021:LFI

[244] Ali Carl Ehsan. Limnology and fisheries of inland waters. Fisheries and Aquaculture Journal, 12(1):1, ???? 2021. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/limnology-and-
fisheries-of-inland-waters.pdf.

Bjarnason:2021:CTL

[245] Arnar Bjarnason and Soffia Karen Magnusdottir. A C-type lectin with antibacterial activity in weather loach. Fisheries and Aquaculture Journal, 12(2):1–6, ???? 2021. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/a-ctype-lectin-with-
antibacterial-activity-in-weather-loach.pdf.

Karen:2021:MDA

[246] Soffia Karen. Macrophyte diversity alters invertebrate community and fish diet. Fisheries and Aquaculture Journal, 12(2):1–6, ???? 2021. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/macrophyte-diversity-alters-invertebrate-community-and-
fish-diet.pdf.

J:2021:TLS

[247] Ander J. Testicular leiomyoma and spermatogenia failure syndrome. Fisheries and Aquaculture Journal, 12(2):1, ???? 2021. CODEN FAJIAO.

ISSN 2150-3508. URL https://www.longdom.org/open-access/study-
of-size-and-climatic-requirements-of-freshwater-phytoplankton.
pdf.

P:2021:PEF

[248] Kishan P. Probiotic effect on feeding technology and their use in aquaculture. Fisheries and Aquaculture Journal, 12(2):1, ???? 2021. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
risk-management-strategies-in-pond-auaculture.pdf.

J:2021:LVF

[249] Ander J. Litopenaeus vannamei feeding activity and development on soybean based diets with added feeding effectors. Fisheries and Aquaculture Journal, 12(2):1, ???? 2021. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/testicular-
leiomyoma-and-spermatogenia-failure-syndrome.pdf.

Desalegn:2021:FRC

[250] Tamenut Desalegn and Takele Shitaw. Fishery resources, conservation challenges and management strategies in Ethiopia. Fisheries and Aquaculture Journal, 12(3):1–5, ???? 2021. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/fishery-
resources-conservation-challenges-and-management-strategies-
in-ethiopia.pdf.

Jayawardena:2021:PAY

[251] A. W. Jayawardena. Prestigious award for young researchers at Oceanography 2020. Fisheries and Aquaculture Journal, 12(3):1, ???? 2021. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/the-effects-of-ocean-acidification-and-warming-on-marine-
fish-health.pdf.

Thompson:2021:DCC

[252] Olaniran A. Thompson, S. F. Arifalo, and A. A. Atejioye. Determinants of climate change risk management strategies among the aquaculture fish farmers in Nigeria using multinomial logit model. Fisheries and Aquaculture Journal, 12(3):1–3, ???? 2021. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/determinants-
of-climate-change-risk-management-strategies-among-the-aquaculture-
fish-farmers-in-nigeria-using-multinomia.pdf.

Agbugui:2021:IFF

[253] Marian Agbugui, L. I. Emawunegbe, and A. C. Yaro. Indigenous fish feed resources in Nigeria. Fisheries and Aquaculture Journal, 12(3): 1–2, ???? 2021. CODEN FAJIAO. ISSN 2150-3508. URL https:/

/www.longdom.org/open-access/indigenous-fish-feed-resources-
in-nigeria.pdf.

Hosain:2021:PSF

[254] Mohammad Aslam Hosain, Kalim Ullah, Mohammad Abdullah Al Sayam, Kazi Mohiuddin, and Emon Rahman. Present status and future direction of Bangladeshi shrimp resources. Fisheries and Aquaculture Journal, 12(3):1–6, ???? 2021. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/present-status-
and-future-direction-of-bangladeshi-shrimp-resources.pdf.

Labelle:2021:MLE

[255] Marc Labelle, Richard Bussanich, and Ryan Benson. Maximum likelihood estimation of Kokanee and sockeye salmon spawners in a stream using visual survey data, 2003–2017. Fisheries and Aquaculture Journal, 12 (4):1–9, ???? 2021. CODEN FAJIAO. ISSN 2150-3508. URL https:

//www.longdom.org/open-access/maximum-likelihood-estimation-
of-kokanee-and-sockeye-salmon-spawners-in-a-stream-using-visual-
survey-data-20032017.pdf.

Jayawardena:2021:AA

[256] A. W. Jayawardena. Award of aquaculture. Fisheries and Aquaculture Journal, 12(4):1, ???? 2021. CODEN FAJIAO. ISSN 2150-3508. URL

https://www.longdom.org/open-access/award-of-aquaculture.pdf.

Nassiri:2021:ICA

[257] Rune Nassiri. 9th International Conference on Aquaculture & Fisheries schedule on October 07–08, 2021 at Vienna, Austria. Fisheries and Aquaculture Journal, 12(4):1–0, ???? 2021. CODEN FAJIAO.

ISSN 2150-3508. URL https://www.longdom.org/open-access/9th-
international-conference-on-aquaculture--fisheries-schedule-
on-october-0708-2021-at-vienna-austria.pdf.

Oluwafemi:2021:SDA

[258] Funmilola Oluwafemi. Save the date Aquaculture 2021 will be arriving! Fisheries and Aquaculture Journal, 12(4):1, ???? 2021. CODEN FAJIAO.

ISSN 2150-3508. URL https://www.longdom.org/open-access/save-
the-date-aquaculture-2021-will-be-arriving.pdf.

Begum:2021:MAF

[259] Alsolami Begum. Market analysis for fisheries in Europe 2020. Fisheries and Aquaculture Journal, 12(4):1, ???? 2021. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
market-analysis-for-fisheries-in-europe-2020.pdf.

Tareke:2021:DSJ

[260] Abebe Tesfaye Tareke, Abebe Getahun Gubale, and Tadesse Fetahi Hailu. Dietary shifts in juvenile and adult Nile tilapia, Oreochromis niloticus (L.) (Pisces: Cichlidae) in Lake Ziway, Ethiopia. Fisheries and Aquaculture Journal, 12(5):1–6, ???? 2021. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/dietary-shifts-in-
juvenile-and-adult-nile-tilapia-oreochromis-niloticus-lpisces-
cichlidae-in-lake-ziway-ethiopia.pdf.

Shitaw:2021:ESC

[261] Takele Shitaw, Lammessa Berisa, and Berhan Asmamaw. Effect of season on the concentration of nutrients in the three highland lakes of Ethiopia. Fisheries and Aquaculture Journal, 12(5): 1–8, ???? 2021. CODEN FAJIAO. ISSN 2150-3508. URL

https://www.longdom.org/open-access/the-effect-of-season-on-
the-concentration-of-nutrients-in-the-three-highland-lakes-of-
ethiopia.pdf.

Sam:2021:AMF

[262] Louis Sam. Advances in marine farming: a commentary. Fisheries and Aquaculture Journal, 12(5):1, ???? 2021. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/advances-
in-marine-farming-a-commentary.pdf.

Ehsan:2021:NCF

[263] Ali Carl Ehsan. A note on catfish farming. Fisheries and Aquaculture Journal, 12(5):1, ???? 2021. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/a-note-on-
catfish-farming.pdf.

Sam:2021:IRS

[264] Louis Sam. Insights of recreational (sport) fishing. Fisheries and Aquaculture Journal, 12(5):1, ???? 2021. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/insights-
of-recreational-sport-fishing.pdf.

Ehsan:2021:TA

[265] Ali Carl Ehsan. Tropical aquaculture. Fisheries and Aquaculture Journal, 12(5):1, ???? 2021. CODEN FAJIAO. ISSN 2150-3508. URL https://

www.longdom.org/open-access/tropical-aquaculture.pdf.

Stryjeckin:2021:MIA

[266] Rodgers M. Stryjeckin. Market influences and apparatus to examine fisheries. Fisheries and Aquaculture Journal, 12(6):1, ???? 2021. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
market-influences-and-apparatus-to-examine-fisheries.pdf.

Oliver:2021:MCC

[267] Daniel Oliver. Major constrains in costal mariculture. Fisheries and Aquaculture Journal, 12(6):1–2, ???? 2021. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/major-constrains-in-
costal-mariculture.pdf.

Stryjeckin:2021:PGC

[268] Rodgers M. Stryjeckin. Population genetics in conservation and aquaculture. Fisheries and Aquaculture Journal, 12(6):2, ???? 2021. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
population-genetics-in-conservation-and-aquaculture.pdf.

Shah:2021:TLF

[269] Md Imran Shah. Transportation of live fish seed. Fisheries and Aquaculture Journal, 12(6):1–3, ???? 2021. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/transportation-
of-live-fish-seed.pdf.

Oliver:2021:CPG

[270] Daniel Oliver. A commentary on population genetics of hatchery and wild fish. Fisheries and Aquaculture Journal, 12(6):1–2, ???? 2021.

CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/

open-access/a-commentary-on-population-genetics-of-hatchery-
and-wild-fish.pdf.

Kela:2021:EPC

[271] Esther Kela, Amos O. Sogbesan, and Umar Buba Wakil. Evaluation of phytochemical composition of ginger extracts. Fisheries and Aquaculture Journal, 12(7):1–5, ???? 2021. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/evaluation-of-
phytochemical-composition-of-ginger-extracts.pdf.

Oliver:2021:ENC

[272] Daniel Oliver. An editorial note on commercial fish feed. Fisheries and Aquaculture Journal, 12(7):1, ???? 2021. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/an-editorial-
note-on-commercial-fish-feed.pdf.

Oliver:2021:BND

[273] Daniel Oliver. A brief note on deep-sea fish. Fisheries and Aquaculture Journal, 12(7):1, ???? 2021. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/a-brief-note-
on-deepsea-fish.pdf.

Stryjeckin:2021:BNR

[274] Rodgers M. Stryjeckin. Brief note on recirculating aquaculture systems. Fisheries and Aquaculture Journal, 12(7):1, ???? 2021. CODEN FAJIAO.

ISSN 2150-3508. URL https://www.longdom.org/open-access/brief-
note-on-recirculating-aquaculture-systems.pdf.

Ehsan:2021:BNI

[275] Ali Carl Ehsan. Brief note on insights on environmental impact of fishing. Fisheries and Aquaculture Journal, 12(7):1, ???? 2021. CODEN FAJIAO.

ISSN 2150-3508. URL https://www.longdom.org/open-access/brief-
note-on-insights-on-environmental-impact-of-fishing.pdf.

Oliver:2021:BNM

[276] Daniel Oliver. Brief note on marine habitats. Fisheries and Aquaculture Journal, 12(7):1, ???? 2021. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/brief-note-on-
marine-habitats.pdf.

Bezabih:2022:BNF

[277] Kidanie Bezabih. Brief note on fish-handler’s disease. Fisheries and Aquaculture Journal, 13(1):1–2, ???? 2022. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/brief-note-on-
fishhandlers-disease.pdf.

Shah:2022:CPP

[278] Imran Shah, Niraj Kumar, and Shivani Kumari. Construction and pond preparation in aquaculture. Fisheries and Aquaculture Journal, 13(1):1–4, ???? 2022. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/construction-
and-pond-preparation-in-aquaculture.pdf.

Chan:2022:MRT

[279] Yen Yew Chan, Siou Ning Aileen See, and Subha Bhassu. A mini-review on technological advancement of diagnostic and therapeutic measures in disease control used in aquaculture. Fisheries and Aquaculture Journal, 13(1):1–4, ???? 2022. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
a-minireview-on-technological-advancement-of-diagnostic-and-
therapeutic-measures-in-disease-control-used-in-aquaculture.
pdf.

Bezabih:2022:CMA

[280] Kidanie Bezabih. Commentary on microorganisms in aquaculture. Fisheries and Aquaculture Journal, 13(1):1–2, ???? 2022. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
commentary-on-microorganisms-in-aquaculture.pdf.

Dabi:2022:AWQ

[281] Shambel Boki Dabi. Assessment of water quality parameters for aquaculture uses: The case of Guder River Main Tributaries of Nile, Ethiopia. Fisheries and Aquaculture Journal, 13(1):1–5, ???? 2022. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
assessment-of-water-quality-parameters-for-aquaculture-uses-
the-case-of-guder-river-main-tributaries-of-nile-ethiopia.pdf.

Olaniyi:2022:VCC

[282] Oluwafemi Zaccheaus Olaniyi, Ayanboye Abolupe Oluyemisi, and Oluwafemi Benedicta Motunrayo. Value chain of catfish products in Ibadan Metropolis, Oyo State, Nigeria. Fisheries and Aquaculture Journal, 13(1):1–6, ???? 2022. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/value-chain-of-
catfish-products-in-ibadan-metropolis-oyo-state-nigeria.pdf.

Presas-Basalo:2022:PHF

[283] Francisco Xose Presas-Basalo. Potassium homeostasis and fish welfare in coupled aquaponic systems. Fisheries and Aquaculture Journal, 13(2): 1–5, ???? 2022. CODEN FAJIAO. ISSN 2150-3508. URL https:/

/www.longdom.org/open-access/potassium-homeostasis-and-fish-
welfare-in-coupled-aquaponic-systems.pdf.

Moslemi:2022:STF

[284] Mehran Moslemi, Mohamadreza Ahmadi, and Ali Barzegar. Investigating feeding regimen of brown trout (Salmo trutta fario) in Tonekabon River, Northern Iran. Fisheries and Aquaculture Journal, 13(2): 1–3, ???? 2022. CODEN FAJIAO. ISSN 2150-3508. URL https:

//www.longdom.org/open-access/investigating-feeding-regimen-
of-brown-trout-fario-in-tonekabon-river-northern-iran.pdf.

Justin:2022:BNF

[285] Pablo Justin. Brief note on fish oil. Fisheries and Aquaculture Journal, 13(2):1, ???? 2022. CODEN FAJIAO. ISSN 2150-3508. URL https://

www.longdom.org/open-access/brief-note-on-fish-oil.pdf.

Pavel:2022:PHH

[286] E Garlov Pavel. Participation of the hypothalamic–hypophysial neurosecretory system in the occurrence of a migration impulse on fish. Fisheries and Aquaculture Journal, 13(2):1–6, ???? 2022. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
participation-of-the-hypothalamichypophysial-neurosecretory-
system-in-the-occurrence-of-a-migration-impulse-on-fish.pdf.

Chalchisa:2022:OFF

[287] Teferi Chalchisa. An overview of fermented fish products. Fisheries and Aquaculture Journal, 13(2):1–2, ???? 2022. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/an-overview-of-
fermented-fish-products.pdf.

VS:2022:STBa

[288] Ajay VS. Sharp teeth, body full of thorns; antenna rising high on forehead, strange devil fish landing on the shore. Fisheries and Aquaculture Journal, 13(2):1, ???? 2022. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/sharp-teeth-

body-full-of-thorns-antenna-rising-high-on-forehead-strange-
devil-fish-landing-on-the-shore.pdf.

Olaniyi:2022:OTU

[289] Oluwafemi Zaccheaus Olaniyi. An overview of techniques used in aquaculture. Fisheries and Aquaculture Journal, 13(3):1–2, ???? 2022. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/an-overview-of-techniques-used-in-aquaculture.pdf.

Greenhalgh:2022:BNO

[290] Colin A. Greenhalgh. Brief note on oyster culture. Fisheries and Aquaculture Journal, 13(3):1, ???? 2022. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/brief-note-on-
oyster-culture.pdf.

Khatua:2022:SFF

[291] Rohit Khatua. A synopsis of fish farming. Fisheries and Aquaculture Journal, 13(3):1–2, ???? 2022. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/a-synopsis-of-
fish-farming.pdf.

Justin:2022:HMS

[292] Pablo Justin. Harvesting methods in shrimp farming. Fisheries and Aquaculture Journal, 13(3):1–2, ???? 2022. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/harvesting-
methods-in-shrimp-farming.pdf.

Ariyati:2022:NSA

[293] Armitage Ariyati. Note on sustainable aquaculture. Fisheries and Aquaculture Journal, 13(3):1–2, ???? 2022. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/note-on-sustainable-
aquaculture.pdf.

Oliver:2022:RVC

[294] Daniel Oliver. Role of vaccines in control and prevention of infectious disease in aquaculture. Fisheries and Aquaculture Journal, 13(4):1– 2, ???? 2022. CODEN FAJIAO. ISSN 2150-3508. URL https://

www.longdom.org/open-access/role-of-vaccines-in-control-and-
prevention-of-infectious-disease-in-aquaculture.pdf.

Howell:2022:TBP

[295] Ryan Howell. Types of bacterial pathogens in aquaculture. Fisheries and Aquaculture Journal, 13(4):1–2, ???? 2022. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/types-of-
bacterial-pathogens-in-aquaculture.pdf.

Olaniyi:2022:APA

[296] Oluwafemi Zaccheaus Olaniyi. Applications of probiotics in aquaculture. Fisheries and Aquaculture Journal, 13(4):1, ???? 2022. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
applications-of-probiotics-in-aquaculture.pdf.

Khatua:2022:BMU

[297] Rohit Khatua. Biosecurity measures used in aquaculture. Fisheries and Aquaculture Journal, 13(4):1, ???? 2022. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
biosecurity-measures-used-in-aquaculture.pdf.

Tang:2022:ELC

[298] Fujiang Tang. Evidence from living creature: The Cope’s Rule. Fisheries and Aquaculture Journal, 13(4):1–2, ???? 2022. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/evidence-
from-living-creature-the-copes-rule.pdf.

Chalchisa:2022:IFP

[299] Teferi Chalchisa. Impact of foxhunt production on women in fishing community. Fisheries and Aquaculture Journal, 13(4):1–2, ???? 2022. CODEN FAJIAO. ISSN 2150-3508. URL https://

www.longdom.org/open-access/impact-of-foxhunt-production-on-
women-in-fishing-community.pdf.

Islam:2022:PSC

[300] Saiful Islam, MA Azadi, Munira Nasiruddin, and Muhammad Mohiuddin Sarker. Plankton species composition, abundance and diversity indices in three ponds of Chittagong University Campus, Bangladesh. Fisheries and Aquaculture Journal, 13(4):1–9, ???? 2022. CODEN

FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/plankton-species-composition-abundance-and-diversity-
indices-in-threeponds-of-chittagong-university-campus-bangladesh.
pdf.

Takele:2022:UHP

[301] Shitaw Takele, Akwake Geremew, and Abebe Getahun. Unlocking the hidden potential of plant proteins in fish nutrition using technology. Fisheries and Aquaculture Journal, 13(5):1–6, ???? 2022. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/unlocking-
the-hidden-potential-of-plant-proteins-in-fish-nutrition-using-
technology.pdf.

Sahu:2022:EAF

[302] Rohit Kumar Sahu, Ajay Tegar, and Mukesh Kumar Anant. An economic analysis of fish production of SHG’s and fishermen cooperative groups in Mungeli District of Chhattisgarh. Fisheries and Aquaculture Journal, 13(5):1–4, ???? 2022. CODEN FAJIAO. ISSN 2150-3508. URL

https://www.longdom.org/open-access/an-economic-analysis-of-
fish-production-of-shgs-and-fishermen-cooperative-groups-in-
mungeli-district-of-chhattisgarh.pdf.

Pilli:2022:PWQ

[303] Saif Pilli. Probiotics in water quality management and disease resistance in P. vannamei farms of Razole, East Godavari, Andhra Pradesh. Fisheries and Aquaculture Journal, 13(5):1–10, ???? 2022. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/probiotics-
in-water-quality-management-and-disease-resistance-in-p-vannamei-
-farms-of-razole-east-godavari-andhra-prades.pdf.

Chalchisa:2022:BHI

[304] Teferi Chalchisa. Bacteria host interactions in columnaris disease. Fisheries and Aquaculture Journal, 13(5):1, ???? 2022. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
bacteria-host-interactions-in-columnaris-disease.pdf.

Dumith:2022:ISR

[305] Michelle Torres Dumith and Alejandra FGN Santos. Ichthyofauna structure at risk due to the bio invasion of Clarias gariepinus in a river at Guapimirim Environmental Protection Area, Southeastern Brazil. Fisheries and Aquaculture Journal, 13(5):1–10, ???? 2022. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
ichthyofauna-structure-at-risk-due-to-the-bio-invasion-of-clarias-
gariepinus-in-a-river-at-guapimirim-environmental-prot.pdf.

Olaniyi:2022:NUP

[306] Oluwafemi Zaccheaus Olaniyi. Note on uses of probiotics in aquaculture. Fisheries and Aquaculture Journal, 13(5):1, ???? 2022. CODEN FAJIAO.

ISSN 2150-3508. URL https://www.longdom.org/open-access/note-
on-uses-of-probiotics-in-aquaculture.pdf.

Mwainge:2022:OPF

[307] Venny M. Mwainge, Nicholas Outa, Caleb Ogwai, Veronica Ombwa, Nathan Lenjo, Evans Abich, and Kennedy Oyier. Occurrence and prevalence of fish parasites and the interaction with water quality parameters in selected small water bodies in Western Kenya. Fisheries and Aquaculture Journal, 13(6):1–6, ???? 2022. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/occurrence-and-
prevalence-of-fish-parasites-and-the-interaction-with-water-
quality-parameters-in-selected-small-water-bo.pdf.

Alim:2022:RAA

[308] Fauzan Syaiful Alim. Review on the analysis and application of blue economy in Indonesia. Fisheries and Aquaculture Journal, 13(6):1–4, ???? 2022. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/review-on-the-analysis-and-application-of-blue-
economy-in-indonesia.pdf.

Chalchisa:2022:BIM

[309] Teferi Chalchisa. Benefits of integrated multi-trophic aquaculture systems. Fisheries and Aquaculture Journal, 13(6):1, ???? 2022. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
benefits-of-integrated-multitrophic-aquaculture-systems.pdf.

Stryjecki:2022:ADT

[310] Rodgers M. Stryjecki. Advantages and different techniques of fish farming in aquaculture industry. Fisheries and Aquaculture Journal, 13(6):1–2, ???? 2022. CODEN FAJIAO. ISSN 2150-3508. URL https://www.

longdom.org/open-access/advantages-and-different-techniques-
of-fish-farming-in-aquaculture-industry.pdf.

Ahmed:2022:IPE

[311] Aziz Ahmed, Rehana Kauser, Hsina Basharat, Ramzan Ali, and Anila Naz Soomro. Evaluation of sperm mobility activation media for egg fertilization in locally bred American channel catfish (Ictalurus punctatus). Fisheries and Aquaculture Journal, 13(6):1–5, ???? 2022. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
evaluation-of-sperm-mobility-activation-media-for-egg-fertilization-
in-locally-bred-american-channel-catfish-ictalurus-p.pdf.

Tareke:2022:DSJ

[312] Abebe Tesfaye Tareke, Abebe Getahun Gubale, and Tadesse Fetahi Hailu. Dietary shifts in juvenile and adult Nile tilapia, Oreochromis niloticus (L.) (Pisces: Cichlidae) in Lake Ziway, Ethiopia. Fisheries and Aquaculture Journal, 13(6):1–6, ???? 2022. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/dietary-shifts-in-
juvenile-and-adult-nile-tilapia-oreochromis-niloticus-lpisces-
cichlidae-in-lake-ziway-ethiopia.pdf.

Kela:2022:EPC

[313] Esther Kela, Amos O. Sogbesan, and Umar Buba Wakil. Evaluation of phytochemical composition of ginger extracts. Fisheries and Aquaculture Journal, 13(6):1–5, ???? 2022. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/evaluation-of-
phytochemical-composition-of-ginger-extracts.pdf.

VS:2022:STBb

[314] Ajay VS. Sharp teeth, body full of thorns; antenna rising high on forehead, strange devil fish landing on the shore. Fisheries and Aquaculture Journal, 13(6):1, ???? 2022. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/sharp-teeth-
body-full-of-thorns-antenna-rising-high-on-forehead-strange-
devil-fish-landing-on-the-shore.pdf.

Chalchisa:2023:BNC

[315] Teferi Chalchisa. Brief note on Catla catla and its ecological importance. Fisheries and Aquaculture Journal, 14(1):1–2, ???? 2023. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
brief-note-on-catla-catla-and-its-ecological-importance.pdf.

Justin:2023:ARS

[316] Pablo Justin. Aquaculture’s role in sustainable development, building communities and protecting oceans. Fisheries and Aquaculture Journal, 14(1):1–2, ???? 2023. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/aquacultures-role-
in-sustainable-development-building-communities-and--protecting-
oceans.pdf.

Stryjecki:2023:DAS

[317] Rodgers M Stryjecki. Diversification and adaptability study of catfish. Fisheries and Aquaculture Journal, 14(1):1, ???? 2023. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
diversification-and-adaptability-study-of-catfish.pdf.

Chalchisa:2023:CIC

[318] Teferi Chalchisa. Challenges and innovations for conservation of cod fish aquatic biota. Fisheries and Aquaculture Journal, 14(1):1, ???? 2023. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/challenges-and-innovations-for-conservation-
of-cod-fish-aquatic-biota.pdf.

Tefiear:2023:CEI

[319] Chalchis Tefiear. Characters and ecological importance of shoals fish. Fisheries and Aquaculture Journal, 14(1):1, ???? 2023. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
characters-and-ecological-importance-of-shoals-fish.pdf.

Oliver:2023:CTE

[320] Daniel Oliver. Cultivation techniques and economic implications of eel fish in aquaculture. Fisheries and Aquaculture Journal, 14(1):1, ???? 2023. CODEN FAJIAO. ISSN 2150-3508. URL https://www.

longdom.org/open-access/cultivation-techniques-and-economic-
implications-of-eel-fish-in-aquaculture.pdf.

Islam:2023:PSC

[321] Saiful Islam, MA Azadi, Munira Nasiruddin, and Muhammad Mohiuddin Sarker. Plankton species composition, abundance and diversity indices in three ponds of Chittagong University Campus, Bangladesh. Fisheries and Aquaculture Journal, 14(1):1–9, ???? 2023. CODEN

FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/plankton-species-composition-abundance-and-diversity-
indices-in-threeponds-of-chittagong-university-campus-bangladesh.
pdf.

Hailu:2023:ACS

[322] Addisu Hailu and Alemayehu Abebe. Assessment of current status of fishermen cooperatives at selected Oromia water bodies, Ethiopia. Fisheries and Aquaculture Journal, 14(1):1–19, ???? 2023. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/assessment-
of-current-status-of-fishermen-cooperatives-at-selected-oromia-
water-bodies-ethiopia.pdf.

Olaniyi:2023:VCC

[323] Oluwafemi Zaccheaus Olaniyi, Ayanboye Abolupe Oluyemisi, and Oluwafemi Benedicta Motunrayo. Value chain of catfish products in Ibadan Metropolis, Oyo State, Nigeria. Fisheries and Aquaculture Journal, 14(1):1–6, ???? 2023. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/value-chain-of-
catfish-products-in-ibadan-metropolis-oyo-state-nigeria.pdf.

Olaniyi:2023:NUP

[324] Oluwafemi Zaccheaus Olaniyi. Note on uses of probiotics in aquaculture. Fisheries and Aquaculture Journal, 14(1):1, ???? 2023. CODEN FAJIAO.

ISSN 2150-3508. URL https://www.longdom.org/open-access/note-
on-uses-of-probiotics-in-aquaculture.pdf.

Ehsan:2023:EIC

[325] Ali Carl Ehsan. Economic importance and conservation of red snapper. Fisheries and Aquaculture Journal, 14(2):1–2, ???? 2023. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
economic-importance-and-conservation-of-red-snapper.pdf.

Shah:2023:EIA

[326] Md Imran Shah. Ecological importance of aquatic animals and threats faced by aquatic species. Fisheries and Aquaculture Journal, 14(2):1, ???? 2023. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/ecological-importance-of-aquatic-animals-and-
threats-faced-by-aquatic-species.pdf.

Chalchisa:2023:EFA

[327] Teferi Chalchisa. External features and adaptations of crabs to their surrounding environment. Fisheries and Aquaculture Journal, 14(2):1–2, ???? 2023. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/external-features-
and-adaptations-of-crabs-to-their-surrounding-environment.pdf.

Ehsan:2023:RES

[328] Ali Carl Ehsan. Reproduction and ecology of seahorse. Fisheries and Aquaculture Journal, 14(2):1, ???? 2023. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/reproduction-
and-ecology-of-seahorse.pdf.

Shah:2023:CFF

[329] Md Imran Shah. Conservation of freshwater fishes and their ecological significance. Fisheries and Aquaculture Journal, 14(2):1, ???? 2023. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/conservation-of-freshwater-fishes-and-their-
ecological-significance.pdf.

Sam:2023:SBA

[330] Louis Sam. Squids behavioral adaptations for predatory life. Fisheries and Aquaculture Journal, 14(2):1, ???? 2023. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
squids-behavioral-adaptations-for-predatory-life.pdf.

Justin:2023:CRS

[331] Pablo Justin. Characteristics and reproductive system of the octopus. Fisheries and Aquaculture Journal, 14(2):1, ???? 2023. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
characteristics-and-reproductive-system-of-the-octopus.pdf.

Olaniyi:2023:BNH

[332] Oluwafemi Zaccheaus Olaniyi. Brief note on halibut characteristics and anatomy. Fisheries and Aquaculture Journal, 14(2):1, ???? 2023. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/brief-note-on-halibut-characteristics-and-anatomy.pdf.

Stryjeckin:2023:MBP

[333] Rodgers M. Stryjeckin. Mariculture benefits in promoting sustainable aquatic food production for a growing world. Fisheries and Aquaculture Journal, 14(2):1–2, ???? 2023. CODEN FAJIAO. ISSN 2150-3508. URL

https://www.longdom.org/open-access/mariculture-benefits-in-
promoting-sustainable-aquatic-food-production-for-a--growing-
world.pdf.

Stryjecki:2023:FST

[334] Rodgers M Stryjecki. Fish scales: Types and waste management. Fisheries and Aquaculture Journal, 14(3):1–2, ???? 2023. CODEN FAJIAO.

ISSN 2150-3508. URL https://www.longdom.org/open-access/fish-
scales-types-and-waste-management.pdf.

Chalchisa:2023:FAF

[335] Teferi Chalchisa. Functions and adaptations of fish fins. Fisheries and Aquaculture Journal, 14(3):1, ???? 2023. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/functions-and-
adaptations-of-fish-fins.pdf.

Oliver:2023:AFS

[336] Daniel Oliver. Anatomy and functions of the swim bladder. Fisheries and Aquaculture Journal, 14(3):1–2, ???? 2023. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
anatomy-and-functions-of-the-swim-bladder.pdf.

Olaniyi:2023:SSB

[337] Oluwafemi Zaccheaus Olaniyi. Significance of schooling behavior in fishes. Fisheries and Aquaculture Journal, 14(3):1, ???? 2023. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
significance-of-schooling-behavior-in-fishes.pdf.

Justin:2023:BNC

[338] Pablo Justin. Brief note on characteristics and behaviour of pomfret fish. Fisheries and Aquaculture Journal, 14(3):1, ???? 2023. CODEN FAJIAO.

ISSN 2150-3508. URL https://www.longdom.org/open-access/brief-
note-on-characteristics-and-behaviour-of-pomfret-fish.pdf.

Stryjecki:2023:MTA

[339] Rodgers M. Stryjecki. Morphological traits of apex predator barracuda. Fisheries and Aquaculture Journal, 14(3):1–2, ???? 2023. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
morphological-traits-of-apex-predator-barracuda.pdf.

Islam:2023:OTF

[340] Saiful Islam. An overview on tuna fish: Diversity and sustainability. Fisheries and Aquaculture Journal, 14(3):1, ???? 2023. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
an-overview-on-tuna-fish-diversity-and-sustainability.pdf.

Sam:2023:SIF

[341] Louis Sam. Structure and important functions of fish gills. Fisheries and Aquaculture Journal, 14(3):1, ???? 2023. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/structure-
and-important-functions-of-fish-gills.pdf.

Pawar:2023:IMT

[342] Lokesh Pawar, Mayuri Nag, and M. Junaid Sidiq. Integrated multitrophic aquaculture systems (IMTA)-A sustainable approach for better resource utilization. Fisheries and Aquaculture Journal, 14(3):1–4, ???? 2023. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/integrated-multitrophic-aquaculture-systems-
imtaa-sustainable-approach-for-better-resource-utilization.pdf.

Yesuf:2023:PSI

[343] Arebu Yesuf, Shimeles Abegaz, Nesibu Awol, and Getachew Gugsa. Prevalence study of internal and external parasitic disease of fish in selected lakes of South Wollo Zone, North East Ethiopia. Fisheries and Aquaculture Journal, 14(3):1–8, ???? 2023. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/prevalence-study-of-
internal-and-external-parasitic-disease-of-fish-in-selected-
lakes-of-south-wollo-zone-north-east-eth.pdf.

Ahmad:2023:CAF

[344] Ashfaq Ahmad. Comparative analysis of fish fauna of District Mardan and Swabi, Khyber Pakhtunkhwa, Pakistan. Fisheries and Aquaculture Journal, 14(3):1–2, ???? 2023. CODEN FAJIAO. ISSN 2150-3508. URL

https://www.longdom.org/open-access/comparative-analysis-of-
fish-fauna-of-district-mardan-and-swabi-khyber-pakhtunkhwa-pakistan.
pdf.

Yadav:2023:FFP

[345] Manish Yadav, Rahul Ranjan, and Tapendra Bahadur Shah. Fish farming practices and disease occurrence in the fish farms of Dhanusha District, Nepal. Fisheries and Aquaculture Journal, 14(3):1–18, ???? 2023. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/fish-farming-practices-and-disease-occurrence-in-the-
fish-farms-of-dhanusha-district-nepal.pdf.

Sam:2023:NBA

[346] Iouis Sam. Note on benefits of aquaculture. Fisheries and Aquaculture Journal, 14(3):1, ???? 2023. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/analysis-of-
food-and-feeding-habits-of-channa-punctata--from-ashihali-beel-
of-hailakandi-district-of-assam-india.pdf.

Sravani:2023:ADF

[347] Savva Sravani and S Ravichandran. An article on disease free and pollution free ponds. Fisheries and Aquaculture Journal, 14(3):1, ???? 2023. CODEN FAJIAO. ISSN 2150-3508. URL https:/

/www.longdom.org/open-access/an-article-on-disease-free-and-
pollution-free-ponds.pdf.

Sackey:2021:OOM

[348] Anthony Djaba Sackey, Benjamin Lantei Lamptey, Patrick Ofori-Danson, Beshiru Yahaya, and Abigail Dede Sackey. Outlining the offshore marine environment and mammal habitation with the changing climate in the Gulf of Guinea: The case of Ghana’s waters. Fisheries and Aquaculture Journal, 0(0):1–12, ???? 2021. CODEN

FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/outlining-the-offshore-marine-environment-and-mammal-
habitation-with-the-changing-climate-in-the-gulf-of-guinea-the-

case.pdf. Special issue on Biological and Environmental Factors Effecting Marine Organisms.

Carmel:2021:CTF

[349] Ivanov Carmel. A commentary on tuna fish farming. Fisheries and Aquaculture Journal, 0(0):1, ???? 2021. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/a-commentary-

on-tuna-fish-farming.pdf. Special issue on Biological and Environmental Factors Effecting Marine Organisms.

Carmel:2021:RBT

[350] Ivanov Carmel. Role of bio toxins in marine life. Fisheries and Aquaculture Journal, 0(0):1, ???? 2021. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/role-of-

bio-toxins-in-marine-life.pdf. Special issue on Biological and Environmental Factors Effecting Marine Organisms.

Ariyati:2021:NFD

[351] Armitage Ariyati. A note on fish disease and parasites. Fisheries and Aquaculture Journal, 0(0):1, ???? 2021. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/a-note-on-fish-

disease-and-parasites.pdf. Special issue on Biological and Environmental Factors Effecting Marine Organisms.

Ariyati:2021:NMB

[352] Armitage Ariyati. A note on marine biology. Fisheries and Aquaculture Journal, 0(0):1, ???? 2021. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/a-note-on-

marine-biology.pdf. Special issue on Biological and Environmental Factors Effecting Marine Organisms.

Williams:2021:ECC

[353] Emilio Dayani Williams. Effects of climate changes on coral. Fisheries and Aquaculture Journal, 0(0):1, ???? 2021. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/effects-of-

climate-changes-on-coral.pdf. Special issue on Biomass, Bioproduct and Bioenergy From Marine Based Industry.

Greenhalgh:2021:CEO

[354] Colin A. Greenhalgh. A commentary on electric organs in fishes. Fisheries and Aquaculture Journal, 0(0):1, ???? 2021. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/a-commentary-
on-electric-organs-in-fishes.pdf. Special issue on Biomass, Bio-

product and Bioenergy From Marine Based Industry.

Greenhalgh:2021:BTT

[355] Colin A. Greenhalgh. Behavioural toxicology and teleost fish. Fisheries and Aquaculture Journal, 0(0):1, ???? 2021. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/behavioural-
toxicology-and-teleost-fish.pdf. Special issue on Biomass, Bioprod-

uct and Bioenergy From Marine Based Industry.

Williams:2021:CFB

[356] Emilio Dayani Williams. A commentary on fish bacterial and viral diseases. Fisheries and Aquaculture Journal, 0(0):1–2, ???? 2021. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/a-commentary-on-fish-bacterial-and-viral-diseases.pdf.

Special issue on Biomass, Bioproduct and Bioenergy From Marine Based Industry.

Autier:2021:TFL

[357] Vincent Autier and J. A. Heindel. Technical fishway limitations and common misconceptions. Fisheries and Aquaculture Journal, 0(0):1– 9, ???? 2021. CODEN FAJIAO. ISSN 2150-3508. URL https:

//www.longdom.org/open-access/technical-fishway-limitations-

and-common-misconceptions.pdf. Special issue on Biomass, Bioproduct and Bioenergy From Marine Based Industry.

deSouzaMotta:2021:FAP

[358] Jonas Henrique de Souza Motta, Leonardo Siqueira Glória, André Batista de Souza, Marcelo Fanttini Polese, Fernando Bosisio, and Manuel Vazquez Vidal. Freshwater angelfish Pterophyllum scalare (Lichtenstein; Pisces: Cichlidae) phenotypic plasticity during fasting periods. Fisheries and Aquaculture Journal, 0(0):1–3, ???? 2021. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/freshwater-
angelfish-pterophyllum-scalare-lichtenstein-pisces-cichlidae-
phenotypic-plasticity-during-fasting-periods.pdf. Special issue

on Climatic Changes and Its Impact On Ocean Bodies.

Makaras:2021:IAS

[359] Tomas Makaras and Milda Stankevičiūtė. The impact of acclimation on the swimming behaviour of salmonids (Oncorhynchus mykiss and Salmo salar). Fisheries and Aquaculture Journal, 0(0):1, ???? 2021. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
the-impact-of-acclimation-on-the-swimming-behaviour-of-salmonids-
oncorhynchus-mykiss-and-salmo-salar.pdf. Special issue on Climatic

Changes and Its Impact On Ocean Bodies.

Justin:2021:MEP

[360] Pablo Justin. Marine ecology and production process. Fisheries and Aquaculture Journal, 0(0):1, ???? 2021. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/marine-ecology-

and-production-process.pdf. Special issue on Climatic Changes and Its Impact On Ocean Bodies.

Justin:2021:CPH

[361] Pablo Justin. A commentary on ponds harvesting techniques. Fisheries and Aquaculture Journal, 0(0):1, ???? 2021. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
a-commentary-on-ponds-harvesting-techniques.pdf. Special issue on

Climatic Changes and Its Impact On Ocean Bodies.

Ariyati:2021:ENA

[362] Armitage Ariyati. An editorial note on aquatic ecosystem. Fisheries and Aquaculture Journal, 0(0):1, ???? 2021. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/an-editorial-

note-on-aquatic-ecosystem.pdf. Special issue on Climatic Changes and Its Impact On Ocean Bodies.

Ajay:2021:DCS

[363] V. S. Ajay and A. R. Krishnan. Design characteristics and specifications of cast net operated along the lower stretches of Vembanad Wetlands. Fisheries and Aquaculture Journal, 0(0):1–9, ???? 2021. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
design-characteristics-and-specifications-of-cast-net-operated-
along-the-lowerstretches-of-vembanad-wetlands.pdf. Special issue

on Genetic Parameters in Aquaculture.

Jiang:2021:MTG

[364] Hui Gong Jiang, Jarupan Channarong, Lertluk Ngernsiri, and Akarapong Swatdipong. Microsatellite techniques in Guam’s specific-pathogen-free Penaeus vannamei stock: Genetic variance and parentage identification. Fisheries and Aquaculture Journal, 0(0):1–12, ???? 2021. CODEN FA-

JIAO. ISSN 2150-3508. URL https://www.longdom.org/open-access/
microsatellite-techniques-in-guams-specificpathogenfree-penaeus-
vannamei-stock-genetic-variance-and-parentage-identifica.pdf.

Special issue on Genetic Parameters in Aquaculture.

Mamuru:2021:SSF

[365] Gobeng Likambo Seme Mamuru, Daud Bona Lubani, Flora Magara Emmanuel, Rafile Talamuk, Johnson Jiribi Bali, John Sebit Benansio, and Ngoepe Tlou Kevin. South Sudan fish processing and value chain analysis of Juba City. Fisheries and Aquaculture Journal, 0(0):1–13, ???? 2021. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/south-sudan-fish-processing-and-value-chain-

analysis-of-juba-city.pdf. Special issue on Genetic Parameters in Aquaculture.

Ehsan:2021:OA

[366] Ali Carl Ehsan. Organic aquaculture. Fisheries and Aquaculture Journal, 0(0):1–2, ???? 2021. CODEN FAJIAO. ISSN 2150-3508. URL https:

//www.longdom.org/open-access/organic-aquaculture.pdf. Special

issue on Genetic Parameters in Aquaculture.

Ehsan:2021:MPA

[367] Ali Carl Ehsan and Ali Carl Ehsan. Marine protected areas help fisheries and ocean ecosystems. Fisheries and Aquaculture Journal, 0(0):1–2, ???? 2021. CODEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.

org/open-access/marine-protected-areas-help-fisheries-and-ocean-

ecosystems.pdf. Special issue on Genetic Parameters in Aquaculture.

Endebu:2021:DPC

[368] Megerssa Endebu, Abebe Getahun, and Misikire Tessema. Differences in phenotypic characters of Nile tilapia (Oreochromis niloticus L.) in three Ethiopian Rift Valley lakes; screening strains for aquaculture. Fisheries and Aquaculture Journal, 0(0):1–6, ???? 2021. CODEN

FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/differences-in-phenotypic-characters-of-nile-tilapia-
oreochromis-niloticus-l-in-three-ethiopian-rift-valley-lakes-

chamo-.pdf. Special issue on Microorganisms in aquaculture.

Ajay:2021:IDV

[369] V. S. Ajay. Ichthyofaunal diversity in the Varapuzha Wetlands of Vembanad Lake, Kerala, India: Comprehensive study on the living status, biodiversity assessment and fishing methods. Fisheries and Aquaculture Journal, 0(0):1–9, ???? 2021. CODEN FAJIAO. ISSN 2150-

3508. URL https://www.longdom.org/open-access/ichthyofaunal-
diversity-in-the-varapuzha-wetlands-of-vembanad-lake-kerala-
india-comprehensive-study-on-the-living-status.pdf. Special is-

sue on Microorganisms in aquaculture.

Jayawardena:2021:ICA

[370] A. W. Jayawardena. International conference on aquaculture research and fisheries. Fisheries and Aquaculture Journal, 0(0):1, ???? 2021. CO-

DEN FAJIAO. ISSN 2150-3508. URL https://www.longdom.org/open-
access/international-conference-on-aquaculture-research-and-

fisheries.pdf. Special issue on Microorganisms in aquaculture.

Ehsan:2021:FTM

[371] Ali Carl Ehsan. Fishing technology and methods. Fisheries and Aquaculture Journal, 0(0):1, ???? 2021. CODEN FAJIAO. ISSN

2150-3508. URL https://www.longdom.org/open-access/fishing-

technology-and-methods.pdf. Special issue on Microorganisms in aquaculture.

Ehsan:2021:MAE

[372] Ali Carl Ehsan. Market analysis of European Summit on Aquaculture, Fisheries and Horticulture. Fisheries and Aquaculture Journal, 0(0):1–2, ???? 2021. CODEN FAJIAO. ISSN 2150-3508.

URL https://www.longdom.org/open-access/market-analysis-of-
european-summit-on-aquaculture-fisheries-and-horticulture.pdf.

Special issue on Microorganisms in aquaculture.